### v1

# Invariants for $\mathbb{G}_{(r)}$-modules

Abstract.
Following earlier work of C. Bendel, J. Petvsova, P. Sobaje, A. Suslin, and the author, we further investigate refined invariants for finite dimensional representations $M$ of infinitesimal group schemes $\mathbb{G}$ over a field $k$ of characteristic $p>0$. In particular, we explore invariants of modules for Frobenius kernels $\mathbb{G}_{(r)}$ of linear algebraic groups $\mathbb{G}$ of exponential type. Such finite group schemes admit a refinement of support theory, including Jordan type functions and constructions of vector bundles for special classes of $\mathbb{G}_{(r)}$-modules. We reformulate earlier results for $\mathbb{G}_{(r)}$-modules in terms which are both more explicit and more accessible to computation, yet preserve the usual support theory for $\mathbb{G}_{(r)}$-modules.
Key words and phrases:
Jordan type, infinitesimal group scheme, vector bundles
2020 Mathematics Subject Classification:
20G05, 14L17
partially supported by the Simons Foundation
0. Introduction
In the present work, we utilize methods and results of support theory to add to our understanding of invariants finer than support theory for certain classes of infinitesimal group schemes over an arbitrary base field $k$ of characteristic $p$. This is in contrast to efforts devoted to establish properties of one or more categories of representations of such a finite group scheme. Our techniques do not apply to finite groups, for we utilize the support theory developed in collaboration with A. Suslin and C. Bendel in [22], [23] which replaces cohomological varieties by varieties of 1-parameter subgroups.
In particular, we investigate the “local Jordan type function” as briefly considered by J. Pevtsova and the author in [13, Defn 2.3]. For a given affine group scheme $\mathbb{G}$ of finite type over $k$ and a finite dimensional $\mathbb{G}$-module $M$ for the $r$-th Frobenius kernel $\mathbb{G}_{(r)}$ of $\mathbb{G}$, we formulate the continuous function  $JT_{\mathbb{G},r,M}(-):\mathbb{P}V_{r}(\mathbb{G})\to\mathcal{Y}$ from the scheme-theoretic points of $\mathbb{P}V_{r}(\mathbb{G})$ representing height $r$ 1-parameter subgroups to the partially ordered set $\mathcal{Y}$ of Young diagrams with $p$ columns. This “theory” $M\mapsto JT_{\mathbb{G},r,M}(-)$ captures much more information about a $\mathbb{G}_{(r)}$-module $M$ than does the support theory $M\mapsto\mathbb{P}V_{r}(\mathbb{G})_{M}$, yet possesses many good properties satisified by the more familiar cohomological support theory $M\mapsto\mathbb{P}H^{*}(\mathbb{G}_{(r)},k)_{M}$. For example, if $p$ does not divide the dimension of $M$, then $\mathbb{P}H^{*}(\mathbb{G}_{(r)},k)_{M}$ equals $\mathbb{P}H^{*}(\mathbb{G}_{(r)},k)$ and therefore is independent of $M$, whereas $M\ \mapsto\ JT_{\mathbb{G},r,M}(-)$ distinguishes many such $\mathbb{G}_{(r)}$-modules $M$ of a given dimension.
As a prelude to establishing the continuity of $JT_{\mathbb{G},r,M}(-)$, we reprove from first principles the compatibility of specialization of $M$ at a scheme-theoretic point of $\mathbb{P}V_{r}(\mathbb{G})$ with specialization of the Jordan type of $M$ at that scheme-theoretic point. A central role in this present work is played by the “universal $p$-nilpotent operator” $\Theta_{\mathbb{G},r}\in k[V_{r}(\mathbb{G})]\otimes k\mathbb{G}_{(r)}$ constructed by Pevtsova and the author in [14]. Our first use of $\Theta_{\mathbb{G},r}$ is to complete the proof of the continuity of the function $JT_{\mathbb{G},r,M}(-)$ using the “projective extension” $\tilde{\Theta}_{\mathbb{G},r}$ of $\Theta_{\mathbb{G},r}$.
In order to make the Jordan type function more explicit and computationally accessible, we restrict our attention to affine group schemes $\mathbb{G}$ which are of exponential type of height $r$; for such $\mathbb{G}$, height $r$ 1-parameter subgroups of $\mathbb{G}$ are parametrized by the variety $\mathcal{C}_{r}(\mathfrak{g})$ of $r$-tuples of $p$-nilpotent, pair-wise commuting elements of the Lie algebra $\mathfrak{g}$ of $\mathbb{G}$. This enables us to work with the explicitly defined variety $\mathcal{C}_{r}(\mathfrak{g})$ rather than the scheme $V_{r}(\mathbb{G})$ which is defined by the functor it represents.
For a given finite dimensional $\mathbb{G}_{(r)}$-module $M$, the Jordan type function becomes a continuous function $JT_{\mathfrak{g},r,M}(-):\mathbb{P}\mathcal{C}_{r}(\mathfrak{g})\to\mathcal{Y}$ associated to the “projectivization” $\tilde{\Theta}_{\mathfrak{g},r}$ of the $p$-nilpotent element in $\Theta_{\mathfrak{g},r}\in k[\mathcal{C}_{r}(\mathfrak{g}]\otimes k\mathbb{G}_{(r)}$ corresponding to $\Theta_{\mathbb{G},r}$.
We proceed to simplify this construction, replacing the function $JT_{\mathfrak{g},r,M}(-)$ by its linearization $JT_{\mathfrak{g},r,M}^{exp}(-):\mathbb{P}\mathcal{C}_{r}(\mathfrak{g})\to\mathcal{Y}$ defined by replacing $\Theta_{\mathfrak{g},r}$ by its linear part (denoted $\Theta_{\mathfrak{g},r}^{exp}$) with respect to the grading on $\mathcal{C}_{r}(\mathfrak{g})$ transferred from the natural grading on $V_{r}(\mathbb{G})$. In the very special case that $\mathbb{G}$ equals $\mathbb{G}_{a(r)}$, this simplification is closely related to J. Carlson’s use of cyclic shifted subgroups to identify support varieties for $(\mathbb{Z}/p)^{\times r}$-modules (see [3]). Unlike Carlson’s cyclic shifted subgroups for $(\mathbb{Z}/p)^{\times r}$, $\Theta_{\mathfrak{g}_{a},r}^{exp}$ does not require a choice of generators for the augmentation ideal of $k\mathbb{G}_{a(r)}$.
The suitability of $\Theta_{\mathfrak{g},r}^{exp}$ follows from work of Pevtsova, Suslin, and the author [15] and subsequent work of P. Sobaje [20]. The replacement of $JT_{\mathbb{G},r,M}(-)$ by $JT_{\mathfrak{g},r,M}^{exp}(-):\mathbb{P}\mathcal{C}_{r}(\mathfrak{g})\to\mathcal{Y}$ does not affect certain refined support varieties considered in previous work of Pevtsova and the author in [13] and [14]. Moreover, we verify that vector bundles associated to $\mathbb{G}_{(r)}$-modules $M$ of constant Jordan type using $\tilde{\Theta}_{\mathfrak{g},r}^{exp}$ (the projectivization of $\Theta_{\mathfrak{g},r}^{exp}$) are $\mathbb{A}^{1}$-equivalent to vector bundles previously constructed in [14] using $\tilde{\Theta}_{\mathbb{G},r}$.
We provide sample computations which suggest the accessibility of $JT_{\mathfrak{g},r,M}^{exp}(-)$ for certain $\mathbb{G}_{(r)}$-modules. We also include some relevant discussion of weighted projective varieties and of the connection between 1-parameter subgroups for $\mathbb{G}$ and operators on $\mathbb{G}$-modules.
We thank Julia Pevtsova and Paul Sobaje for their insights into the mathematics we present.
1. Jordan types and specialization
We begin by recalling some basic properties of Jordan forms of $p$-nilpotent endomorphisms.
Definition 1.1.
Let $M$ be a $k[t]/t^{p}$-module of dimension $m$ given by the map $\rho_{M}:k[t]/t^{p}\to M_{m,m}(k)$ of $k$-algebras. Let $\underline{a}=(a_{1},\ldots,a_{p})\in\mathbb{N}^{\times p}$ be defined by setting $a_{i}$ to be the number of blocks of size $i$ of the Jordan canonical form of the matrix $\rho_{M}(t)$. We say that the Jordan type of $M$ is $\underline{a}$, expressing this as $JT(M)=\underline{a}$.
The following proposition recalls familiar interpretations of Jordan types.
Proposition 1.2.
Let $M$ be a $k[t]/t^{p}$-module of dimension $m$ such that $JT(M)=\underline{a}$ (often written $JT(M)=a_{p}[p]+a_{p-1}[p-1]+\cdots+a_{1}[1]$).
  1. (1)
    As a $k[t]/t^{p}$-module, $M$ is isomorphic to  $\bigoplus_{i=1}^{p}(k[t]/t^{i})^{\oplus a_{i}}$.
  2. (2)
    The partition of $m=\sum_{i=1}^{p}a_{i}\cdot i$ with $a_{i}$ subsets of size $i$ is naturally associated to $\underline{a}=(a_{1},\ldots,a_{p})$.
  3. (3)
    The usual partial ordering of partitions determines the partial ordering on ×$\mathbb{N}^{\times p}$ given by
    $\underline{a}\ \leq\ \underline{b}\quad\iff\quad\forall s,1\leq s\leq p,\quad(\sum_{i=s}^{p}a_{i}\cdot i)\leq(\sum_{i=s}^{p}b_{i}\cdot i).$
  4. (4)
    Uniquely associated to $\underline{a}$ is a Young diagram with $p$-columns such that the diagram has $a_{i}$ rows with $i$ boxes and the rows of the diagram are left-adjusted with descending length. Thus, the partial ordering $\underline{a}\ \leq\ \underline{b}$ requires for each $\ell$ that the number of boxes in the top $\ell$ rows of $\underline{a}$ is less or equal to the number of boxes in the top $\ell$ rows of $\underline{b}$.
  5. (5)
    The Jordan type of $t^{j}$ acting on $M\simeq\oplus_{i=1}^{p}(k[t]/t^{i})^{\oplus a_{i}}$ (denoted by $\underline{a}^{j}\in\mathbb{N}^{\times p}$) has associated Young diagram obtained from the Young diagram for $\underline{a}$ by removing the first (i.e., left-most) $j-1$ columns of $\underline{a}$ and adding rows with each having only 1 box so that the number of boxes of $\underline{a}^{j}$ equals $m$.
The lemma below recalls some properties of the partial ordering on Jordan types. We denote by $rk(t^{j}:M\to M)$ the dimension of the subspace $t^{j}(M)\subset M$, the dimension of the image of multiplication by $t^{j}$ on the vector space $M$.
Proposition 1.3.
[…]
  1. (1)
    $rk(t^{s}:M\to M)\ =\ \sum_{i=s+1}^{p}a_{i}(i-1-s)$.
  2. (2)
[…]
  3. (3)
    $\underline{a}\leq\underline{b}\ \iff\underline{a}^{s}\leq\underline{b}^{s},\ \forall s,1\leq s<p$.
  4. (4)
[…]
Proof.
Assertions (1) and (2) are verified by an easy computation. Assertion (3) follows from assertion (2) and Proposition 1.2(3). Assertion (4) follows easily from assertion (2). ∎
Since a surjective map $f:M\to N$ of $k[t]/t^{p}$-modules induces a surjective map $f:t^{j}(M)\to t^{j}(N)$ for any $j$ with $1\leq j<p$, Proposition 1.3(4) implies the following additional corollary.
Corollary 1.4.
Let $f:M\to N$ be a surjective map of finite dimensional $k[t]/t^{p}$-modules. Then  $JT(N)\leq JT(M)$, with equality if and only if $f$ is an isomorphism.
We formulate the partially ordered abelian monoid $\mathcal{Y}_{\leq}$ of “Young diagrams” equipped with a (tensor) product.
Definition 1.5.
Let ×$\mathbb{N}^{\times p}$ denote the monoid of $p$-tuples of non-negative integers $\underline{a}=(a_{1},\ldots,a_{p})$. Denote by $\mathcal{Y}_{\leq}$ the monoid ×$\mathbb{N}^{\times p}$ provided with the partial ordering of Proposition 1.2(3), and denote by $\mathcal{Y}_{\leq,n}$ the partially ordered subset of $\mathcal{Y}_{\leq}$ whose objects are $p$-tuples $\underline{a}$ with $\sum_{i=1}^{p}a_{i}\cdot i$ equal to $n$.
We equip $\mathcal{Y}_{\leq}$ with the bi-additive pairing
$(-)\otimes(-):\ \mathcal{Y}_{\leq}\times\mathcal{Y}_{\leq}\quad\to\quad\mathcal{Y}_{\leq}\ ,$
where $\underline{a}\otimes\underline{b}$ is the $p$-tuple $\underline{c}$ defined by the isomorphism
$\bigoplus_{i=1}^{p}(k[t]/t^{i})^{\oplus c_{i}})\quad\simeq\quad(\bigoplus_{i=1}^{p}(k[t]/t^{i})^{\oplus a_{i}})\ \bigotimes\ (\bigoplus_{i=1}^{p}(k[t]/t^{i})^{\oplus b_{i}})$
of $k[t]t^{p}$-modules (see [5, Cor10.3]).
We recall the lower semi-continuity of Jordan type, an easy consequence of Nakayama’s Lemma (see [15, Prop 3.2]).
Proposition 1.6.
Consider finitely generated field extensions $k\subset K\subset L$ and a noetherian, local domain $R\subset L$ with field of fractions $L$ and residue field $K$. Let $M_{R}$ be a free $R$-module of finite rank equipped with a $p$-nilpotent endomorphism $\theta_{R}\in End_{R}(M_{R},M_{R})$. Denote $K\otimes_{R}M$ by $M_{K}$ and $L\otimes_{R}M$ by $M_{L}$, and set
$\theta_{K}=K\otimes_{R}\theta_{R}\in End_{K}(M_{K},M_{K}),\quad\theta_{L}=L\otimes_{R}\theta_{R}\in End_{L}(M_{L},M_{L}).$
So defined, $M_{K}$ is a $K[t]/t^{p}$-module with the action of $t$ given by $\theta_{K}$ and $M_{L}$ is an $L[t]/t^{p}$-module with the action of $t$ given by $\theta_{L}$. We say that the $L[t]/t^{p}$-module $M_{L}$ specializes along $R$ to the $K[t]/t^{p}$-module $M_{K}$.
For such a specialization,  $JT(M_{K})\ \leq\ JT(M_{L})$.
Proof.
By applying Nakayma’s Lemma [19, Lem4.3] to the cokernels of the maps $(\theta_{R})^{i}:M_{R}\to M_{R}$, we conclude that $dim(coker(\theta_{L}^{i}))\leq dim(coker(\theta_{K}^{i}))$ so that $rk(\theta_{K}^{i})\leq rk(\theta_{L}^{i})$ for each $i,1\leq i<p$. Thus, the proposition follows from Proposition 1.3. ∎
Definition 1.7.
Let $X$ be a scheme of finite type over $k$. For two scheme-theoretic points $x,y$ of $X$, we write $x\leq y$ if $x$ is in the (Zariski) closure of $y$. If $\operatorname{Spec}\nolimits R\subset X$ is an affine open subset of $X$ containing $x$, then the scheme-theoretic point $x\in\operatorname{Spec}\nolimits R$ corresponds to the prime ideal $\mathfrak{p}_{x}\subset R$. The structure sheaf $\mathcal{O}_{X}$ of $X$ has stalk the local ring $R_{(\mathfrak{p}_{x})}$ with residue field $k(x)=R_{(\mathfrak{p}_{x})}/\mathfrak{p}_{x}R_{(\mathfrak{p}_{x})}$. We identify the scheme-theoretic point $x\in X$ with the associated map of schemes $\chi_{x}:\operatorname{Spec}\nolimits k(x)\to X$.
We define $X_{\leq}$ to be the set of scheme-theoretic points of $X$ with this partial ordering.
Lemma 1.8.
Let $x,y$ be two scheme-theoretic points of $X$. Then $x\leq y$ if and only if there exists a local domain $R$ with field of fractions $k(y)$ and residue field $k(x)$ and a morphism $\operatorname{Spec}\nolimits R\to X$ whose restriction to $\operatorname{Spec}\nolimits k(y)$ is $\chi_{y}:\operatorname{Spec}\nolimits k(y)\to X$ and whose restriction to $\operatorname{Spec}\nolimits k(x)$ is $\chi_{x}:\operatorname{Spec}\nolimits k(x)\to X$.
Proof.
We may assume $x\not=y$. Let $\operatorname{Spec}\nolimits A$ be a Zariski open subset of $X$ containing $x$ with corresponding prime $\mathfrak{p}_{x}\subset A$. Since $x$ is in the closure of $y$, the prime ideal $\mathfrak{p}_{y}$ is contained in $\mathfrak{p}_{x}$. Set $S$ equal to the quotient $A/\mathfrak{p}_{y}$ and let $R$ be the localization of $S$ at (the image of ) $\mathfrak{p}_{x}$. So defined, $R$ is a local domain with residue field $k(x)$ and field of fractions $k(y)$. Moreover, $\chi_{x}:\operatorname{Spec}\nolimits k(x)\to X$ and $\chi_{y}:\operatorname{Spec}\nolimits k(y)\to X$ factor through $\operatorname{Spec}\nolimits R\subset\operatorname{Spec}\nolimits S\subset\operatorname{Spec}\nolimits A\subset X$.
Conversely, assume given a local domain $R$ with field of fractions $k(y)$ and residue field $k(x)$ and a morphism $\operatorname{Spec}\nolimits R\to X$ whose restriction to $\operatorname{Spec}\nolimits k(x)$ is $\chi_{x}:\operatorname{Spec}\nolimits k(x)\to X$ and whose restriction to $\operatorname{Spec}\nolimits k(y)$ is $\chi_{y}:\operatorname{Spec}\nolimits k(y)\to X$. Then the kernel of $R\to k(x)$ equals the prime ideal $\mathfrak{p}_{x}$ and contains $\mathfrak{p}_{y}$, the kernel of $R\to k(y)$. Thus, $x\leq y$. ∎
The following proposition will be employed with $A=k[V_{r}(\mathbb{G})]$ as in Definition 2.1 and with the action of $t$ given by various powers of the universal $p$-nilpotent operator of Theorem 2.11. We point out that in this proposition we consider the tensor product over the commutative algebra $A$ of $A$-modules; in discussion below, we shall frequently consider tensor product over $k$ of $\mathbb{G}$-modules using the Hopf algebra structure of $[\mathbb{G}]$.
Proposition 1.9.
Let $A$ be a Noetherian commutative $k$-algebra and consider a finitely generated, projective $A$-module $M_{A}$ equipped with a $p$-nilpotent endomorphism $\theta_{A}\in Hom_{A}(M_{A},M_{A})$. Then sending a scheme-theoretic point $x\in\operatorname{Spec}\nolimits A$ (i.e., a prime ideal $\mathfrak{p}_{x}\subset A$) to the Jordan type of $k(x)\otimes_{A}M_{A}$ as a $k(x)[t]/t^{p}$-module determines a map of posets
(1.9.1) $JT_{A,M_{A}}(-):(\operatorname{Spec}\nolimits A)_{\leq}\quad\to\quad\mathcal{Y}_{\leq},\quad x\mapsto JT(k(x)\otimes_{A}M_{A}).$
If $M_{A},N_{A}$ are finitely generated $A[t]/t^{p}$-modules, then for all $x\in\operatorname{Spec}\nolimits A$
$JT_{A,M_{A}}(x)\oplus JT_{A,N_{A}}(x)\ =\ JT_{A,M_{A}\oplus N_{A}}(x),\quad JT_{A,M_{A}}(x)\otimes JT_{N_{A}}(x)\ =\ JT_{A,M_{A}\otimes_{A}N_{S}}(x).$
Proof.
The assertion that $JT_{A,M_{A}}(-)$ is a map of posets is the assertion that $x\leq y\in\operatorname{Spec}\nolimits A$ implies that $JT_{A,M_{A}}(x)\leq JT_{N,N_{A}}(y)$. As in Lemma 1.8, we may restrict $M_{A}$ along a map from $A$ to $R$ with a local domain with field of fractions $k(y)$ and residue field $k(x)$. Thus, we reduce to the case that $A=R$ is a local domain with field of fractions $k(y)$ and residue field $k(x)$. In this case, Proposition 1.6 applies to prove $JT_{A,M_{A}}(x)\leq JT_{A,M_{A}}(y)$.
The fact that the function $JT_{A,-}(x)$ (sending a finitely generated $A[t]/t^{p}$-module $M_{A}$ to $JT_{A,M_{A}}(x)\in\mathcal{Y}_{\leq}$) respects both addition and tensor product is immediate from the definitions of addition and tensor product for $\mathcal{Y}_{\leq}$. ∎
Proposition 1.9 does not immediately imply that $JT_{A,M_{A}}(-):\operatorname{Spec}\nolimits A\ \to\ \mathcal{Y}_{\leq,m}$ is continuous when $\operatorname{Spec}\nolimits A$ is given the Zariski topology. This is because arbitrary unions of closed subsets (i.e., subsets closed under specialization) are often not closed in the Zariski topology. We supplement Proposition 1.9 with another consequence of Nakayama’s Lemma in commutative algebra. (It seems more convenient for this lemma to refer to the scheme-theoretic points of $\operatorname{Spec}\nolimits A$ as prime ideals $\mathfrak{p}\subset A$.)
Lemma 1.10.
Let $A$ be a Noetherian commutative $k$-algebra and consider a projective $A$-module $M_{A}$ of rank $m$ equipped with a $p$-nilpotent endomorphism $\theta_{A}:M_{A}\to M_{A}$. Then for any $d>0$,
(1.10.1) $C_{d}\ \equiv\ \{\mathfrak{p}\in\operatorname{Spec}\nolimits A:\ rk(\theta_{\mathfrak{p}}:k(\mathfrak{p})\otimes_{A}M_{A}\to k(\mathfrak{p})\otimes_{A}M_{A})\leq d\}$
is a closed subset of $\operatorname{Spec}\nolimits R$.
Proof.
By considering the intersection of $C_{d}$ with each member of an affine open covering of $\operatorname{Spec}\nolimits A$ restricted to which $M_{A}$ is free, it suffices to prove the lemma assuming $M_{A}$ is free. Moreover, if $\mathfrak{q}_{1},\ldots,\mathfrak{q}_{r}$ are the minimal primes of $\operatorname{Spec}\nolimits A$, then the closures of $\{\mathfrak{q}_{i}\}$ are the (closed) irreducible components of $\operatorname{Spec}\nolimits A$. Thus, by considering the intersection of $C_{d}$ with each of these closures, it suffices to assume that $A$ is a domain.
We view $\theta_{A}$ as an element of $M_{m,m}(A)$. Then $C_{d}$ is the closed subset defined as the zero locus of the determinants of all $(d+1\times d+1)$-submatrices of $\theta_{A}$. ∎
We remind the reader that an Alexandrov discrete topology on a $T_{0}$ space is a topology in which the arbitrary union of closed subsets is closed. (See [1] and [18].) The topology on $\mathcal{Y}$ which we utilize in the following theorem is such a topology.
Theorem 1.11.
Let $A$ be a Noetherian commutative $k$-algebra and consider a projective $A$-module $M_{A}$ of rank $m$ equipped with a $p$-nilpotent endomorphism $\theta_{A}$. As usual, we give $\operatorname{Spec}\nolimits A$ the Zariski topology. Give $\mathcal{Y}$ the topology whose closed subsets are those subsets $S\subset\mathcal{Y}$ satisfying the condition that if $s\in S$ and $s^{\prime}\leq s$ then $s^{\prime}\in S$.
Then
$JT_{A,M_{A}}(-):\operatorname{Spec}\nolimits A\quad\to\quad\mathcal{Y}_{\leq,m}$
is continuous.
Proof.
A subset $C\subset\mathcal{Y}_{\leq,m}$ is closed if and only
$C\quad=\quad\bigcup_{\underline{a}\in C}\overline{\{\underline{a}\}},$
where $\overline{\{\underline{a}\}}$ is the closure of the singleton subset $\{\underline{a}\}$. Thus, it suffices to prove that $JT_{A,M_{A}}^{-1}(\overline{\{\underline{a}\}})$ is closed for all $\underline{a}\in\mathcal{Y}_{\leq,m}$ (since $\mathcal{Y}_{\leq,m}$ is finite). By definition, $JT_{A,M_{A}}^{-1}(\overline{\{\underline{a}\}})$ consists of those $x\in\operatorname{Spec}\nolimits A$ such that $JT_{A,M_{A}}(x)\ \leq\ \underline{a}$ in $\mathcal{Y}_{\leq,m}.$ By Proposition 1.3(4), $JT_{A,M_{A}}(x)\ \leq\ \underline{a}$ if and only if the rank of $(\theta_{x})^{s}:k(x)\otimes_{A}M_{A}\to k((x)\otimes_{A}M_{A}$ is less than or equal to the rank of $t^{s}$ on $\oplus_{i=1}^{p}(k[t]/t^{i})^{\oplus a_{i}}$ (which equals $\sum_{i=s+1}^{p}a_{i}\cdot(i-1-s)$) for all $s$, $1\leq s<p$.
Applying Lemma 1.10 to each of $\theta^{s},1\leq s<p$, we conclude the set of those $x\in\operatorname{Spec}\nolimits A$ such that $JT_{A,M_{A}}(x)\ \leq\ \underline{a}$ is closed in $\operatorname{Spec}\nolimits A$. ∎
Remark 1.12.
Considering stable Jordan types (i.e., disregarding blocks of size $p$) can be useful, but has the disadvantage that the stable version of $JT_{A,M_{A}}(-)$ is not continuous.
2. Jordan type functions for $\mathbb{G}$-modules
For an affine group scheme $\mathbb{G}$ of finite type over $k$, we shall be interested in formulation invariants of finite dimensional $\mathbb{G}$-modules. We denote by $F^{r}:\mathbb{G}\to\mathbb{G}^{(r)}$ the $r$-th iterate of the Frobenius map $F:\mathbb{G}\to\mathbb{G}^{(1)}$, and we set
$\mathbb{G}_{(r)}\ \equiv\ ker\{F^{r}:\mathbb{G}\to\mathbb{G}^{(r)}\},$
an infinitesimal group scheme of height $\leq r$ with coordinate algebra $k[\mathbb{G}_{(r)}]$. Thus, $k[\mathbb{G}_{(r)}]$ is a finite dimensional Hopf algebra whose $k$-linear dual we denote by $k\mathbb{G}_{(r)}$.
For any commutative $k$-algebra $R$, we denote by $\mathbb{G}_{R}$ the group scheme $\mathbb{G}\times\operatorname{Spec}\nolimits R$ over $\operatorname{Spec}\nolimits R$ defined as the base change of the affine group scheme $\mathbb{G}$. Thus, the coordinate algebra of $\mathbb{G}_{R}$ equals $R\otimes k[\mathbb{G}]$. If $\mathbb{G}$ is a finite group scheme, then the “group algebra” of $\mathbb{G}_{R}$ (namely, $Hom_{R}(R\otimes k[\mathbb{G}],R)$ equals $R\otimes k\mathbb{G}$.
Definition 2.1.
Let $\mathbb{G}$ be an affine group scheme of finite type over $k$ and let $r$ be a positive integer. As proved in [22, Thm1.5], the functor of commutative $k$-algebras sending an algebra $A$ to the set of maps of group schemes $\mu:\mathbb{G}_{a(r),A}\to\mathbb{G}_{A}$ over $A$ (i.e., the set of height $r$ 1-parameter groups of $\mathbb{G}_{A}$) is representable by the spectrum $V_{r}(\mathbb{G})$ of the finitely generated commutative $k$-algebra $k[V_{r}(\mathbb{G})]$. Each such map $\mu$ factors through $\mathbb{G}_{(r),A}\hookrightarrow\mathbb{G}_{A}$.
Given a $\mathbb{G}_{(r)}$-module $M$ and a scheme theoretic point $x\in V_{r}(\mathbb{G})$, base change provides a $\mathbb{G}_{(r),k(x)}$-module structure on $k(x)\otimes M$. If $\mu_{x}:\mathbb{G}_{a(r),k(x)}\to\mathbb{G}_{(r),k(x)}$ is the 1-parameter subgroup over $k(x)$ parametrized by $\chi_{x}:\operatorname{Spec}\nolimits k(x)\to V_{r}(G)$, then restricting along $\mu_{x}$ determines a $\mathbb{G}_{a(r),k(x)}$-module structure on $k(x)\otimes M$.
In other words, the map of “group algebras” $\mu_{x,*}:k(x)\otimes k\mathbb{G}_{a(r)}\to k(x)\otimes k\mathbb{G}_{(r)}$ determines by restriction along $\mu_{x,*}$ the structure of a $k(x)\otimes k\mathbb{G}_{a(r)}$-module on $k(x)\otimes M$.
Remark 2.2.
A theorem of Suslin, Bendel, and the author ([23, Thm 5.2] asserts that there is a “canonical homomorphism $\psi:H^{ev}(\mathbb{G}_{(r)},k)\to k[V_{r}(\mathbb{G})]$ whose kernel is nilpotent and whose image contains every $p^{r}$-th power. This enables the authors to formulate in terms of 1-parameter subgroups a “rank variety interpretation” of cohomological support varieties for $\mathbb{G}_{(r)}$-modules.
For general infinitesimal group schemes we lack a concrete description of $V_{r}(\mathbb{G})$. In Section 4, we restrict our attentions to $\mathbb{G}$ of exponential type of height $r$ for which we do have such a description of $V_{r}(\mathbb{G})$.
The following definition makes explicit how the map of group algebras $\mu_{x,*}:k(x)\otimes k\mathbb{G}_{a(r)}\to k(x)\otimes k\mathbb{G}_{(r)}$ naturally determines by “restricting along $\epsilon_{r}:k[u]/u^{p}\to k\mathbb{G}_{a(r)}$” the map of $k(x)$-algebras $(\mu_{x})_{*}\circ\epsilon_{r}:k(x)[u]/u^{p}\to\ k(x)\otimes k\mathbb{G}_{(r)}$. We remind the reader that group algebra $k\mathbb{G}_{a(r)}$ (i.e., the dual of the coordinate algebra $k[\mathbb{G}_{a(r)}]=k[t]/t^{p^{r}}$) is naturally identified as the commutative $k$-algebra $k[u_{0},\ldots,u_{r-1}]/(u_{i}^{p})$ where $u_{i}$ is the linear dual to $t^{p^{i}}$.
Definition 2.3.
Let $\mathbb{G}$ be an affine group scheme of finite type over $k$. For $r\geq 1$, we denote by $\epsilon_{r}:k[u]/u^{p}\to k\mathbb{G}_{a(r)}$ the map of $k$-algebras (but not of Hopf algebras if $r>1$) which sends $u$ to the element $u_{r-1}\in k\mathbb{G}_{a(r)}$. A scheme-theoretic point $x\in V_{r}(\mathbb{G})$ corresponding to the 1-parameter subgroup $\mu_{x}:\mathbb{G}_{a(r),k(x)}\to\mathbb{G}_{(r),k(x)}$ determines the map of $k(x)$-algebras
(2.3.1) $(\mu_{x})_{*}\circ\epsilon_{r}:k(x)[u]/u^{p}\to k(x)\otimes k\mathbb{G}_{a(r)}\to k(x)\otimes k\mathbb{G}_{(r)}\quad u\mapsto(\mu_{x})_{*}(u_{r-1}).$
Thus, if $M$ is a $\mathbb{G}$-module, then $((\mu_{x})_{*}\circ\epsilon_{r})^{*}(k(x)\otimes M)$ is equipped with a $k(x)[u]/u^{p}$-module structure such that $u$ acts as $(\mu_{x})_{*}(u_{r-1})$ on $k(x)\otimes M$.
To unpack $(\mu_{x})_{*}(u_{r-1})\in k(x)\otimes k\mathbb{G}_{(r)}$, we observe that this is the $k(x)$-linear function associated the the $k$-linear function $k[\mathbb{G}]\to k$ sending $f\in k[\mathbb{G}]$ to the coefficient of $t^{p^{r-1}}$ in the expansion of $\mu_{x}^{*}(f)\in k(x)\otimes k[\mathbb{G}_{a(r)}]=k(x)[t]/t^{p^{r}}$.
Maps of the form (2.3.1) are $\pi$-points of $\mathbb{G}$; $\pi$-point equivalence classes of these maps are in natural 1-1 correspondence with the scheme theoretic points of the $\Pi$-point scheme $\Pi(\mathbb{G})$ which is isomorphic to $\operatorname{Proj}\nolimits(-)$ applied to $H^{*}(\mathbb{G},k)$ (see [12]).
In the following definition, we recall the natural grading of $V_{r}(\mathbb{G})$ as discussed in [22, §1]. This is particularly relevant because $\mathbb{P}V_{r}(\mathbb{G})\equiv\operatorname{Proj}\nolimits k[V_{r}(\mathbb{G})]$ is naturally homeomorphic to the $\pi$-point scheme $\Pi(\mathbb{G})$ (see [15, Prop 3.1]).
Definition 2.4.
We consider $A$-points of $V_{r}(\mathbb{G}_{a(r)})$ as additive polynomials in $A[t]/t^{p^{r}}\to A[t]/t^{p^{r}}$ (i.e., maps of $k$ algebras given by $t\mapsto\sum_{i}a_{i}t^{p^{i}})$ and consider the monoid action $V_{r}(\mathbb{G}_{a(r)})\times\mathbb{A}^{1}\to V_{r}(\mathbb{G}_{a(r)})$ sending $(\sum_{i=0}^{r-1}a_{i}t^{p^{i}},\ b)$ to $\sum_{i=0}^{r-1}a_{i}(bt)^{p^{i}}.$
Let $\mathbb{G}$ be an affine group scheme of finite type over $k$. The right action of $\mathbb{A}^{1}$ on $V_{r}(\mathbb{G}_{a(r)})$ induces a right action of $\mathbb{A}^{1}$ on $V_{r}(\mathbb{G})$ with associated grading on $k[V_{r}(\mathbb{G})]$. We set
$\mathbb{P}V_{r}(\mathbb{G})\quad\equiv\quad\operatorname{Proj}\nolimits k[V_{r}(\mathbb{G})].$
If $\mu:\mathbb{G}_{(r),R}\to\mathbb{G}_{R}$ is an $R$-point of $V_{r}(\mathbb{G})$ for some commutative $k$-algebra $R$ and $s$ an $R$-point of $A^{1}$ (i.e., an element $s\in R$), then $(s\circ\mu)_{*}(u_{i})=s^{p^{i}}\mu_{*}(u_{i})$ (see the proof of [23, Prop 6.5]).
The following proposition is a rephrasing of Lemma 1.12 of [22].
Proposition 2.5.
The graded commutative algebra $k[V_{r}(GLN)]$ is a graded quotient of the graded algebra $k\{X_{i,j}^{\ell};1\leq i,j,\leq N,0\leq\ell<r\}$ provided that $X_{i,j}^{\ell}$ is given degree $p^{\ell}$. Moreover, if $\mathbb{G}\subset GL_{N}$ is a closed immersion, then $k[V_{r}(\mathbb{G})]$ is a graded quotient of $k[V_{r}(GLN)]$.
[…]
To emphasize the subtlety of the geometry of $\mathbb{P}V_{r}(\mathbb{G})$, we analyze the very special example of $\mathbb{P}V_{r}(\mathbb{G}_{a})$.
Example 2.6.
We identify the graded coordinate algebra of $V_{r}(\mathbb{G}_{a})=V_{r}(\mathbb{G}_{(r)})$ as the graded polynomial algebra $k[T_{0},T_{1},\ldots,T_{r-1}]$, where $T_{i}$ is given degree $p^{i}$. Then $\operatorname{Proj}\nolimits$ applied to the graded algebra is the weighted projective space $w\mathbb{P}(1,p,\ldots,p^{r-1})$. Let $k[t_{0},t_{1},\ldots,t_{r-1}]$ denote the graded polynomial algebra with each $t_{i}$ of degree 1. The graded map of $k$-algebras
$\phi^{*}:k[T_{0},T_{1},\ldots,T_{r-1}]\ \to\ k[t_{0},t_{1},\ldots,t_{r-1}],\quad T_{i}\mapsto t_{i}^{p^{i}}$
determine a ramified covering map
$\Phi:\mathbb{P}^{r-1}\quad\to\quad w\mathbb{P}(1,p,\ldots,p^{r-1}).$
Similarly, the map of $k$-algebras
$\psi^{*}:k[t_{0},t_{1},\ldots,t_{r-1}]\ \to\ k[T_{0},T_{1},\ldots,T_{r-1}],\quad t_{i}\mapsto T_{i}^{p^{r-i-1}}$
multiplies all degrees by $p^{r-1}$ and thus induces a morphism
$\Psi:w\mathbb{P}(1,p,\ldots,p^{r-1})\quad\to\quad\mathbb{P}^{r-1}.$
Moreover,
$\Psi\circ\Phi=F^{r-1}:\mathbb{P}^{r-1}\ \to\mathbb{P}^{r-1},\quad\Phi\circ\Psi=F^{r-1}:w\mathbb{P}(1,p,\ldots,p^{r-1})\to w\mathbb{P}(1,p,\ldots,p^{r-1}).$
This will enable us to compare Jordan type functions for $(\mathbb{G}_{a}^{\times r})_{1}$ and $\mathbb{G}_{a(r)}$ in Example 5.2.
Example 2.7.
Consider the case $r=3$, so that $\mathbb{P}V_{3}(\mathbb{G}_{a})\ =\ w\mathbb{P}(1,p,p^{2})$ is 2-dimensional, covered by three affine open subsets:
$U_{0}\ =\ \operatorname{Spec}\nolimits k[T_{1}/T_{0}^{p},T_{2}/T_{0}^{p^{2}}]\ \simeq\ \mathbb{A}^{2},$
$U_{1}\ =\ \operatorname{Spec}\nolimits k[T_{0}^{p}/T_{1},T_{2}/T_{1}^{p}]\ \simeq\ \mathbb{A}^{2},$
[…]
Observe that $U_{2}\ \subset\ w\mathbb{P}(1,p,p^{2})$ has an isolated singularity at $T_{0}=0=T_{1}$; the maximal ideal at this singular point is generated by the $p+1$ elements of
$\{T_{0}^{p^{2}}/T_{2},T_{0}^{p^{2}-p}T_{1}/T_{2},\ldots,T_{0}^{p}T_{1}^{p-1}/T_{2},T_{1}^{p}/T_{2}\}$.
The representability by $k[V_{r}(\mathbb{G})]$ of the functor sending a commutative $k$-algebra $A$ to set of all height $r$ 1-parameter subgroups $\mathbb{G}_{a(r),A}\to\mathbb{G}_{A}$ (over $\operatorname{Spec}\nolimits A$) leads to the universal 1-parameter subgroup $\mathcal{U}_{\mathbb{G},r}$ as recalled in the following definition.
Definition 2.8.
([14, §2]) Let $\mathbb{G}$ be an affine group scheme of finite type over $k$ and let $r$ be a chosen positive integer. Then the universal 1-parameter subgroup over $V_{r}(\mathbb{G})$
(2.8.1) $\mathcal{U}_{\mathbb{G},r}:\mathbb{G}_{a(r),k[V_{r}(\mathbb{G})]}\ \to\mathbb{G}_{(r),k[V_{r}(\mathbb{G})]}$
of height $r$ is the 1-parameter subgroup of the group scheme $\mathbb{G}_{(r),k[V_{r}(\mathbb{G})]}$ over $V_{r}(\mathbb{G})$ corresponding to the identity map $V_{r}(\mathbb{G})\to V_{r}(\mathbb{G})$. By the universality of $\mathcal{U}_{\mathbb{G},r}$, for any scheme-theoretic point $x\in V_{r}(\mathbb{G})$ the restriction along $\chi_{x}:\operatorname{Spec}\nolimits k(x)\to V_{r}(G)$ of $\mathcal{U}_{\mathbb{G},r}$ equals the map $\mu_{x}:\mathbb{G}_{a(r),k(x)}\ \to\mathbb{G}_{(r),k(x)}$ of Definition 2.1.
We denote by
$(\mathcal{U}_{\mathbb{G},r})_{*}:k[V_{r}(\mathbb{G})]\otimes k\mathbb{G}_{a(r)}\quad\to\quad k[V_{r}(\mathbb{G})]\otimes k\mathbb{G}_{(r)}$
the map of group algebras over $k[V_{r}(\mathbb{G})]$ induced by $\mathcal{U}_{\mathbb{G},r}$. In other words, $(\mathcal{U}_{\mathbb{G},r})_{*}$ is the $k[V_{r}(\mathbb{G})]$ dual of the map on coordinate algebras $\mathcal{U}_{\mathbb{G},r}^{*}:k[V_{r}(\mathbb{G})]\otimes k[\mathbb{G}_{(r)}]\to k[V_{r}(\mathbb{G})]\otimes k[\mathbb{G}_{a(r)}]$.
We readily observe that a closed embedding $i:\mathbb{H}\hookrightarrow\mathbb{G}$ of affine group schemes of finite type over $k$ determines the commutative square
(2.8.2) ${\mathbb{G}_{a(r),k[V_{r}(\mathbb{H})]}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{(\mathcal{U}_{\mathbb{H},r})_{*}}$${\mathbb{H}_{(r),k[V_{r}(\mathbb{H})]}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathbb{G}_{a(r),k[V_{r}(\mathbb{G})]}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{(\mathcal{U}_{\mathbb{G},r})_{*}}$${\mathbb{G}_{(r),k[V_{r}(\mathbb{G})]}].}$
We recall and reformulate the “universal $p$-nilpotent operator” $\Theta_{\mathbb{G},r}$ for $\mathbb{G}_{(r)}$-modules of [14, Defn 2.1].
Definition 2.9.
Let $\mathbb{G}$ be an affine group scheme of finite type over $k$ and let $r$ be a chosen positive integer. We define
(2.9.1) $\Theta_{\mathbb{G},r}\quad\equiv\quad(\mathcal{U}_{\mathbb{G},r})_{*}(1\otimes u_{r-1})\quad\in\quad k[V_{r}(\mathbb{G})]\otimes k\mathbb{G}_{(r)}\ \equiv\ k\mathbb{G}_{(r),k[V_{r}(\mathbb{G})]},$
where $(\mathcal{U}_{\mathbb{G},r})_{*}$ is the map of “group algebras” for the map $\mathcal{U}_{\mathbb{G},r}$ of group schemes defined in Definition 2.8.
In [14, Ex 2.6], the reader will find a few explicit computations of $\Theta_{\mathbb{G},r}$. We mention two such examples consistent with the indexing used in this paper.
Example 2.10.
Consider $\mathbb{G}=\mathbb{G}_{a}^{\times r}$ whose Lie algebra is the dimension $r$ Lie algebra $\mathfrak{g}_{a}^{\oplus r}$ (which is commutative, with $p$-th power map trivial). Identify $V_{1}(\mathbb{G}_{a}^{\times r})$ with $\mathbb{A}^{r}=\operatorname{Spec}\nolimits k[t_{0},\ldots,t_{r-1}]$ and $k(\mathbb{G}_{a}^{\times r})_{!}$ with $k[x_{0},\ldots,x_{r-1}]/(x_{i}^{p})$. Then
$\Theta_{\mathbb{G}_{a}^{\times r},1}\ =\ \sum_{i=0}^{r-1}t_{i}\otimes x_{i}\ \in\ k[V_{1}(\mathbb{G}_{a}^{\times r}]\otimes\mathfrak{u}(\mathfrak{g}_{a}^{\oplus r}).$
Consider $\mathbb{G}=\mathbb{G}_{a(r)}$ whose Lie algebra is $\mathfrak{g}_{a}$. Identify $V_{r}(\mathbb{G}_{(r)})$ with $\operatorname{Spec}\nolimits k[T_{0},\ldots,T_{r-1}]$ with $T^{i}$ given grading $p^{i}$ and $k\mathbb{G}_{(r)}$ with $k[u_{0},\ldots,u_{r-1}]/(u_{i}^{p})$. Then
$\Theta_{\mathbb{G}_{a},r}\ =\ \sum_{i=0}^{r-1}T_{i}^{p^{r-1-i}}\otimes u_{i}\ +\ \cdots\quad\in\ k[V_{r}(\mathbb{G}_{a})]\otimes k\mathbb{G}_{a(r)},$
where the non-explicit, remaining terms of $\Theta_{\mathbb{G}_{a},r}$ each consist of a polynomial homogeneous of weighted degree $p^{r-1}$ in the $T_{i}$’s tensor a product of two or more $u_{i}$’s.
The somewhat subtle aspect of Theorem 2.11 is its last assertion of homogeneity of degree $p^{r-1}$ whose proof is given in [14] (reflected in the examples above).
Theorem 2.11.
([14]) Let $\mathbb{G}$ be an affine group scheme of finite type over $k$ and let $r$ be a chosen positive integer. Then $\Theta_{\mathbb{G},r}$ as given in Definition 2.9 satisfies the following properties.
  1. (1)
    For any scheme-theoretic point $x\in V_{r}(\mathbb{G})$, the specialization along $\chi_{x}:\operatorname{Spec}\nolimits k(x)\to V_{r}(\mathbb{G})$ of $\Theta_{\mathbb{G},r}\in k[V_{r}(\mathbb{G})]\otimes k\mathbb{G}_{(r)}$ equals $(\mu_{x})_{*}(1\otimes u_{r-1})\quad\in\ k(x)\otimes k\mathbb{G}_{(r)}$, where $\mu_{x}$ is the 1-parameter subgroup of Definition 2.1.
  2. (2)
    Let $\phi:\mathbb{H}\hookrightarrow\mathbb{G}$ be a closed subgroup scheme determining $\phi^{*}:k[V_{r}(\mathbb{G})]\twoheadrightarrow k[V_{r}(\mathbb{H})]$ and $\phi_{(r)*}:k\mathbb{H}_{(r)}\hookrightarrow k\mathbb{G}_{(r)}$. Then the restriction of $(\phi^{*}\otimes 1)(\Theta_{\mathbb{G},r})\in k[V_{r}(\mathbb{H})]\otimes k\mathbb{G}_{(r)}$ equals the image of $(1\otimes\phi_{(r)*})(\Theta_{\mathbb{H},r})$.
  3. (3)
    For any $\mathbb{G}_{(r)}$-module $M$, tensor product with $\Theta_{\mathbb{G},r}$ determines a $p$-nilpotent, $k[V_{r}(\mathbb{G})]$-linear operator $\Theta_{\mathbb{G},r,M}:k[V_{r}(\mathbb{G})]\otimes M\ \to\ k[V_{r}(\mathbb{G})]\otimes M$.
  4. (4)
    The base change along a scheme theoretic point $x\in V_{r}(\mathbb{G})$ of $\Theta_{\mathbb{G},r,M}$ equals the action of $(\mu_{x})_{*}(1\otimes u_{r-1})$ on $k(x)\otimes M$.
  5. (5)
    $\Theta_{\mathbb{G},r}$ is of the form $\sum f_{i}\otimes\phi_{i}\in k[V_{r}(\mathbb{G})]\otimes k\mathbb{G}_{(r)}$ with each $f_{i}$ homogeneous of degree $p^{r-1}$ in $k[V_{r}(\mathbb{G})]$ and each $\phi_{i}\in k\mathbb{G}_{(r)}$ an element of the augmentation ideal $I_{\mathbb{G}_{(r)}}$ of $k\mathbb{G}_{(r)}$.
Proof.
Since $\mu_{x}$ is the pull-back along $\chi_{x}$ of $\mathcal{U}_{\mathbb{G},r}$, the following diagram commutes:
(2.11.1) ${k[V_{r}(\mathbb{G})]\otimes k\mathbb{G}_{a(r)}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\chi^{*}}$$\scriptstyle{(\mathcal{U}_{\mathbb{G},r})_{*}}$${k[V_{r}(\mathbb{G})]\otimes k\mathbb{G}_{(r)}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\chi^{*}}$${k(x)\otimes k\mathbb{G}_{a(r)}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{(\mu_{x})_{*}}$${k(x)\otimes k\mathbb{G}_{(r)}.}$
Consequently, the restriction of $\Theta_{\mathbb{G},r}=\mathcal{U}_{\mathbb{G},r})_{*}(1\otimes u_{r-1})$ equals $(\mu_{x})_{*}(1\otimes u_{r-1})$.
The naturality of $\mathbb{G}\mapsto\Theta_{\mathbb{G},r}$ with respect to closed embeddings $\phi:\mathbb{H}\hookrightarrow\mathbb{G}$ as in assertion (2) follows from the commutativity of (2.8.2) as verified in [14, Prop 2.9].
The action by $\Theta_{\mathbb{G},r,M}$ is given by the restriction to $\{\Theta_{\mathbb{G},r}\}\otimes(k[V_{r}(\mathbb{G})]\otimes M)$ of the tensor product pairing
(2.11.2) $(k[V_{r}(\mathbb{G})]\otimes k\mathbb{G}_{(r)})\otimes(k[V_{r}(\mathbb{G})]\otimes M)\quad\to\quad k[V_{r}(\mathbb{G})]\otimes M.$
The fact that this action is $p$-nilpotent follows immediately from the observation that the $i$-th power of $\Theta_{\mathbb{G},r}$ equals $(\mathcal{U}_{\mathbb{G},r})_{*}((1\otimes u_{r-1})^{i})\ =\ ((\mathcal{U}_{\mathbb{G},r})_{*}(1\otimes u_{r-1}))^{i}$. Assertion (4) follows directly from assertion (1).
The last assertion is stated and proved in [14, Prop 2.11]. The proof proceeds by first observing that it suffices to verify that the restriction of the $k[V_{r}(\mathbb{G})]$-linear map $\mathcal{U}_{\mathbb{G},r}^{*}:k[V_{r}(\mathbb{G})]\otimes k[\mathbb{G}_{(r)}]\to k[V_{r}(\mathbb{G})]\otimes k[\mathbb{G}_{a(r)}]$ to $1\otimes k[\mathbb{G}_{(r)}]$ reads off the coefficient of $t^{p^{r-1}}$ in $k[V_{r}(\mathbb{G})]\otimes k\mathbb{G}_{a(r)}\simeq k[V_{r}(\mathbb{G})]\otimes k[t]/t^{p^{r}}$. The authors then verify that the actions of $\mathbb{A}^{1}$ on $\mathbb{G}_{a(r)}$ and $V_{r}(\mathbb{G})$ are appropriately compatible. ∎
The following is an immediate corollary of Theorem 2.11.
Corollary 2.12.
The Jordan type of the $k(x)\otimes k[u]/u^{p}$-module obtained as the restriction of the $k(x)\otimes k\mathbb{G}_{(r)}$-module $k(x)\otimes M$ along $(\mu_{x})_{*}\circ\epsilon_{r}$ (as in Definiton 2.3) equals the Jordan type of  $k(x)\otimes_{k[V_{r}(\mathbb{G})]}\Theta_{\mathbb{G},r,M}\ \in\ End_{k(x)}(k(x)\otimes M)$.
For any graded $R$-module $M_{*}$, we denote by $(M_{*})^{\sim}$ the associated quasi-coherent sheaf on $\operatorname{Proj}\nolimits(R)$. We shall often consider graded $R$-modules of the form $R\otimes V$ where $V$ is a $k$-vector space and $R\otimes V$ inherits the grading of $R$.
If $M_{*}$ is a graded $R$-module, we denote by $M_{*}(n)$ the graded module obtained from $M_{*}$ by “shifting down by $n$” the degrees of $M_{*}$; thus, $(M_{*}(n))_{d}=M_{n+d}$. We use $(M_{*})^{\sim}(n)$ to denote $(M_{*}(n))^{\sim}$. The operator $\Theta_{\mathbb{G},r,M}$ of Theorem 2.11(3) multiplies degrees by $p^{r-1}$, so that viewed as an operator on graded modules one should write $\Theta_{\mathbb{G},r,M}$ as
$\Theta_{\mathbb{G},r,M}:k[V_{r}(\mathbb{G})]\otimes M\quad\to\quad(k[V_{r}(\mathbb{G})]\otimes M)(p^{r-1}).$
Because $k$ has positive characteristic, the informative survey by M. Beltrametti and L. Robbiano in [2] does not apply to our consideration of $\operatorname{Proj}\nolimits V_{r}(\mathbb{G})$. Because $V_{r}(\mathbb{G})$ is usually not generated by homogeneous elements of degree 1, we supplement the clear exposition of R. Hartshorne in [16, §II.5] with [8, Ex 9.5] when considering the relationship between graded modules over a graded, commutative $k$-algebra $R$ and sheaves of modules on $\operatorname{Proj}\nolimits(R)$. We also point out that our weighted projective spaces are not obtained by taking a quotient by an action of $\mathbb{G}_{m}$.
Proposition 2.13.
Consider a polynomial algebra $k[T_{0},\ldots,T_{N}]$ with each $T_{i}$ given a “weighted” degree $d_{i}$. Assume that $I_{Y}\subset k[T_{0},\ldots,T_{N}]$ is an ideal generated by polynomials $F_{j}(\underline{T})$ which are (weighted) homogenous. Then, the zero locus $Y=Z(I_{Y})\ \subset\operatorname{Proj}\nolimits k[T_{0},\ldots,T_{N}]$ is a weighted projective variety associated to the graded algebra $k[T_{0},\ldots,T_{N}]/I_{Y}$.
Let $d$ be a positive integer divisible by each $d_{i}$. The graded $k[T_{0},\ldots,T_{N}]/I_{Y}$-module $(k[T_{0},\ldots,T_{N}]/I_{Y})(d)$ is a locally free, coherent rank 1 sheaf of $\mathcal{O}_{Y}$-modules on $Y$. We denote this sheaf by $\mathcal{O}_{Y}(d)$. For each $i$ such that $T_{i}$ is not in the radical of $I_{Y}$, the restriction $(\mathcal{O}_{Y})_{|V_{i}}$ of $\mathcal{O}_{Y}$ to  $V_{i}\ \equiv(\operatorname{Proj}\nolimits k[T_{0},\ldots,T_{N}])-Z(T_{i}))\cap Y$ is generated by the image of $T_{i}^{d/d_{i}}$ viewed as a global section of $\mathcal{O}_{Y}(d)$.
Proof.
It suffices to assume that $I_{Y}$ is a radical ideal . Moreover, by considering each irreducible component of $Y$, we may assume that $Y$ is irreducible. Then [8, Ex 9.5] applies to verify that the variey $Z(I_{Y})$ is isomorphic to its image in $\operatorname{Proj}\nolimits S$, where $S\ =k[T_{0},\ldots,T_{N}]_{(d)}\subset k[T_{0},\ldots,T_{N}]$ is the graded subalgebra generated by (weighted) homogeneous polynomials each of whose weights is divisible by $d$ and replacing $I_{Y}$ by the homogenous ideal in $S$ defining the image of $Y$ in $\operatorname{Proj}\nolimits S$. Thus, replacing $k[T_{0},\ldots,T_{N}]$ by $S$ with grading modified by dividing degrees by $d$, we may apply [16, Prop II.5.12] (which is stated for graded submodules of a graded commutative algebra generated in degree 1). ∎
In particular, Proposition 2.13 applies to the graded commutative algebra $k[V_{r}(\mathbb{G})]$ which is generated by elements homogeneous of degree $p^{i},1\leq i<r$. (See Proposition 2.5.)
We next present the projectivization of the universal operator $\Theta_{\mathbb{G},r}$. This operator has been used in [14] to produce global invariants of $\mathbb{G}$-modules (especially, non-trivial vector bundles on $\mathbb{P}\mathbb{G}_{(r)}$) in contrast to the Jordan type function $JT_{\mathbb{G},r,M}(-)$ which is a “pointwise on $\mathbb{P}\mathbb{G}_{(r)}$” invariant of $M$.
Theorem 2.14.
The universal operator $\Theta_{\mathbb{G},r}$ of Theorem 2.11 determines a global section
$\tilde{\Theta}_{\mathbb{G},r}\quad\in\quad\Gamma(\mathcal{O}_{\mathbb{P}V_{r}(\mathbb{G})}(p^{r-1})\otimes k\mathbb{G}_{(r)})$
of the coherent, locally free sheaf $\mathcal{O}_{\mathbb{P}V_{r}(\mathbb{G})}(p^{r-1})\otimes k\mathbb{G}_{(r)})$ on $\mathbb{P}V_{r}(\mathbb{G})$ satisfying the following properties.
  1. (1)
    For any $\mathbb{G}_{(r)}$-module $M$, tensor product with $\tilde{\Theta}_{\mathbb{G},r}$ determines the map of $\mathcal{O}_{\mathbb{P}V_{r}(\mathbb{G})}$-modules
    $\tilde{\Theta}_{\mathbb{G},r,M}:(k[V_{r}(\mathbb{G})]\otimes M)^{\sim}\quad\to\quad(k[V_{r}(\mathbb{G})]\otimes M)^{\sim}(p^{r-1}).$
  2. (2)
    For any non-zero divisior $h\in k[V_{r}(\mathbb{G})]_{p^{r-1}}$, the natural isomorphism of $(\mathcal{O}_{\mathbb{P}V_{r}(\mathbb{G})})_{U_{h}}$-modules
    $h^{-1}\otimes 1:(\mathcal{O}_{\mathbb{P}V_{r}(\mathbb{G})}(p^{r-1})\otimes k\mathbb{G}_{(r)})_{|U_{h}}\quad\stackrel{{\scriptstyle\sim}}{{\to}}\quad(\mathcal{O}_{\mathbb{P}V_{r}(\mathbb{G})}\otimes k\mathbb{G}_{(r)})_{|U_{h}}$
    sends
    $(\tilde{\Theta}_{\mathbb{G},r})_{|U_{h}}\ \in\ \Gamma(U_{h},\mathcal{O}_{\mathbb{P}V_{r}(\mathbb{G})}(p^{r-1})\otimes k\mathbb{G}_{(r)})\ =\ (k[V_{r}(\mathbb{G})][1/h])_{p^{r-1}}\otimes k\mathbb{G}_{(r)}$
    to
    $(h^{-1}\otimes 1)\cdot\Theta_{\mathbb{G},r}\in k[V_{r}(\mathbb{G})][1/h]\otimes k\mathbb{G}_{(r)}.$
  3. (3)
    If $\overline{x}$ is a scheme-theoretic point of $\mathbb{P}V_{r}(\mathbb{G})$ satisfying the condition that $h(\overline{x})\not=0$ and if $x$ is a scheme theoretic point of $V_{r}(\mathbb{G})$ projecting to $\overline{x}$, then the restriction $k(x)\otimes_{\mathcal{O}_{\mathbb{P}V_{r}(\mathbb{G})}}\tilde{\Theta}_{\mathbb{G},r}$ along $\chi_{\overline{x}}:\operatorname{Spec}\nolimits k(x)\to\mathbb{P}V_{r}(\mathbb{G})$ of $\tilde{\Theta}_{\mathbb{G},r}$ is given by
    (2.14.1) $k(x)\otimes_{\mathcal{O}_{\mathbb{P}V_{r}(\mathbb{G})}}\tilde{\Theta}_{\mathbb{G},r}\quad=\quad c\cdot(\mu_{x})_{*}(1\otimes u_{r-1})\in k(x)\otimes k\mathbb{G}_{(r)}$
    for some $0\not=c\in k(x)$.
  4. (4)
    For any $\mathbb{G}_{(r)}$-module $M$, the restriction of $\tilde{\Theta}_{\mathbb{G},r,M}$ along $\chi_{\overline{x}}:\operatorname{Spec}\nolimits k(x)\to\mathbb{P}V_{r}(\mathbb{G})$ is a non-zero scalar multiple of the image of $(\mu_{x})_{*}(1\otimes u_{r-1})$ in
    $End_{k(x)}(k(x)\otimes M$).
Proof.
The global sections of the sheaf $\mathcal{O}_{\mathbb{P}V_{r}(\mathbb{G})}(p^{r-1})\otimes k\mathbb{G}_{(r)}$ on $\mathbb{P}V_{r}(\mathbb{G})$ are those elements of degree 0 of the graded $k[V_{r}(\mathbb{G})]$-module $k[V_{r}(\mathbb{G})](p^{r-1})\otimes k\mathbb{G}_{(r)}$. Since $\Theta_{\mathbb{G},r}\in k[V_{r}(\mathbb{G})]\otimes k\mathbb{G}_{(r)}$ is homogeneous of degree $p^{r-1}$, it determines the global section $\Theta_{\mathbb{G},r}$ of $\mathcal{O}_{\mathbb{P}V_{r}(\mathbb{G})}(p^{r-1})\otimes k\mathbb{G}_{(r)}$.
The first assertion of the theorem follows from Theorem 2.11(3) and the correspondence between graded $k[V_{r}(\mathbb{G})]$-modules and quasi-coherent sheaves on $\mathbb{P}V_{r}(\mathbb{G})$. The second assertion follows from the standard description of principal open subsets of $\operatorname{Proj}\nolimits R$ for a graded ring $R$ in terms of the degree 0 part of $R[h^{-1}]$ for homogenous elements $h\in R$.
The equality (2.14.1) follows from Theorem 2.11(1) and the identification of assertion (2). Assertion (4) follows immediately from assertion (3) (see Theorem 2.11(4)). ∎
Definition 2.15.
We denote by $Hom_{cont}(\mathbb{P}V_{r}(\mathbb{G}),\mathcal{Y}_{\leq,m})$ the set of continuous maps from $\mathbb{P}V_{r}(\mathbb{G})$ with the Zariski topology to $\mathcal{Y}_{\leq,m}$ with the Alexandrov discrete topology (as in Theorem 1.11). For $F,G\in Hom_{cont}(\mathbb{P}V_{r}(\mathbb{G}),\mathcal{Y}_{\leq,m})$, we write $F\leq G$ if and only if $F(\overline{x})\leq G(\overline{x})\in\mathcal{Y}_{\leq,m}$ for all scheme-theoretic points $\overline{x}\in\mathbb{P}V_{r}(\mathbb{G})$.
The Jordan type function in the following theorem was first formulated in [13, Defn 2.3] and its continuity briefly specified in [13, Prop 2.8.3].
Theorem 2.16.
Let $\mathbb{G}$ be an affine group scheme of finite type over $k$ and let $r$ be a positive integer. If $M$ is a $\mathbb{G}_{(r)}$-module of dimension $m$ and if $\chi_{x}:\operatorname{Spec}\nolimits k(x)\to V_{r}(\mathbb{G})$ is a lift of a scheme-theoretic point $\chi_{\overline{x}}:\operatorname{Spec}\nolimits k(\overline{x})\to\mathbb{P}V_{r}(\mathbb{G})$, we define
$JT_{\mathbb{G},r,M}(\overline{x})\quad\equiv\quad JT((\mu_{x})_{*}\circ\epsilon_{r})^{*}(k[V_{r}(\mathbb{G})]\otimes M))$
to be the Jordan type of the $k(x)[u]/u^{p}$-module obtained by restricting $k[V_{r}(\mathbb{G})]\otimes M$ along $(\mu_{x})_{*}\circ\epsilon_{r}:k(x)[u]/u^{p}\to k(x)\otimes k\mathbb{G}_{a(r)}\to k[V_{r}(\mathbb{G})]\otimes k\mathbb{G}_{(r)}$ (equivalent to the Jordan type of $(\mu_{x})_{*}(1\otimes u_{r-1})$ acting on $k(x)\otimes M$).
  1. (1)
    The association $\overline{x}\mapsto JT_{\mathbb{G},r,M}(\overline{x})$ determines a continuous map
    $JT_{\mathbb{G},r,M}(-):\mathbb{P}V_{r}(\mathbb{G})\quad\to\quad\mathcal{Y}_{\leq,m};$
    in other words, $JT_{\mathbb{G}_{(r)},M}(-)\ \in\ Hom_{cont}(\mathbb{P}V_{r}(\mathbb{G}),\mathcal{Y}_{\leq,m})$.
  2. (2)
    For any $\overline{x}\in\mathbb{P}V_{r}(\mathbb{G})$, $M\in mod(\mathbb{G}_{(r)},m)$, and $N\in mod(\mathbb{G}_{(r)},n)$,
    $JT_{\mathbb{G},r,M\oplus N}(\overline{x})\ =\ JT_{\mathbb{G},r,M}(\overline{x})+JT_{\mathbb{G},r,N}(\overline{x})\ \in\ \mathcal{Y}_{\leq,m+n}.$
  3. (3)
    If $r=1$ or if $\overline{x}\in\mathbb{P}V_{r}(\mathbb{G})$ is a generic point, if $M\in mod(\mathbb{G}_{(r)},m)$, and if $N\in mod(\mathbb{G}_{(r)},n)$, then
    $JT_{\mathbb{G},r,M\otimes N}(\overline{x})\ =\ JT_{\mathbb{G},r,M}(\overline{x})\otimes JT_{\mathbb{G},r,N}(\overline{x})\ \in\ \mathcal{Y}_{\leq,m\cdot n}.$
  4. (4)
    If $f:M\to N$ is a surjective map of $\mathbb{G}_{(r)}$-modules, then $JT_{\mathbb{G},r,N}(\overline{x})\ \leq\ JT_{\mathbb{G},r,M}(\overline{x})$ for any $\overline{x}\in\mathbb{P}V_{r}(\mathbb{G})$; in other words, $JT_{\mathbb{G},r,N}(-)\leq JT_{\mathbb{G},r,M}(-)$, where the partial ordering of functions is that inherited by the partial ordering of $\mathcal{Y}$.
  5. (5)
    Consider a short exact sequence $\mathcal{E}:0\to N\ \to\ M\ \to\ Q\to 0$ of finite dimensional $\mathbb{G}_{(r)}$-modules. Then $\mathcal{E}$ is locally split if and only if $JT_{\mathbb{G},r,M}(-)=JT_{\mathbb{G},r,N}(-)+JT_{\mathbb{G},r,Q}(-).$
  6. (6)
    If $\phi:\mathbb{H}\to\mathbb{G}$ is a morphism of affine group schemes of finite type over $k$, then
    $JT_{\mathbb{H},r,\phi^{*}(M)}(-):\mathbb{P}V_{r}(\mathbb{H})\ \to\ \mathcal{Y}_{\leq,m}$
    equals
    $JT_{\mathbb{G},r,M}(-)\circ\phi_{*}:\mathbb{P}V_{r}(\mathbb{H})\ \to\ \mathbb{P}V_{r}(\mathbb{G})\ \to\ \mathcal{Y}_{\leq,m}.$
Proof.
The fact that $JT_{\mathbb{G},r,M}(\overline{x})$ is well defined (namely, independent of the choice of $x\in V_{r}(\mathbb{G})$ projecting to $\overline{x}\in\mathbb{P}V_{r}(\mathbb{G})$), follows from the equality (2.14.1) of Theorem 2.14.
To prove the asserted continuity of assertion (1) it suffices to prove continuity of $JT_{\mathbb{G},r,M}(-)$ when restricted to each affine open subset of the form $\operatorname{Spec}\nolimits(k[V_{r}(\mathbb{G})][1/h])_{0}\\ \subset\mathbb{P}V_{r}(\mathbb{G})$ where $h\in k[V_{r}(\mathbb{G})]_{p^{r-1}}$ is a power of some homogeneous generator of $k[V_{r}(\mathbb{G})]$. This follows from the continuity proved in Theorem 1.11 for $A=(k[V_{r}(\mathbb{G})][1/h])_{0}$.
The fact that $JT_{\mathbb{G},r,M\oplus N}(-)=JT_{\mathbb{G},r,M}(-)+JT_{\mathbb{G},r,N}(-)$ is self evident.
Since $(\mu_{x})_{*}:k(x)\otimes k\mathbb{G}_{a(r)}\to k(x)\otimes k\mathbb{G}_{(r)}$ is a map of Hopf algebras (over $k(x)$), the restriction map $\mu_{x}^{*}:mod(k(x)\otimes k\mathbb{G}_{(r)})\to mod(k(x)\otimes k\mathbb{G}_{a(r)})$ commutes with tensor products. Since $\epsilon_{1}$ is the identity, this implies assertion (3) whenever $r=1$. On the other hand, if $x$ is a generic point of $V_{r}(\mathbb{G})$ for any $r>0$, the fact that $M\ \mapsto\ JT_{\mathbb{G},r,M}(x)$ commutes with tensor products is given by [15, Prop4.7].
A surjective map $f:M\to N$ of finite dimensional $\mathbb{G}_{(r)}$-modules induces a surjective map $k(x)\otimes M\ \to\ k(x)\otimes N$ of $k(x)[u]/t^{p}$-modules. Consequently, assertion (4) follows from Corollary 1.4. Assertion (5) follows from the definition of a locally split exact sequence (see [4]) and the fact for any field $K$ and any short exact sequence $0\to N_{K}\to M_{K}\to Q_{K}\to 0$ that the Jordan type of $M_{K}$ equals the sum of the Jordan types of $N_{K}$ and $Q_{K}$ if and only if the short exact sequence splits,
Assertion (6) is equivalent to the assertion that $JT_{\mathbb{G},r,M}(\phi_{*}(y))=JT_{\mathbb{H},r,\phi^{*}(M)}(y)$ for every $y\in\mathbb{P}V_{r}(\mathbb{H})$. This is immediate from the identification (with $x=\phi_{*}(y)$)
$JT((\mu_{x})_{*}\circ\epsilon_{r})^{*}(k(x)\otimes M)\quad=\quad JT((\mu_{y})_{*}\circ\epsilon_{r})^{*}(k(y)\otimes\phi^{*}(M)).$
Remark 2.17.
For any $\mathbb{G}_{(r)}$-module $M$ of dimension $m$, $JT_{\mathbb{G},r,M}(-)=JT_{\mathbb{G},r,M^{*}}(-):\mathbb{P}V_{r}(\mathbb{G})\to\mathcal{Y}_{\leq,m}$ where $M^{*}$ is the $k$-linear dual of $M$. Moreover, if $I$ and $J$ restrict to injective $\mathbb{G}_{(r)}$ modules of the same dimension, then $JT_{\mathbb{G},r,I}(-)=JT_{\mathbb{G},r,J}(-)$. Thus, $M\mapsto JT_{\mathbb{G},r,M}(-)$ is far from faithful.
Proposition 2.18.
Consider a linear algebraic group $\mathbb{G}$ and denote by $\mathbb{P}V_{r}(\mathbb{G})(k)/\mathbb{G}(k)$ the quotient of the set of $k$-points of $\mathbb{P}V_{r}(\mathbb{G})$ by the adjoint action of the discrete group $\mathbb{G}(k)$. For any $\mathbb{G}$-module $M$ of dimension $m$, the functor $JT_{\mathbb{G},r,M}(-)$ of Theorem 2.16(1) induces a set map
$JT_{\mathbb{G},r,M}(-):\mathbb{P}V_{r}(\mathbb{G})(k)/\mathbb{G}(k)\quad\to\quad\mathcal{Y}_{\leq,m}.$
Proof.
Consider two $k$-points $x,y\in V_{r}(\mathbb{G})$ and $g\in\mathbb{G}(k)$ with $y=g\circ x$ for the adjoint action of $\mathbb{G}(k)$ on $V_{r}(\mathbb{G})$. Let $M^{g}$ denote the $\mathbb{G}$-module with the same underlying $k$-vector space structure as that of $M$ and with $\mathbb{G}$-action as that given by the $\mathbb{G}$-action on $M$ preceded by conjugation by $g$. Observe that action by $g$ determines an isomorphism $g:M\to M^{g}$.
We have a commutative diagram
(2.18.1) ${k[V_{r}(\mathbb{G})]\otimes M^{g}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{g}$$\scriptstyle{\Theta_{\mathbb{G},r,M^{g}}}$${k[V_{r}(\mathbb{G})]\otimes M\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\Theta_{\mathbb{G},r,M}}$${k[V_{r}(\mathbb{G})]\otimes M^{g}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{g}$$\scriptstyle{\chi_{y}^{*}}$${k[V_{r}(\mathbb{G})]\otimes M\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\chi_{x}^{*}}$${k(x)\otimes M^{g}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{g}$${k(y)\otimes M.}$
This implies that the action of $(\mu_{y})_{*}(u_{r-1})=(\chi_{y}^{*}\circ\Theta_{\mathbb{G},r})_{*}(1\otimes u_{r-1})$ on $M^{g}$ is isomorphic to the action of $(\mu_{x})_{*}(u_{r-1})=(\chi_{x}^{*}\circ\Theta_{\mathbb{G},r})_{*}(1\otimes u_{r-1})$ on $M$. ∎
The following theorem associates to a finite dimensional $\mathbb{G}$-module $M$ the configuration $\{(\mathbb{P}V_{r}(\mathbb{G}))_{M,\leq\underline{a}}\}$ which gives considerably more information about $M$ than does the support variety $\mathbb{P}V_{r}(\mathbb{G})_{M}\subset\mathbb{P}V_{r}(\mathbb{G})$ of $M$. The fact that $(V_{r}(\mathbb{G}))_{M,\leq\underline{a}}\subset V_{r}(\mathbb{G})$ is closed is given with a sketch of proof in [13, Prop 2.8.3].
Theorem 2.19.
(see Proposition 2.8.3 of [13]) Let $\mathbb{G}$ be an affine group of finite type over $k$ and let $r$ be a positive integer. Consider a $\mathbb{G}_{(r)}$-module $M$ of dimension $m$. For each $\underline{a}\in\mathcal{Y}_{\leq,m}$, we define the subset
$(\mathbb{P}V_{r}(\mathbb{G}))_{M,\leq\underline{a}}\ \equiv\ \{\overline{x}:JT_{\mathbb{G},r,M}(\overline{x})\leq\underline{a}\}\ \subset\ \mathbb{P}V_{r}(\mathbb{G}).$
Then $\{(\mathbb{P}V_{r}(\mathbb{G}))_{M,\leq\underline{a}};\ \underline{a}\in\mathcal{Y}_{\leq,m}\}$ is a partially ordered configuration of closed subspaces of $\mathbb{P}V_{r}(\mathbb{G})$.
With this notation, the support variety $\mathbb{P}V_{r}(\mathbb{G})_{M}$ of [23] equals the union of those subspaces $(\mathbb{P}V_{r}(\mathbb{G}))_{M,\underline{a}\leq}$ indexed by $\underline{a}\in\mathcal{Y}_{\leq,m}$ which satisfy the condition that $a_{i}\not=0$ for some $i<p$.
The non-maximal support variety $\Gamma(\mathbb{G}_{(r)})_{M}$ of [13] equals the union of those subspaces $(\mathbb{P}V_{r}(\mathbb{G}))_{M,\underline{a}\leq}$ indexed by the (finitely many) $\underline{a}\in\mathcal{Y}_{\leq,m}$ which satisfy the condition that there exists some $\underline{b}\in\mathcal{Y}_{\leq,m}$ with $\underline{a}\leq\underline{b}$ such that $(\mathbb{P}V_{r}(\mathbb{G}))_{M,\leq\underline{a}}$ is strictly contained in $(\mathbb{P}V_{r}(\mathbb{G}))_{M,\leq\underline{b}}$
Proof.
To prove that $(\mathbb{P}V_{r}(\mathbb{G}))_{M,\leq\underline{a}}$ is closed in $\mathbb{P}V_{r}(\mathbb{G})$, it suffices to verify the affine analogue: $V_{r}(\mathbb{G})_{M,\leq\underline{a}}$ is closed in $V_{r}(\mathbb{G})$, where $V_{r}(\mathbb{G})_{M,\leq\underline{a}}\equiv\{x:JT_{\mathbb{G}_{(r)},M}(x)\leq\underline{a}\}$. This follows from Theorem 1.11 with $A=k[V_{r}(\mathbb{G}]$, $M_{A}=A\otimes M$, and $\theta_{A}$ equal to $\Theta_{\mathbb{G},r}$ (of Theorem 2.11).
The identification of the subspaces $\mathbb{P}V_{r}(\mathbb{G})_{M}=\cup_{\underline{a}\not=n[p]}(\mathbb{P}V_{r}(\mathbb{G})_{M,\leq\underline{a}})$ and $\Gamma(\mathbb{G}_{(r)})_{M}$ (as asserted above) inside $\mathbb{P}V_{r}(\mathbb{G})$ follows immediately from the definitions of these subspaces. ∎
We make explicit the following corollary first stated and proved in [13, Cor 2.15]. The configuration $\{(\mathbb{P}V_{r}(\mathbb{G}))_{M,\leq\underline{a}};\ \underline{a}\in\mathcal{Y}_{\leq,m}\}$ gives a geometric “picture” of the array of Jordan types of a $\mathbb{G}$-module of dimension $m$. One may view this configuration as giving a geometric interpretation of $JT_{\mathbb{G},r,M}(-)\in Hom_{cont}(\mathbb{P}V_{r}(\mathbb{G}),\mathcal{Y})$.
Corollary 2.20.
Let $M$ be a $\mathbb{G}_{(r)}$-module of dimension $m$, and let $\underline{a}$ be an object on $\mathcal{Y}_{\leq,m}$, The subset
$(\mathbb{P}V_{r}(\mathbb{G}))_{M,=\underline{a}}\ \equiv\ \{x:JT_{\mathbb{G},r,M}(x)=\underline{a}\}\ \subset\ \mathbb{P}V_{r}(\mathbb{G})$
is locally closed. More precisely, $(\mathbb{P}V_{r}(\mathbb{G}))_{M,=\underline{a}}$ is the complement inside $(\mathbb{P}V_{r}(\mathbb{G}))_{M,\leq\underline{a}}$ of the union of those $(\mathbb{P}V_{r}(\mathbb{G}))_{M,\leq\underline{b}}$ such that $\underline{b}\leq\underline{a}$ and $\underline{b}\not=\underline{a}$,
$(\mathbb{P}V_{r}(\mathbb{G}))_{M,=\underline{a}}\quad=\quad(\mathbb{P}V_{r}(\mathbb{G}))_{M,\leq\underline{a}}-(\cup_{\underline{a}\not=\underline{b}\leq\underline{a}}(\mathbb{P}V_{r}(\mathbb{G}))_{M,\leq\underline{b}}).$
Remark 2.21.
In [13, Defn 4.1], generalized support varieties $\Gamma^{j}(\mathbb{G})_{M}$ were introduced for finite dimensional $\mathbb{G}$-modules $M$ for an arbitrary group scheme $\mathbb{G}$. For $\mathbb{G}$ infinitesimal, one can recover $\Gamma^{j}(\mathbb{G})_{M}$ as in Theorem 2.19 by replacing $\Theta_{\mathbb{G},r,M}$ in (2.9.1) in Theorem 2.11 by its $j$-th power, then replacing $JT((\mu_{x})_{*}\circ\epsilon_{r})^{*}(k[V_{r}(\mathbb{G})]\otimes M)$ in Theorem 2.16 by the Jordan type of the $j$-th power of $(\mu_{x})_{*}(u_{r-1})$ acting on $k(x)\otimes M$.
3. First examples
In this section, we consider a few general examples of computations of $JT_{\mathbb{G},r,M}(-)$ (some found in [13, §2]). More specific computations for the simpler $JT^{exp}_{\mathbb{G},r,M}(-)$ are given in Section 5. For the examples below, we consider $\mathbb{G}$ to be an infinitesimal group scheme of height $\leq r$ for a specific $r>0$, modifying our notation in these case by using $JT_{\mathbb{G},M}(-)$ to denote $JT_{\mathbb{G},r,M}(-)$.
Example 3.1.
Let $\mathbb{G}$ be an arbitrary infinitesimal group scheme of height $\leq r$. A $\mathbb{G}$-module $M$ of dimension $m$ is of constant Jordan type $\underline{a}$ if and only if $JT_{\mathbb{G},M}(-)$ is the constant function sending each $\overline{x}\in\mathbb{P}V_{r}(\mathbb{G})$ to $\underline{a}\in\mathcal{Y}_{\leq,m}$. (See [5] for numerous examples.)
In particular, if $I$ is an injective $\mathbb{G}$-module of dimension $m$, then $p$ divides $m$ and $JT_{\mathbb{G},I}(-)$ sends each $\overline{x}\in\mathbb{P}V_{r}(\mathbb{G})$ to $\frac{m}{p}[p]\in\mathcal{Y}_{\leq,m}$ (i.e., $JT_{\mathbb{G},I}(\overline{x})=\frac{m}{p}[p]+0[p-1]+\cdots+0[1]$).
Example 3.2.
Let $\mathbb{G}$ be an arbitrary infinitesimal group scheme of height $\leq r$ and $\overline{x}\in\mathbb{P}V_{r}(\mathbb{G})$. Consider $0\not=\zeta\in H^{2d}(\mathbb{G},k)$ and take $M$ to be the Carlson $L_{\zeta}$-module as first constructed in [3]; namely, $M$ is the kernel of a map $\tilde{\zeta}:\Omega^{2d}(k)\to k$ representing $\zeta$.
Then $JT_{\mathbb{G},M}(\overline{x})=\frac{n}{p}[p]$ if $((\mu_{x})_{*}\circ\epsilon_{r})^{*}(\zeta))\not=0\in H^{2d}(k(x)[u]/u^{p},k)$, and
$J_{\mathbb{G},M}(\overline{x})=(\frac{n}{p}-1)[p]+[p-1]+[1]$ if $((\mu_{x})_{*}\circ\epsilon_{r})^{*}(\zeta))=0$.
Example 3.3.
(See [13, Prop 2.9]) Once again, let $\mathbb{G}$ be an arbitrary infinitesimal group scheme of height $\leq r$. Let $M$ be a finite dimensional $\mathbb{G}$-module and $P\twoheadrightarrow M$ a finite dimensional projective cover of $M$ with kernel $\Omega(M)$. If $JT_{\mathbb{G},M}(\overline{x})\ =\underline{a}$ for some $\overline{x}\in\mathbb{P}V_{r}(\mathbb{G})$, then $JT_{\mathbb{G},\Omega(M)}(\overline{x})\ =\underline{a}^{\perp}+n[p]$ for some integer $n$, where $\underline{a}^{\perp}=\sum_{i=1}^{p-1}a_{p-i}[i]$.
Example 3.4.
(See [17, §I.3.8].) Consider $\mathbb{G}_{1},\ \mathbb{G}_{2}$, infinitesimal group schemes of height $\leq r$. Let $\mathbb{G}=\mathbb{G}_{1}\times\mathbb{G}_{2}$ and $M_{1}$ be a a finite dimensional $\mathbb{G}_{1}$-module. Set $M$ equal to $ind_{\mathbb{G}_{1}}^{\mathbb{G}}M_{1}\simeq M_{1}\otimes k[\mathbb{G}_{2}]$. Let $\overline{(x,y)}$ be a scheme theoretic point in $\mathbb{P}V_{r}(\mathbb{G})$ represented by the scheme-theoretic point $(x,y)\in V_{r}(\mathbb{G}_{1})\times V_{r}(\mathbb{G}_{2})$. Then
$JT_{\mathbb{G},M}(\overline{(x,0}))\ =\ dim(k[\mathbb{G}_{2}])\cdot JT_{\mathbb{G}_{1},M_{1}}((\overline{x})),$
$JT_{\mathbb{G},M}(\overline{(x,y}))\ =\ dim(M)/p\cdot[p],\quad y\not=0.$
Example 3.5.
Consider the special case $r=1$. Then applying the functor sending a group scheme to its Lie algebra determines an equivalence between the category of finite group schemes of height 1 and the category of finite dimensional restricted Lie algebra; this sends $\mathbb{G}$ to $\mathfrak{g}$, with $k\mathbb{G}$ identified with the restricted enveloping algebra $\mathfrak{u}(\mathfrak{g})$ of $\mathfrak{g}$. Consequently, a 1-parameter subgroup $\mathbb{G}_{a(1)}\otimes A\to\mathbb{G}_{A}$ corresponds to a choice of element $\alpha\in A\otimes\mathfrak{g}$ with $\alpha^{[p]}=0$. Thus, $V_{1}(\mathbb{G})$ is identified with the $p$-nilpotent cone $\mathcal{N}_{p}(\mathfrak{g})\subset\mathfrak{g}$.
Let $M$ be a faithful representation of $\mathbb{G}$ of dimension $N$, and associate to $M$ an embedding $\phi:\mathbb{G}\hookrightarrow GL_{N(1)}\hookrightarrow GL_{N}$ with the property that $M=\phi^{*}(V_{N})$ (where $V_{N}$ is the defining representation for $GL_{N}$). For any $\alpha\in\mathcal{N}_{p}(\mathfrak{g})(A)$, the associated 1-parameter subgroup of $\mathbb{G}$ sends a $p$-nilpotent element $s\in A$ to the $N\times N$ matrix $1+s\cdot\alpha+\frac{s^{2}}{2}\alpha^{2}+\cdots+\frac{s^{p-1}}{(p-1)!}\alpha^{p-1}\in GL_{N}(A)$ (which is in the image of $\mathbb{G}(A)\to GL_{N}(A)$). Moreover, if the scheme-theoretic point $x\in\mathcal{N}_{p}(\mathfrak{g})$ corresponds to $\alpha\in k(x)\otimes\mathfrak{g}$, then $JT_{\mathbb{G},M}(x)$ is the Jordan type of $\alpha\in\mathcal{N}_{p}(\mathfrak{g})(A)\subset M_{N,N}(A)$.
In particular, if $M=\mathfrak{g}l_{N}$, the adjoint representation of $GL_{N}$, then $JT_{GL_{N},1,\mathfrak{g}l_{N}}(-)$ is the class function taking a $p$-nilpotent element of $\mathfrak{g}l_{N}$ to its Jordan block structure viewed as a Young diagram.
Example 3.6.
Assume $\mathbb{G}=\mathbb{G}_{a(r)}$ for some $r>0$. Then $\mathbb{G}$ admits an embedding in $GL_{2}$ (as the $r$-th Frobenius kernel of the unipotent subgroup $U_{2}$ of strictly upper triangular matrices). As in Example 2.6, $\mathbb{P}V_{r}(\mathbb{G})$ is a weighted projective space with weights $(1,p,\cdots,p^{r-1})$ associated to the graded polynomial algebra $k[T_{0},T_{1},\ldots,T_{r-1}]$ where $T_{i}$ has weight $p^{i}$. A scheme-theoretic point $\underline{x}\in\mathbb{P}V_{r}(\mathbb{G})$ lifts to a scheme-theoretic point $x\in V_{r}(X)$ represented by a non-zero $r$-tuple $\underline{a}=(a_{0},\ldots,a_{r-1})$ of elements in $k(x)$ with associated 1-parameter subgroup $\mu_{x}:\mathbb{G}_{a(r),k(x)}\to\mathbb{G}_{a(r),k(x)}\simeq k(x)[u_{0},\ldots,u_{r-1}]/(u_{i}^{p})$. An explicit (but quite complicated) formula for $\mu_{x*}(u_{r-1})$ for $x$ represented by $(a_{0},\ldots,a_{r-1})$ is given in [23, (6.5.12)].
The sum of terms linear in $u_{i}$ in the expression for $\mu_{x*}(u_{r-1})$ is $\sum_{i=0}^{r-1}a_{r-1-i}^{p^{i}}u_{i}\in k(x)\otimes k\mathbb{G}$. See Example 2.10.
4. Exponential Jordan type
In this section, we consider affine group schemes whose 1-parameter subgroups are given by exponentiating $r$-tuples of $p$-nilpotent, pair-wise commuting elements of $\mathfrak{g}\equiv Lie(\mathbb{G})$. The more explicit nature of these 1-parameter subgroups allow us to simplify our Jordan type functors, not necessarily preserving their values but preserving their close relationship with support varieties. The key to this simplification is a fundamental computation of [15, Thm 1.12] which tells us that only linear terms of commuting generators of group actions given as polynomials in these generators are needed to determine support varieties.
Definition 4.1.
Let $\mathbb{G}$ be an affine group scheme of finite type over $k$ with Lie algebra $\mathfrak{g}$. The functor on commutative $k$ algebras sending $A$ to an $r$-tuple $\underline{B}=(B_{0},\ldots,B_{r-1})$ of elements in $A\otimes\mathfrak{g}$ which satisfy the conditions that $B_{s}^{[p]}=0$ for all $i,0\leq s<r$ and that $[B_{s},B_{s^{\prime}}]=0$ for all $s,s^{\prime}$ with $0\leq s,s^{\prime}<r$ is represented by an affine scheme $\mathcal{C}_{r}(\mathfrak{g})$. In contrast to [11, 1.11.2], we give $\mathcal{C}_{r}(\mathfrak{g})$ the monoid action $\mathbb{A}^{1}\times\mathcal{C}_{r}(\mathfrak{g})\to\mathcal{C}_{r}(\mathfrak{g})$ by sending $(\alpha;B_{0},\ldots,B_{r-1})$ to $(\alpha\cdot B_{0},\alpha^{p}\cdot B_{1},\ldots,\alpha^{p^{r-1}}\cdot B_{r-1})$.
For $\mathbb{G},\mathcal{E}_{(r)}$ an affine group scheme of exponential type of height $r$ as in Definitions 4.2 and 4.6, the equality $(\mathcal{U}_{\mathfrak{g},r})_{*}(u_{r-1})\ =(\Phi_{\mathfrak{g},r}\otimes 1)^{*}(\Theta_{\mathbb{G},r})$ of Proposition 4.5(2) below gives an identification of the universal $p$-nilpotent operator $\Theta_{\mathbb{G},r}\in k[V_{r}(\mathbb{G})]\otimes\mathbb{G}_{(r)}$ utilizing $\mathcal{U}_{\mathfrak{g},r}$ of (4.3.2) in place of the universal height $r$ 1-parameter subgroup $\mathcal{U}_{\mathbb{G},r}$ This identification has the advantage that $\mathcal{U}_{\mathfrak{g},r}$ is given in terms of the “explicit” representation of 1-parameter subgroups of $\mathbb{G}_{(r)}$ in terms of $r$-tuples of $p$-nilpotent, pairwise elements of $\mathfrak{g}$.
Definition 4.2.
Let $\mathbb{G}$ be an affine group scheme of finite type over $k$ with Lie algebra $\mathfrak{g}$. Fix a positive integer $r$. Assume that $\mathbb{G}$ is equipped with a morphism
$\mathcal{E}_{(r)}:\mathbb{G}_{a(r)}\times\mathcal{N}_{p}(\mathfrak{g})\quad\to\quad\mathbb{G}_{(r)}$
which satisfies the following conditions for any commutative $k$-algebra $A$, any $p$-nilpotent $B\in A\otimes\mathfrak{g}$, and any $\alpha,\beta\in\mathbb{G}_{a(r)}(A)$:
  1. (1)
    The specialization at $B$, $\mathcal{E}_{(r),B}:\mathbb{G}_{a(r),A}\to\mathbb{G}_{(r),A}$, is a 1-parameter subgroup of $\mathbb{G}_{(r),A}$.
  2. (2)
    If $B^{\prime}\in A\otimes\mathfrak{g}$ commutes with $B$, then
    $\mathcal{E}_{(r),B}(\alpha)\cdot\mathcal{E}_{(r),B^{\prime}}(\beta)=\mathcal{E}_{(r),B^{\prime}}(\beta)\cdot\mathcal{E}_{(r),B}(\alpha)\in\mathbb{G}_{(r)}(A).$
  3. (3)
    $\mathcal{E}_{(r),\alpha\cdot B}(-)\ =\ \mathcal{E}_{(r),B}(\alpha\cdot-):\mathbb{G}_{a(r),A}\to\mathbb{G}_{(r),A}$.
Definition 4.3.
Let $\mathbb{G}$ be an affine group scheme of finite type over $k$ equipped with $\mathcal{E}_{(r)}:\mathbb{G}_{a(r)}\times\mathcal{N}_{p}(\mathfrak{g})\to\mathbb{G}_{(r)}$ satisfying the conditions of Definition 4.2. For any $s,0\leq s<r$, we define
(4.3.1) $\mathcal{E}_{r|s}\quad\equiv\quad\mathcal{E}_{(r),B}\circ(1\times p_{s})\circ(F^{s}\times 1)$
to be the composition
$\mathbb{G}_{a(r)}\times\mathcal{C}_{r}(\mathfrak{g})\ \to\ \mathbb{G}_{a(r)}\times\mathcal{C}_{r}(\mathfrak{g})\to\mathbb{G}_{a(r)}\times\mathcal{N}_{p}(\mathfrak{g})\to\mathbb{G}_{(r)}$
where $p_{s}$ is projection onto the $s$-th factor and $F^{s}$ is the $s$-th iterate of the Frobenius morphism.
We define
(4.3.2) $\mathcal{E}_{\mathfrak{g},r}\quad\equiv\quad\prod\circ\ (\times_{s=0}^{r-1}\mathcal{E}_{r|s})\circ Diag$
to be the composition
$\mathbb{G}_{a(r)}\times\mathcal{C}_{r}(\mathfrak{g})\ \to(\mathbb{G}_{a(r)}\times\mathcal{C}_{r}(\mathfrak{g}))^{\times r}\ \to\ (\mathbb{G}_{(r)})^{\times r}\ \to\mathbb{G}_{(r)}$
where the product is taken with respect to the group structure of $\mathbb{G}_{(r)}$.
We define
(4.3.3) $\mathcal{U}_{\mathfrak{g},r|s}\ \equiv\ \mathcal{E}_{r|s}\times pr_{2},\quad\mathcal{U}_{\mathfrak{g},r}\ \equiv\ \mathcal{E}_{\mathfrak{g},r}\times pr_{2},$
the “scalar extensions” of $\mathcal{E}_{r|s},\ \mathcal{E}_{(r)}$ along $k\to k[\mathcal{C}_{r}(\mathfrak{g})]$.
Remark 4.4.
Observe that for any scheme-theoretic point $\underline{B}=(B_{0},\ldots,B_{r-1})$, $(\mathcal{E}_{r|s})_{*}(u_{r-1})$ determines
(4.4.1) $t^{p^{r-1}}\in\mathbb{G}_{a(r)}(K[t]/t^{p^{r}})\ \mapsto\ t^{p^{r-s-1}}\in\mathbb{G}_{a(r)}(K[t]/t^{p^{r}})\ \mapsto(\mathcal{E}_{(r),B_{s}})_{*}(t^{r-s-1})\in\mathbb{G}_{(r)}(K[t]/t^{p^{r}})$
where $K=k(\underline{B})$ and we identify $t^{p^{r-1}}\in\mathbb{G}_{a(r)}(k[t]/t^{p^{r}})=k[t]/t^{p^{r}}$ with $u_{r-1}\in(k[t]/t^{p^{r}})^{*}$. (See [22, Eqn 1.3.1] with $S=k[t]/t^{p^{r}}$.)
Thus, $(\mathcal{E}_{\mathfrak{g},r})_{*}(u_{r-1})$ determines
$t^{p^{r-1}}\quad\mapsto\quad\prod_{s=0}^{r-1}(\mathcal{E}_{(r),B_{s}})_{*}(t^{r-s-1})\ \in\ \mathbb{G}_{(r)}(K[t]/t^{p^{r}}).$
The following proposition “specializes” the structures of Definition 2.8 for an affine group scheme $\mathbb{G}$ satisfying the condition that it is equipped with $\mathcal{E}{(r)}:\mathbb{G}_{a(r)}\times\mathcal{N}(\mathfrak{g})\to\mathbb{G}_{(r)}$ as in Definition 4.2.
Proposition 4.5.
Let $\mathbb{G}$ and $\mathcal{E}_{(r)}$ satisfy the conditions of Definition 4.2.
  1. (1)
    The morphisms
    $\mathcal{U}_{\mathfrak{g},r|s},\ \mathcal{U}_{\mathfrak{g},r}:\mathbb{G}_{a(r),k[\mathcal{C}_{r}(\mathfrak{g})]}\quad\to\quad\mathbb{G}_{(r),k[\mathcal{C}_{r}(\mathfrak{g})]}$
    are 1-parameter subgroups.
  2. (2)
    These 1-parameter subgroups determine morphisms
    $\Phi_{\mathfrak{g},r|s},\ \Phi_{\mathfrak{g},r}:\mathcal{C}_{r}(\mathfrak{g})\quad\to\quad V_{r}(\mathbb{G})$
    with the property that their pull-backs of the universal 1-parameter group  $\mathcal{U}_{\mathbb{G},r}:\mathbb{G}_{a(r),k[V_{r}(\mathbb{G})]}\to\ \mathbb{G}_{(r),k[V_{r}(\mathbb{G})]}$ equal  $\mathcal{U}_{\mathfrak{g},r|s},\ \mathcal{U}_{\mathfrak{g},r}$.
  3. (3)
    $\Phi_{\mathfrak{g},r}^{*}(\mathcal{U}_{\mathbb{G},r})\ =\ \mathcal{U}_{\mathfrak{g},r}$, where $\mathcal{U}_{\mathbb{G},r}$ is given in Definition 2.9.
  4. (4)
    Each $\Phi_{\mathfrak{g},r|s}$ commutes with the monoid actions of $\mathbb{A}^{1}$ on $\mathcal{C}_{r}(\mathfrak{g})$ and $V_{r}(\mathbb{G})$ in the sense that the follow squares commute:
[…]
Proof.
Over any point $\underline{B}$ of $\mathcal{C}_{r}(\mathfrak{g})$, $\mathcal{U}_{\mathfrak{g},r|s}and\ \mathcal{U}_{\mathfrak{g},r}$ are morphisms of group schemes over $\underline{B}$; this implies assertion (1). Assertions (2) and (3) follow from the universal property of $\mathcal{U}_{\mathbb{G},r}:\mathbb{G}_{a(r),k[V_{r}(\mathbb{G})]}\quad\to\quad\mathbb{G}_{(r),k[V_{r}(\mathbb{G})]}$.
To check the commutativity of (4.5.1), it suffices to observe for any $\alpha\in A$ and for any $A$-point $\mathcal{C}_{r}(\mathbb{G})$ of the form $(0,\ldots,B_{s},\ldots 0)$ (and thus of weight $p^{s}$) that $\mathcal{E}_{(r),\alpha\cdot\underline{B}}(t)\ =\mathcal{E}_{(r),(0,\ldots,\alpha^{p^{s}}B_{s},\ldots 0)}(t)=\mathcal{E}_{(r),\underline{B}}(\alpha^{p^{s}}t)=(\alpha\cdot\mathcal{E}_{(r),\underline{B}})(t)$. ∎
In [22], the authors introduce the condition that an embedding $\mathbb{G}\hookrightarrow GL_{N}$ of linear algebraic groups is of exponential type; they observe that many classical groups satisfy this condition, including many classical simple groups, their parabolic subgroups, and unipotent radicals of these parabolic subgroups. Subsequent generalizations of this condition on $\mathbb{G}$ occur in [10] and [11]. A more definitive account of such linear algebraic groups can be found in P. Sobaje’s article [21]. The following definition is general enough to include examples of those papers.
Definition 4.6.
Let $\mathbb{G}$ be an affine group scheme of finite type over $k$ with Lie algebra $\mathfrak{g}$. Assume that $\mathbb{G}$ is equipped with a morphism
$\mathcal{E}_{(r)}:\mathbb{G}_{a(r)}\times\mathcal{N}_{p}(\mathfrak{g})\quad\to\quad\mathbb{G}_{(r)}$
satisfying the conditions of Definition 4.2. We say that $\mathbb{G}$ equipped with $\mathcal{E}_{(r)}$ is of exponential type of height $r$ if the morphism $\Phi_{\mathfrak{g},r}:\mathcal{C}_{r}(\mathfrak{g})\ \to\ V_{r}(\mathbb{G})$ of Theorem 4.5(1) induces a bijection on geometric points.
In [9, Defn 1.6], the notions of a “continuous algebraic map” and a “bicontinuous morphism” were introduced in order to study rational maps between schemes which induce well defined maps on geometric points. In that work, the topologies of interest were either the ètale topology or the analytic topology (in the case that the ground field was the complex numbers).
Lemma 4.7.
Let $(\mathbb{G},\mathcal{E}_{(r)})$ be affine group scheme of exponential type of height $r$. The bicontinuous morphism
(4.7.1) $\Phi_{\mathfrak{g},r}:\mathcal{C}_{r}(\mathfrak{g})\quad\to\quad V_{r}(\mathbb{G}).$
is a homeomorphism of Zariski topological spaces which commutes with $\mathbb{A}^{1}$-actions.
Proof.
The assumption of exponential type of height $r$ implies that $\Phi$ is a bijection on scheme-theoretic points. Since $\Phi_{\mathfrak{g},r}$ is a morphism, it is continuous. Since $\mathbb{P}\Phi_{\mathfrak{g},r}:\mathbb{P}\mathcal{C}_{r}(\mathfrak{g})\to\mathbb{P}V_{r}(\mathbb{G})$ is proper, it is a closed map, thus inducing a bijection on closed subsets, thus a homeomorphism. This implies that the pull-back/restriction $\mathcal{C}_{r}(\mathfrak{g})-\{0\}\to V_{r}(\mathbb{G})-\{0\}$ is a homeomorphism. Since the cone point 00 is closed in both $\mathcal{C}_{r}(\mathfrak{g})$ and $V_{r}(\mathbb{G})$, this implies that $\Phi_{\mathfrak{g},r}$ is also a homeomorphism. ∎
As in (2.9.1) of Theorem 2.11, we consider $p$-nilpotent operators associated to the 1-parameter subgroups  $\mathcal{U}_{\mathfrak{g},r|s},\ \mathcal{U}_{\mathfrak{g},r}$ of Proposition 4.5(2).
Theorem 4.8.
Let $(\mathbb{G},\mathcal{E}_{(r)})$ be affine group scheme of exponential type of height $r$. We define the $p$-nilpotent operators
(4.8.1) $\Theta_{\mathfrak{g},r|s}\ \equiv\ (\mathcal{U}_{\mathfrak{g},r|s})_{*}(1\otimes u_{r-1}),\ \Theta_{\mathfrak{g},r}\ \equiv\ (\mathcal{U}_{\mathfrak{g},r})_{*}(1\otimes u_{r-1})\ \in\ k[\mathcal{C}_{r}(\mathfrak{g})]\otimes\mathbb{G}_{(r)},$
and denote by $\Theta_{\mathfrak{g},r|s,\underline{B}},\ \Theta_{\mathfrak{g},r,\underline{B}}$ their specializations at a scheme-theoretic point $\underline{B}\in\mathcal{C}_{r}(\mathfrak{g})$.
  1. (1)
    $\Theta_{\mathfrak{g},r}\ =\ (\Phi^{*}_{\mathfrak{g},r}\otimes 1)(\Theta_{\mathbb{G},r})$.
  2. (2)
    $\Theta_{\mathfrak{g},r|s,\underline{B}}\ =\ (\mathcal{E}_{(r),B_{s}})_{*}(u_{r-s-1}),\ \Theta_{\mathfrak{g},r,\underline{B}}\ =\prod_{s=0}^{r-1}\Theta_{\mathfrak{g},r|s,\underline{B}}\quad\in\ k(\underline{B})\otimes\mathbb{G}_{(r)}$
  3. (3)
    Both $\Theta_{\mathfrak{g},r|s}^{exp}$ and $\Theta_{\mathfrak{g},r}^{exp}$ are homogeneous of degree $p^{r-1}$ with respect to the grading on $k[\mathcal{C}_{r}(\mathfrak{g})]$.
  4. (4)
    For any $\mathbb{G}_{(r)}$-module $M$, tensor product with $\Theta_{\mathfrak{g},r|s}$ and $\Theta_{\mathfrak{g},r}$ determines $p$-nilpotent, $k[\mathcal{C}_{r}(\mathfrak{g})]$-linear operators
    $\Theta_{\mathfrak{g},r|s,M},\ \Theta_{\mathfrak{g},r,M}:k[\mathcal{C}_{r}(\mathfrak{g})]\otimes M\ \to\ k[\mathcal{C}_{r}(\mathfrak{g})]\otimes M.$
  5. (5)
    The base changes along a scheme theoretic point $\underline{B}\in\mathcal{C}_{r}(\mathfrak{g})$ of the operators $\Theta_{\mathfrak{g},r,|sM},\ \Theta_{\mathfrak{g},r,M}$ are given by the actions of $\Theta_{\mathfrak{g},r,|s\underline{B}},\ \Theta_{\mathfrak{g},r,\underline{B}}$ on $k(\underline{B})\otimes M$.
Proof.
Assertion (1) follows from Proposition 4.5(1) and the formula of (4.4.1). Assertion (2) follows from Proposition 4.5(3). Assertions (3) and (4) and their proofs are direct analogues of Theorem 2.11 (3) and (4). ∎
Remark 4.9.
The reader might wish to compare $\Theta_{g,r}$ of (4.8.1), which is an element in the group algebra of $\mathbb{G}_{(r),k[\mathcal{C}_{r}(\mathfrak{g})]}$, with the image of $t^{p^{r}}$ as given in (4.4.1) which is a $k(\underline{B})[t]/t^{p^{r}}$-point of $\mathbb{G}_{(r)}$.
The following proposition is a mild generalization of a result of Sobaje [20, Prop 2.3] and is closely related to [10, Prop 4.2]. This proposition implies that $\Theta^{exp}_{\mathfrak{g},r}$ (much simpler than $\Theta_{\mathbb{G},r}$) determines the support theory $M\mapsto(\mathbb{P}V_{r}(\mathbb{G}))_{M}$ for finite dimensional $\mathbb{G}_{(r)}$-modules. See Corollary 4.13 below.
Proposition 4.10.
Let $(\mathbb{G},\mathcal{E}_{r})$ be an affine group scheme of exponential type of height $r$, let $M$ be a finite dimensional $\mathbb{G}_{(r)}$-module of dimension $m$, and let $\underline{B}$ be a scheme-theoretic point of $\mathcal{C}_{r}(\mathfrak{g})$. Consider
$\Theta^{exp}_{\mathfrak{g},r}\ \equiv\ \sum_{s=0}^{r-1}\Theta_{\mathfrak{g},r|s}\quad\in\quad k[\mathcal{C}_{r}(\mathfrak{g})]\otimes\mathbb{G}_{(r)},$
$\Theta^{exp}_{\mathfrak{g},r,\underline{B}}\ \equiv\ \sum_{s=0}^{r-1}\Theta_{\mathfrak{g},r|s,\underline{B}}\ =\ \sum_{s=0}^{r-1}(\mathcal{E}_{(r),B_{s}})_{*}(u_{r-s-1})\quad\in\quad k(\underline{B})\otimes\mathbb{G}_{(r)}.$
For any scheme-theoretic point $\underline{B}\in\mathcal{C}_{r}(\mathbb{G})$,
$JT(\Theta^{exp}_{\mathfrak{g},r,\underline{B}},k(\underline{B})\otimes M)\ =\ \frac{m}{p}[p]\quad\iff\quad JT(\Theta_{\mathfrak{g},r,\underline{B}},k(\underline{B})\otimes M)\ =\ \frac{m}{p}[p].$
Proof.
We repeat Sobaje’s proof of [20, Prop 2.3] to show that $\Theta_{\mathfrak{g},r,\underline{B}}-\Theta^{exp}_{\mathfrak{g},r,\underline{B}}$ is a linear combination of terms each of which is a product of two or more factors which are $p$-nilpotent and thus have $p-1$-st power 0. The proposition then follows from [6, Prop 8] (see also the more general [15, Thm 1.12]).
We may write $\Theta_{\mathfrak{g},r,\underline{B}}\ =\prod_{s=0}^{r-1}\Theta_{\mathfrak{g},r|s,\underline{B}}$ as the composition
(×$(\times_{s=0}^{r-1}\ u_{r-s-1})\circ(\otimes_{s=0}^{r-1}(\mathcal{E}_{(r),B_{s}})_{*})\circ(1\otimes\Delta):\ k(\underline{B})\otimes k[\mathbb{G}_{(r)}]\ \to\ k(\underline{B})\otimes k[\mathbb{G}_{(r)}]^{\otimes r}$
$\quad\to\ k(\underline{B})\otimes k[\mathbb{G}_{a(r)}]^{\otimes r}\to k(\underline{B}).$
Sobaje shows that the result equals $\Theta^{exp}_{\mathfrak{g},r,\underline{B}}\ \equiv\ \sum_{s=0}^{r-1}\Theta_{\mathfrak{g},r|s,\underline{B}}$ plus other terms which are scalar multiples of products of at least two factors of the form $\Theta^{exp}_{\mathfrak{g},r,|s,\underline{B}}$, factors which are $p$-nilpotent and pair-wise commuting. ∎
Remark 4.11.
By definition, $\Theta^{exp}_{\mathfrak{g},1}\ =\ \Theta_{\mathfrak{g},1}$. As observed in [20, Prop 2.3] (see also the proof of [23, Prop 6.5]), $\Theta^{exp}_{\mathfrak{g},2}\ =\ \Theta_{\mathfrak{g},2}$. On the other hand, for $r>2$, $\Theta^{exp}_{\mathfrak{g},r}\ \not=\ \Theta_{\mathfrak{g},r}$.
Remark 4.12.
The Jordan type $JT(\Theta^{exp}_{\mathfrak{g},r,\underline{B}},k(\underline{B})\otimes M)$ is confusingly related to the Jordan type for a rational $\mathbb{G}$-module with $\mathbb{G}$ a linear algebraic group (as defined in [10, Defn:4.4]). For linear algebraic groups, the Jordan type is not defined in terms of the push-forward to $k\mathbb{G}$ along 1-parameter groups $\mathbb{G}_{a}\to\mathbb{G}$ of elements of $k\mathbb{G}_{a}$. A consequence of this is that the expression of $\Theta_{\mathfrak{g},r}^{exp}$ as $\sum_{s=0}^{r-1}(\mathcal{E}_{B_{s}})_{*}(u_{r-s-1})$ in Proposition 4.10 is replaced by $\sum_{s\geq 0}(\mathcal{E}_{B_{s}})_{*}(u_{s})$ for linear algebraic groups.
Proposition 4.10 has the following immediate corollary.
Corollary 4.13.
Let $(\mathbb{G},\mathcal{E}_{r})$ be an affine group scheme of exponential type of height $r$. For any $\mathbb{G}$-module $M$ of dimension $m$, the (affine) support variety $V_{r}(\mathbb{G})_{M}\ \subset V_{r}(G)$ is the image under the bicontinuous map $\Phi_{\mathfrak{g},r}$ of
$\{\langle\underline{B}\rangle\in\mathcal{C}_{r}(\mathfrak{g}):JT(\Theta^{exp}_{\mathfrak{g},r,\underline{B}},k(\underline{B})\otimes M)\ \not=\ \frac{m}{p}[p]\}\quad\subset\quad\mathcal{C}_{r}(\mathfrak{g}).$
The proof of Proposition 4.10 (together with Theorem 1.12 of [15]) also proves the following corollary.
Corollary 4.14.
Assume the notation and hypotheses of Proposition 4.10. Assume that $\underline{B}\in\mathcal{C}_{r}(\mathfrak{g})$ is a generic point. Then for any finite dimensional $\mathbb{G}_{(r)}$-module,
$JT(\Theta^{exp}_{\mathfrak{g},r,\underline{B}},k(\underline{B})\otimes M)\quad=\quad JT(\Theta_{\mathfrak{g},r,\underline{B}},k(\underline{B})\otimes M).$
We next formulate the projectivization $\tilde{\Theta}_{\mathfrak{g},r}^{exp}$ of $\Theta_{\mathfrak{g},r}^{exp}$ and give the analogue of Theorem 2.14 for $\tilde{\Theta}^{exp}_{\mathfrak{g},r}$. We denote by $\mathbb{P}\mathcal{C}_{r}(\mathfrak{g})$ the weighted projective variety $\operatorname{Proj}\nolimits k[\mathcal{C}_{r}(\mathfrak{g})]$, parallel to the notation of $\mathbb{P}V_{r}(\mathbb{G})$ for $\operatorname{Proj}\nolimits k[V_{r}(\mathbb{G})]$. As in Theorem 2.14, we retain the notation and conventions introduced prior to Proposition 2.13 relating graded modules for a graded commutative ring $R$ and sheaves on $\operatorname{Proj}\nolimits R$.
Proposition 4.15.
Let $(\mathbb{G},\mathcal{E}_{(r)})$ be an affine group scheme of exponential type of height $r$. The operator $\Theta_{\mathfrak{g},r}^{exp}$ of Proposition 4.10 determines a global section
$\tilde{\Theta}_{\mathfrak{g},r}^{exp}\quad\in\quad\Gamma(\mathcal{O}_{\mathbb{P}\mathcal{C}_{r}(\mathfrak{g})},\mathcal{O}_{\mathbb{P}\mathcal{C}_{r}(\mathfrak{g})}(p^{r-1})\otimes k\mathbb{G}_{(r)})$
of the coherent, locally free sheaf $\mathcal{O}_{\mathbb{P}\mathcal{C}_{r}(\mathfrak{g})}(p^{r-1})\otimes k\mathbb{G}_{(r)}$ of $\mathcal{O}_{\mathbb{P}\mathcal{C}_{r}(\mathfrak{g})}$-modules which satisfies the following properties.
  1. (1)
    For any $\mathbb{G}_{(r)}$-module $M$, tensor product with $\tilde{\Theta}_{\mathfrak{g},r}^{exp}$ determines the map of $\mathcal{O}_{\mathbb{P}\mathcal{C}_{r}(\mathfrak{g})}$-modules
    $\tilde{\Theta}_{\mathfrak{g},r,M}^{exp}:(k[\mathcal{C}_{r}(\mathfrak{g})]\otimes M)^{\sim}\quad\to\quad(k[\mathcal{C}_{r}(\mathfrak{g})]\otimes M)^{\sim}(p^{r-1}).$
  2. (2)
    For any non-zero divisior $h\in k[\mathcal{C}_{r}(\mathfrak{g})]_{p^{r-1}}$, the natural isomorphism of $(\mathcal{O}_{\mathbb{P}\mathcal{C}_{r}(\mathfrak{g})})_{U_{h}}$-modules
    $h^{-1}\otimes 1:(\mathcal{O}_{\mathbb{P}\mathcal{C}_{r}(\mathbb{G})}(p^{r-1})\otimes k\mathbb{G}_{(r)})_{|U_{h}}\quad\stackrel{{\scriptstyle\sim}}{{\to}}\quad(\mathcal{O}_{\mathbb{P}\mathcal{C}_{r}(\mathbb{G})}\otimes k\mathbb{G}_{(r)})_{|U_{h}}$
    sends
    $(\tilde{\Theta}_{\mathfrak{g},r}^{exp})_{|U_{h}}\ \in\ \Gamma(U_{h},\mathcal{O}_{\mathbb{P}\mathcal{C}_{r}(\mathbb{G})}(p^{r-1})\otimes k\mathbb{G}_{(r)})\ =\ (k[\mathcal{C}_{r}(\mathbb{G})][1/h])_{p^{r-1}}\otimes k\mathbb{G}_{(r)}$
    to
    $(h^{-1}\otimes 1)\cdot\Theta_{\mathfrak{g},r}^{exp}\ \in\ k[\mathcal{C}_{r}(\mathbb{G})][1/h]\otimes k\mathbb{G}_{(r)}.$
  3. (3)
    If $\langle\underline{B}\rangle$ is a scheme-theoretic point of $\mathbb{P}\mathcal{C}_{r}(\mathbb{G})$ satisfying the condition that $h(\langle\underline{B}\rangle)\not=0$ and if $\underline{B}$ is a scheme theoretic point of $\mathcal{C}_{r}(\mathbb{G})$ projecting to $\langle\underline{B}\rangle$ with residue field $k(\langle\underline{B}\rangle)\subset k(\underline{B})$, then the restriction along $\operatorname{Spec}\nolimits k(\underline{B})\to\mathbb{P}\mathcal{C}_{r}(\mathbb{G})$ of $\tilde{\Theta}_{\mathfrak{g},r}^{exp}$ is given by
    (4.15.1) $k(\underline{B})\otimes_{\mathcal{O}_{\mathbb{P}\mathcal{C}_{r}(\mathbb{G})}}\tilde{\Theta}_{\mathfrak{g},r}^{exp}\quad=\quad c\cdot\sum_{s=0}^{r-1}(\mathcal{E}_{(r),B_{s}})_{*}(u_{r-s-1})\ \in\ k(\underline{B})\otimes k\mathbb{G}_{(r)}$
    for some $0\not=c\in k(\underline{B})$.
  4. (4)
    For any $\mathbb{G}_{(r)}$-module $M$, the restriction of $\tilde{\Theta}_{\mathfrak{g},r,M}^{exp}$ along $\operatorname{Spec}\nolimits k(\underline{B})\to\mathbb{P}\mathcal{C}_{r}(\mathfrak{g})$ as in (3) is a non-zero scalar multiple of the image of
    $\sum_{s=0}^{r-1}(\mathcal{E}_{(r),B_{s}})_{*}(u_{r-s-1})$ in $End_{k(\underline{B})}(k(\underline{B})\otimes M)$.
Proof.
By Proposition 2.13, $\mathcal{O}_{\mathbb{P}\mathcal{C}_{r}(\mathfrak{g})}(p^{r-1})\otimes k\mathbb{G}_{(r)}$ is a coherent, locally free sheaf of $\mathcal{O}_{\mathbb{P}\mathcal{C}_{r}(\mathfrak{g})}$-modules on $\mathbb{P}\mathcal{C}_{r}$. Proposition 4.5(3) implies that $(\Phi_{r|s})^{*}:k[V_{r}(\mathbb{G})]\to k[\mathcal{C}_{r}(\mathfrak{g})]$ is a map of graded $k$-algebras so that $\Theta_{\mathfrak{g},r}^{exp}$ is homogenous by Theorem 2.11(5) and Theorem 4.8(1). Thus, $\Theta_{\mathfrak{g},r}^{exp}$ determines $\tilde{\Theta}_{\mathfrak{g},r,M}^{exp}$ in
$\Gamma(\mathcal{O}_{\mathbb{P}\mathcal{C}_{r}(\mathbb{G})},\mathcal{O}_{\mathbb{P}\mathcal{C}_{r}(\mathbb{G})}(p^{r-1})\otimes k\mathbb{G}_{(r)})$.
The proofs of assertions (1) - (4) are verified by evident analogues for $\tilde{\Theta}^{exp}_{\mathfrak{g},r,}$ of the proofs given in the proof of Theorem 2.14 for the corresponding statements for $\tilde{\Theta}_{\mathbb{G},r,}$
The following is our simplification of Theorem 2.16 for $(\mathbb{G},\mathcal{E}_{r})$ $(\mathbb{G},\mathcal{E}_{r})$ is an affine group scheme of exponential type of height $r$. The proof of Theorem 4.16 follows from Proposition 4.15 exactly as the proof of Theorem 2.16 follows from Theorem 2.14.
Theorem 4.16.
Let $(\mathbb{G},\mathcal{E}_{(r)})$ be an affine group scheme of exponential type of height $r$. For any finite dimensional $\mathbb{G}_{(r)}$-module $M$ and any scheme-theoretic point $\underline{B}$ of $\mathcal{C}_{r}(\mathfrak{g})$ projecting to $\langle\underline{B}\rangle\in\mathbb{P}\mathcal{C}_{r}(\mathfrak{g})$, we define
$JT_{\mathfrak{g},r,M}^{exp}(\langle\underline{B}\rangle)\quad\equiv\quad JT(\sum_{s=0}^{p-1}(\mathcal{E}_{(r),B_{s}})_{*}(u_{r-s-1}),k(\underline{B})\otimes M),$
the Jordan type of the $p$-nilpotent element  $k(\underline{B})\otimes_{\mathcal{O}_{\mathbb{P}\mathcal{C}_{r}(\mathbb{G})}}\tilde{\Theta}_{\mathfrak{g},r}^{exp}\in k(\underline{B})\otimes\mathbb{G}_{(r)}$ acting on $k(\underline{B})\otimes M$.
  1. (1)
    For any $\mathbb{G}_{(r)}$-module $M$ of dimension $m$, the association $\langle\underline{B}\rangle\ \mapsto JT_{\mathfrak{g},r,M}^{exp}(\underline{B})$ determines a continuous map
    $JT_{\mathfrak{g},r,M}^{exp}(-):\mathbb{P}\mathcal{C}_{r}(\mathfrak{g})\quad\to\quad\mathcal{Y}_{\leq,m}.$
  2. (2)
    For $\mathbb{G}_{(r)}$-modules $M,N$ of dimensions $m,n$ and any scheme-theoretic point $\underline{B}\in\mathbb{P}\mathcal{C}_{r}(\mathfrak{g})$,
    $JT_{\mathfrak{g},r,M\oplus N}^{exp}(\langle\underline{B}\rangle)\ =\ JT_{\mathfrak{g},r,M}^{exp}(\langle\underline{B}\rangle)+JT_{\mathfrak{g},r,N}^{exp}(\langle\underline{B}\rangle)\ \in\ \mathcal{Y}_{\leq,m+n}.$
  3. (3)
    Assume either that $r=1$ or that $\langle\underline{B}\rangle\in\mathbb{P}\mathcal{C}_{r}(\mathfrak{g})$ is a generic point. For $\mathbb{G}_{(r)}$-modules $M,N$ of dimensions $m,n$,
    $JT_{\mathfrak{g},r,M\otimes N}^{exp}(\langle\underline{B}\rangle)\ =\ JT_{\mathfrak{g},r,M}^{exp}(\langle\underline{B}\rangle)\otimes JT_{\mathfrak{g},r,N}^{exp}(\langle\underline{B}\rangle)\ \in\ \mathcal{Y}_{\leq,mn}.$
  4. (4)
    If $f:M\to N$ is a surjective map of $\mathbb{G}_{(r)}$-modules, then $JT_{\mathfrak{g},r,N}^{exp}(\langle\underline{B}\rangle)\ \leq JT_{\mathfrak{g},r,M}^{exp}(\langle\underline{B}\rangle)$ for any $\langle\underline{B}\rangle\in\mathbb{P}\mathcal{C}_{r}(\mathfrak{g})$; in other words, $JT_{\mathfrak{g},r,N}^{exp}(-)\leq JT_{\mathfrak{g},r,M}^{exp}(-)$.
  5. (5)
    A short exact sequence $0\to N\ \to\ M\ \to\ Q\to 0$ of finite dimensional $\mathbb{G}_{(r)}$-modules is locally split if and only if $JT_{\mathfrak{g},r,M}^{exp}(-)=JT_{\mathfrak{g},r,N}^{exp}(-)+JT_{\mathfrak{g},r,Q}^{exp}(-).$
  6. (6)
    If $\phi:\mathbb{H}\to\mathbb{G}$ is a morphism of affine group schemes of exponential type of height $r$, then for any $\mathbb{G}_{(r)}$-module $M$ of dimension $m$
    $JT_{\mathfrak{h},r,\phi*(M)}^{exp}(-):\mathbb{P}\mathcal{C}_{r}(\mathfrak{h})\ \to\ \mathcal{Y}_{\leq,m}$
    equals
    $JT_{\mathfrak{g},r,M}^{exp}(-)\circ\phi_{*}:\mathbb{P}\mathcal{C}_{r}(\mathfrak{h})\ \to\ \mathbb{P}\mathcal{C}_{r}(\mathfrak{g})\ \to\ \mathcal{Y}_{\leq,m}.$
5. Examples of exponential Jordan types
We provide some computations of exponential Jordan types using the simplified Jordan type function $JT^{exp}_{\mathfrak{g},r,M}(-)$ of Theorem 4.16. One challenge to computing Jordan types is the behavior of Jordan types with respect to the sum of two commuting, $p$-nilpotent operators. Another challenge is the fact that for $r>1$ the Jordan type functions $JT_{\mathbb{G},r,M}(-),\ JT_{\mathfrak{g},r,M}^{exp}(-)$ is unlikely to commute with tensor products of $\mathbb{G}_{(r)}$-modules at all points of $\mathbb{P}V_{r}(\mathbb{G}),\ \mathbb{P}\mathcal{C}_{r}(\mathfrak{g})$.
Proposition 5.1.
By Proposition 4.10, Examples 3.1 and 3.2 remain valid if $(\mathbb{G},\mathcal{E}_{(r)})$ is an affine group scheme of exponential type of height $r$ and if $JT_{\mathbb{G},r,M}(-)$ is replaced by $JT^{exp}_{\mathfrak{g},r,M}(-)$. The justification of Example 3.3 also justifies its validity when $JT_{\mathbb{G},r,M}(-)$ is replaced by $JT^{exp}_{\mathfrak{g},r,M}(-)$.
Example 3.5 remains basically unchanged, since the Jordan type of any $p$-nilpotent $\alpha\in\mathfrak{g}_{K}$ is the Jordan type of $exp(\alpha)$.
On the other hand, as we see in the following example, computations much simplify if we replace $JT_{\mathbb{G}_{(r)},M}(-)$ by $JT^{exp}_{\mathfrak{g},M}(-)$.
Example 5.2.
Adopt the notation of Examples 2.6 and 2.10. Let $V$ be an $n$-dimensional $k$-vector space, and let $\alpha_{0},\ldots\alpha_{r-1}$ be be pair-wise commuting $p$-nilpotent endomorphisms of $V$. Denote by $M$ the $(\mathbb{G}_{a}^{r})_{1}$-module such that $x_{i}\in k(\mathbb{G}_{a}^{r})_{1}$ acts as $\alpha_{i}$ and denote by $N$ the $\mathbb{G}_{a(r)}$-module such that $u_{i}\in k\mathbb{G}_{a(r)}$ also acts as $\alpha_{i}$.
Let $\overline{x}$ be a scheme-theoretic point of $\mathbb{P}^{r-1}$, let $\overline{y}=\Psi(\overline{x})$, and let $\overline{x}^{(r-1)}=\Phi((\overline{y})=F^{r-1}(\overline{x})$, where $\Psi:w\mathbb{P}(1,p,\ldots,p^{r-1})\to\mathbb{P}^{r-1},\ \Phi:\mathbb{P}^{r-1}\to w\mathbb{P}(1,p,\ldots,p^{r-1})$ are discussed in Example 2.6. Then
$JT_{\mathfrak{g}_{a},r,N}^{exp}(\overline{y})\quad=\quad JT_{\mathbb{G}_{a}^{\times r},1,M}(\overline{x}^{(r-1)}).$
This is verified by observing that we may identify $\Phi^{*}(\Theta_{\mathbb{G}_{a},r}^{exp})$ with $F^{r-1*}(\Theta_{\mathbb{G}_{a},1})$ in $k[V_{1}(\mathbb{G}_{a}^{\times r})]\otimes k[z_{0},\ldots,z_{r-1}]/(z_{i}^{p})$.
Let $(\mathbb{G},\mathcal{E}_{r})$ be associated with an embedding $i:\mathbb{G}\hookrightarrow GL_{N}$ of exponential type of height $r$, so that $\mathcal{E}_{r}:\mathbb{G}_{a(r)}\to\mathbb{G}_{(r)}$ is the restriction of the exponential map for $GL_{N}$. Thus, $JT^{exp}_{\mathfrak{g},r,M}(-):\mathbb{P}\mathcal{C}_{r}(\mathbb{G})\quad\to\quad\mathcal{Y}_{\leq m}$ sends a scheme-theoretic point $\underline{B}\in\mathcal{C}_{r}(\mathfrak{g})$ to the Jordan type of $\sum_{s=0}^{r-1}(exp_{B_{s}})_{*}(u_{r-s-1})\in GL_{m}(k(\underline{B}))$ acting on $k(\underline{B})\otimes M$.
Example 5.3.
Let $\mathbb{G}=SL_{2(2)}$. Consider the irreducible $SL_{2(2)}$-modules $S_{\lambda}$ for $10\leq\lambda\leq p^{2}-$ 1, so that $S_{\lambda}=S(\lambda_{0})\otimes S(\lambda_{1})^{(1)}$ where $0\leq\lambda_{1},\lambda_{2}<p$ and $\lambda=\lambda_{0}+p\lambda_{1}$. What is implicit in this computation is that $u_{i}$ vanishes on $S(\lambda_{0})$ and on $S(\lambda_{1})$ for all $i\geq 1$. The dimensions of $S(\lambda_{0}),S(\lambda_{1})^{(1)},S(\lambda_{0})\otimes S(\lambda_{1})^{(1)}$ are respectively $m=\lambda_{0}+1,n=\lambda_{1}+1$, and $m\cdot n=(\lambda_{0}+1)\cdot(\lambda_{1}+1)$. To compute $JT_{SL_{2},2,S_{\lambda}}(B_{0},B_{1})$ for $B_{0},B_{1}$ commuting elements in $sl_{2}$, we appeal to Proposition 2.18 to reduce to the case that $B_{0}=a_{0}E,\ B_{1}=a_{1}E$ where $E$ is the $2\times 2$ matrix with $E_{1,2}=1$ and $E_{1,1}=E_{2,1}=E_{2,2}=0$.
Thus, to compute $JT_{sl_{2},2,S(\lambda)}^{exp}(-)$ it suffices to compute the Jordan types of $a_{1}^{p}E\cdot u_{0}+a_{0}E\cdot u_{1}$ on $S(\lambda_{0})\otimes S(\lambda_{1})^{(1)}$. This equals the Jordan type of $a_{0}E$ on $S(\lambda_{1})$ tensor (as Jordan types – see Definition 1.5) the Jordan type of $a_{1}^{p}E$ on $S(\lambda_{0})$. Observe that the Jordan type of $E$ on $S(\lambda_{0})=[m]$ and on $S(\lambda_{1})=[n]$. Thus, if both $a_{0},\ a_{1}$ are non-zero, then the the Jordan type of $a_{1}^{p}E\cdot u_{0}+a_{0}E\cdot u_{1}$ on $S(\lambda)$ is $[m]\otimes[n]$; if $a_{0}\not=0,\ a_{1}=0$ then the Jordan type $S(\lambda)$ is $m\cdot[n]$; and if $a_{0}=0,\ a_{1}\not=0$ then the Jordan type $S(\lambda)$ is $n\cdot[m]$
Example 5.4.
Let $V$ be a polynomial representation of $GL_{N}$ of degree $<p$ and consider the $i$-th Frobenius twist $V^{(i)}$ for some $i\geq 0$. We fix some $r>i$ and proceed to determine $JT_{\mathfrak{g}l_{N},r,V^{(i)}})(-)$. Then the action of $(\mathcal{E}_{B_{s}})_{*}(u_{r-s-1})$ on $V^{(i)}$ is non-trivial if and only if $10\leq s=r-i-$ 1. If $s=r-i-1$ so that $i=r-s-1$, then this action is given by the action of the exponential of $B_{s}$ and thus has Jordan type that of $B_{s}\in\mathfrak{g}l_{N}$ acting on $V$. Consequently, $JT_{\mathfrak{g}l_{N},r,V^{(i)}}^{exp}(\underline{B})=JT(B_{s},V)$ where $s=r-i-1$.
More generally, we can take a polynomial representation $V_{i}$ of $GL_{N}$ for each $i,0\leq i<r$ with each $V_{i}$ of degree $<p$ and consider $M=\otimes_{i=1}^{r-1}V_{i}^{(i)}$. Then we may use the preceding paragraph together with the rule for tensor product of Jordan types to determine $JT_{\mathfrak{g}l_{N},r,M}^{exp}(-)$ as in Example 5.3 in terms of the Jordan types of each $JT(B_{r-i-1},V_{i})$.
Example 5.5.
As a special case of Example 5.4, we take $V=\mathfrak{g}^{ad}$ to be the adjoint representation (i.e., equal to the Lie algebra $\mathfrak{g}L_{N}$ with conjugate action). In this case, $JT_{\mathfrak{g}l_{N},r,\mathfrak{g}^{ad}}^{exp}(\underline{B})$ is the Jordan type of $[-,B_{0}]:\mathfrak{g}l_{N}\to\mathfrak{g}l_{N}$.
This generalizes to $\mathbb{G}$ with $V=\mathfrak{g}^{ad}$ provided that $i:\mathbb{G}\hookrightarrow GL_{N}$ is an embedding of exponential type of height $r$.
6. Vector bundles
In this section we revisit the construction of vector bundles associated to certain $\mathbb{G}$-modules given in [14]. Although the discussion below applies more generally to $\mathbb{G}$-modules of constant $j$-type, we restrict our discussion to modules of constant Jordan type. We refer to [5] for an introduction to modules of constant Jordan type and for numerous examples of these interesting $\mathbb{G}$-modules.
The Jordan type functions we have considered are “pointwise” in the sense that their value at a point $\underline{B}\in\mathcal{C}_{r}(\mathfrak{g})$ depends upon the specialization of $k[\mathcal{C}_{r}(\mathfrak{g})]\otimes M$ along $\operatorname{Spec}\nolimits k(\underline{B})\to\mathcal{C}_{r}(\mathfrak{g})$. In [14], vector bundles were associated to $\mathbb{G}$-modules of constant Jordan type in part because their invariants can sometimes distinguish two $\mathbb{G}$-modules whose Jordan type functions are equal.
We proceed to verify that these vector bundles (i.e., locally free, coherent sheaves) are little changed if we replace the universal $p$-nilpotent operator $\Theta_{\mathbb{G},r}$ by the simpler $\Theta_{\mathfrak{g},r}^{exp}$.
Definition 6.1.
Let $(\mathbb{G},\mathcal{E}_{(r)}$ be an affine group scheme of exponential type of height $r$. Consider the bigraded algebra $k[s,t]\otimes k[\mathcal{C}_{r}(\mathfrak{g})]$ where both $s\otimes 1$ and $t\otimes 1$ are given the bigrading $(1,0)$. We define $\Theta_{\mathfrak{g}/\mathbb{P}^{1},r}$ to be
$\Theta_{\mathfrak{g}/\mathbb{P}^{1},r}\quad\equiv\quad(s\otimes 1)(1\otimes\Theta_{\mathfrak{g},r}^{\exp})\ +(1\otimes t)(1\otimes\Phi_{\mathfrak{g},r}^{*}(\Theta_{\mathbb{G}_{(r)}}))\ \in\ k[s,t]\otimes k[\mathcal{C}_{r}(\mathfrak{g})]\otimes k\mathbb{G}_{(r)}.$
So defined, $\Theta_{\mathbb{G}/\mathbb{P}^{1},r}$ is bihomogenous of bidegree $(1,p^{r-1})$, thereby determining a global section of $\mathcal{O}_{\mathbb{P}^{1}\times\mathbb{P}\mathcal{C}_{r}(\mathfrak{g})}(1,p^{r-1})$.
Remark 6.2.
Specializing Definition 6.1 to $\mathbb{G}=\mathbb{G}_{a}$, we see by Example 2.10 that $\Theta_{\mathbb{G},r}^{exp}$ can be identified with $\Theta_{\mathbb{G}_{a}^{\times r},1}$ as elements of $k[x_{0},\ldots,x_{r-1}]\otimes k[v_{0},\ldots,v_{r-1}]/(v_{i}^{p})$.
Proposition 6.3.
Retain the notation and hypotheses of Definition 6.1. Let $M$ be a finite dimensional $\mathbb{G}_{(r)}$-module. For any scheme theoretic point $(\underline{B},\langle s,t\rangle)\in\mathcal{C}_{r}(\mathfrak{g})\times\mathbb{P}^{1}$,
$JT(\Theta^{exp}_{\mathfrak{g},r,\underline{B}},k(\underline{B})\otimes M)\ =\ \frac{m}{p}[p]\ \ \iff\ \ JT(\Theta_{\mathfrak{g}/\mathbb{P}^{1},r,(\underline{B},\langle s,t\rangle)},k(\underline{B},\langle a,b\rangle\otimes M)\ =\ \frac{m}{p}[p].$
Proof.
Observe that
$\Theta_{\mathfrak{g}/\mathbb{P}^{1},r,(\underline{B},\langle s,t\rangle)}\ =\ s\cdot\Theta_{\mathfrak{g},r,\underline{B}}^{\exp}\ +\ t\cdot\Theta_{\mathfrak{g},r,\underline{B}}\quad\in\quad k(\underline{B},\langle s,t\rangle)\otimes k\mathbb{G}_{(r)}.$
As recalled in the proof of Proposition 4.10, $\Theta_{\mathfrak{g},r,\underline{B}}-\Theta_{\mathfrak{g},r,\underline{B}}^{\exp}$ is a linear combination of terms each of which is a product of at least 2 $p$-nilpotent factors. Thus, proposition follows from [6, Prop 8]. ∎
In the following proposition, we utilize $\Phi_{\mathfrak{g},r}:\mathcal{C}_{r}(\mathfrak{g})\to V_{r}(\mathbb{G})$, the bicontinuous morphism of Lemma 4.7, and the construction in [14, Prop 4.8] of the quasi-coherent sheaf ${\mathcal{K}er}\{(\tilde{\Theta}_{\mathbb{G},r,M})^{j}\})$ on $\mathbb{P}V_{r}(\mathfrak{g})$. As seen [14], these kernel sheaves (as well as cokernel and image sheaves) are algebraic vector bundles provided that $M$ has constant Jordan type.
Proposition 6.4.
Let $(\mathbb{G},\mathcal{E}_{(r)})$ be an affine group scheme of exponential type of height $r$, and let $M$ be a $\mathbb{G}_{(r)}$-module. Then for any $j,0<j<p$, there is a quasi-coherent sheaf ${\mathcal{K}er}\{(\tilde{\Theta}_{\mathfrak{g}/\mathbb{P}^{1},r,M})^{j}\}$ on $\mathbb{P}^{1}\times\mathbb{P}\mathcal{C}_{r}(\mathfrak{g})$ satisfying
  1. (1)
    The fiber over $\langle 0,t\rangle\in\mathbb{P}^{1}$ of ${\mathcal{K}er}\{(\tilde{\Theta}_{\mathfrak{g}/\mathbb{P}^{1},r,M})^{j}\}$ is the quasi-coherent sheaf $\Phi_{\mathfrak{g},r}^{*}({\mathcal{K}er}\{(\tilde{\Theta}_{\mathbb{G},r,M})^{j}\})$, the pull-back along $\Phi_{\mathfrak{g},r}$ of the quasi-coherent sheaf associated to the graded $k[V_{r}(\mathfrak{g})]$-module ${\mathcal{K}er}\{(\tilde{\Theta}_{\mathbb{G},r,M}^{j}:k[V_{r}(\mathfrak{g})]\otimes M\to(k[V_{r}(\mathfrak{g})]\otimes M)(jp^{r-1})\}.$
  2. (2)
    The fiber over $\langle s,0\rangle\in\mathbb{P}^{1}$ of ${\mathcal{K}er}\{(\tilde{\Theta}_{\mathfrak{g}/\mathbb{P}^{1},r,M})^{j}\}$ is the quasi-coherent sheaf associated to the graded $k[\mathcal{C}_{r}(\mathfrak{g})]$-module
    $ker\{(\Theta_{\mathfrak{g},r,M}^{exp})^{j}:k[\mathcal{C}_{r}(\mathfrak{g})]\otimes M\to(k[\mathcal{C}_{r}(\mathfrak{g})]\otimes M)(jp^{r-1})\}.$
  3. (3)
    If $M$ is a finite dimensional $\mathbb{G}_{(r)}$-module of constant Jordan type, then ${\mathcal{K}er}\{(\tilde{\Theta}_{\mathfrak{g}/\mathbb{P}^{1},r,M})^{j}\}$ is a locally free, coherent sheaf which restricts to the locally free coherent sheaves $\Phi_{\mathfrak{g},r}^{*}({\mathcal{K}er}\{(\tilde{\Theta}_{\mathbb{G},r,M})^{j}\}),\ {\mathcal{K}er}\{(\tilde{\Theta}_{\mathfrak{g},r,M}^{exp})^{j}\}$ on $\mathbb{P}\mathcal{C}_{r}(\mathfrak{g})$.
In other words, for any $\mathbb{G}_{(r)}$-module of constant Jordan type, taking kernels of $(\tilde{\Theta}_{\mathbb{G},r,M})^{j}$ and $(\tilde{\Theta}_{\mathfrak{g},r,M}^{exp})^{j}$ yield $\mathbb{A}^{1}$-equivalent vector bundles on $\mathbb{P}\mathcal{C}_{r}(\mathfrak{g})$.
A completely parallel construction applies to taking images of images and cokernels of $(\tilde{\Theta}_{\mathbb{G},r,M})^{j}$ and $(\tilde{\Theta}_{\mathbb{G},r,M}^{exp})^{j}$.
Proof.
We consider the bigraded algebra $k[s,t]\otimes k[\mathcal{C}_{r}(\mathfrak{g})]$ where both $s\otimes 1$ and $t\otimes 1$ are given the bigrading $(1,0)$. We define $\Theta_{\mathfrak{g}/\mathbb{P}^{1},r}$ to be
$\Theta_{\mathfrak{g}/\mathbb{P}^{1},r}\quad\equiv\quad(s\otimes 1)(1\otimes\Theta_{\mathfrak{g},r}^{\exp})\ +(1\otimes t)(1\otimes\Phi_{\mathfrak{g},r}^{*}(\Theta_{\mathbb{G}_{(r)}}))\ \in\ k[s,t]\otimes k[\mathcal{C}_{r}(\mathfrak{g})]\otimes k\mathbb{G}_{(r)}.$
By Theorem 4.8(1), $\Theta_{\mathfrak{g},r}=\Phi_{\mathfrak{g},r}^{*}(\Theta_{\mathbb{G}_{(r)}})$. By Proposition 4.5 and Theorem 4.8, $\Theta_{\mathfrak{g}/\mathbb{P}^{1},r}$ is bihomogeneous of bidegree $(1,p^{r-1})$. Thus, the bigraded module
(6.4.1) $ker\{(\Theta_{\mathfrak{g}/\mathbb{P}^{1},r})^{j}:(k[s,t]\otimes k[\mathcal{C}_{r}(\mathfrak{g})]\otimes M)\ \to\ (k[s,t]\otimes k[\mathcal{C}_{r}(\mathfrak{g})]\otimes M)[1,p^{r-1}]$
determines the quasi-coherent sheaf ${\mathcal{K}er}\{(\tilde{\Theta}_{\mathfrak{g}/\mathbb{P}^{1},r,M})^{j}\}$ on $\mathbb{P}^{1}\times\mathbb{P}\mathcal{C}_{r}(\mathfrak{g})$. The first two assertions follow from the observation that the fiber above $\langle a,b\rangle\in\mathbb{P}^{1}$ of ${\mathcal{K}er}\{(\tilde{\Theta}_{\mathfrak{g}/\mathbb{P}^{1},r,M})^{j}\}$ is the sheaf associated to the bigraded graded module (6.4.1) obtained by extending along $k[s,t]\to k$ sending $s$ to $a$ and $t$ to $b$.
For $M$ a finite dimensional $\mathbb{G}_{(r)}$-module of constant Jordan type, the verification of [14, Thm 4.13] showing that ${\mathcal{K}er}\{(\tilde{\Theta}_{\mathbb{G},r,M})^{j}\}$ is a locally free, coherent sheaf applies to ${\mathcal{K}er}\{(\tilde{\Theta}_{\mathfrak{g}/\mathbb{P}^{1},r,M}\}$ as well as to $\Phi_{\mathfrak{g},r}^{*}({\mathcal{K}er}\{(\tilde{\Theta}_{\mathbb{G},r,M})^{j}\})$ and ${\mathcal{K}er}\{(\tilde{\Theta}_{\mathfrak{g},r,M}^{exp})^{j}\}))$, showing that each of these sheaves is locally free and coherent. ∎