### v2

# Invariants for $\mathbb{G}_{(r)}$-modules

Abstract.
We revisit the constructions given by J. Pevtsova and the author of refined invariants for finite dimensional representations of infinitesimal group schemes $\mathbb{G}_{(r)}$ over a field $k$ of characteristic $p>0$. Our focus is on the universal $p$-nilpotent operator seen as an element in the group algebra of the group scheme $\mathbb{G}_{(r),X}$ over $X$, where $X$ is either the moduli space $V_{r}(\mathbb{G})$ of height $r$ 11-parameter subgroups of $\mathbb{G}$ or the moduli space $\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))$ of $r$-tuples of $p$-nilpotent, pair-wise commuting elements of the Lie algebra of $\mathbb{G}$.
We formalize Jordan type function using several variants of the continuous function $JT_{\mathbb{G},r,M}(-):\mathbb{P}V_{r}(\mathbb{G})\to\mathcal{Y}$ where $\mathcal{Y}$ is the poset of Young diagrams with $p$-columns. One of these variants is designed to be more conducive to computation. The vector bundle construction given by J. Pevtsova and the author is extended to all finite dimensional $\mathbb{G}_{(r)}$-modules, producing coherent sheaves on $X$ which are locally free on the strata of $X$ associated to $JT_{\mathbb{G},r,M}(-)$.
Key words and phrases:
Jordan type, infinitesimal group scheme, universal operators
2020 Mathematics Subject Classification:
20G05, 14L17
partially supported by the Simons Foundation
0. Introduction
In the present work, we utilize methods and results of support theory to investigate explicit invariants for specific finite dimensional modules for certain classes of infinitesimal group schemes over an arbitrary base field $k$ of characteristic $p$. This is in contrast to the role of support theory in establishing properties for various categories of representations of finite group schemes as in [3]. Our techniques do not apply to finite groups, for we utilize the support theory developed in collaboration with A. Suslin and C. Bendel in [26], [27] which replaces cohomological varieties by varieties of 1-parameter subgroups.
We further investigate the “Jordan type function” introduced by J. Pevtsova and the author in [15] and [16]. For a given affine group scheme $\mathbb{G}$ of finite type over $k$ and a finite dimensional $\mathbb{G}$-module $M$ for the $r$-th Frobenius kernel $\mathbb{G}_{(r)}$ of $\mathbb{G}$, we formulate the continuous function $JT_{\mathbb{G},r,M}(-):\mathbb{P}V_{r}(\mathbb{G})\to\mathcal{Y}$ from the scheme-theoretic points of $\mathbb{P}V_{r}(\mathbb{G})$ representing height $r$ 1-parameter subgroups to the partially ordered set $\mathcal{Y}$ of Young diagrams with $p$ columns. By definition, $JT_{\mathbb{G},r,M}(-)=JT_{\mathbb{G}_{(r)},r,M}(-)$. The continuity of $JT_{\mathbb{G},r,M}(-)$ implies that if a point $x\in\mathbb{P}V_{r}(\mathbb{G})$ specializes to $y\in\mathbb{P}V_{r}(\mathbb{G})$ then the Young diagram $JT_{\mathbb{G},r,M}(x)$ is greater or equal to the Young diagram $JT_{\mathbb{G},r,M}(y)$.
This “theory” $M\mapsto JT_{\mathbb{G},r,M}(-)$ constitutes a functor $JT_{\mathbb{G},r}:mod(\mathbb{G}_{(r)})\to Hom_{cont}(\mathbb{P}V_{r}(\mathbb{G}),\mathcal{Y})$ from finite dimensional $\mathbb{G}_{(r)}$-modules to continuous functions $\mathbb{P}V_{r}(\mathbb{G})\to\mathcal{Y}$ which captures much more information about a $\mathbb{G}_{(r)}$-module $M$ than does the cohomological support theory $M\mapsto\mathbb{P}H^{*}(\mathbb{G}_{(r)},k)_{M}$. For example, if $p$ does not divide the dimension of $M$, then $\mathbb{P}H^{*}(\mathbb{G}_{(r)},k)_{M}$ equals $\mathbb{P}H^{*}(\mathbb{G}_{(r)},k)$ and therefore is independent of $M$, whereas $M\ \mapsto\ JT_{\mathbb{G},r,M}(-)$ distinguishes many such $\mathbb{G}_{(r)}$-modules $M$ of a given dimension. As might be expected, determination of $JT_{\mathbb{G},r,M}(-)$ is a challenge. With this in mind, we introduce a simplified Jordan type function $M\ \mapsto\ JT^{exp}_{\mathfrak{g},r,M}(-)$ for affine group schemes $\mathbb{G}$ which are of exponential type of height $r$.
After an elementary discussion of Jordan types, we provide a detailed formalism for the construction of the “universal $p$-nilpotent operator” $\Theta_{\mathbb{G},r}\in k[V_{r}(\mathbb{G})]\otimes k\mathbb{G}_{(r)}\equiv k\mathbb{G}_{(r),V_{r}(\mathbb{G})}$ of height $r$ constructed by J. Pevtsova and the author in [16]. In doing so, we expand upon the connection between 1-parameter subgroups for $\mathbb{G}$ and operators on $\mathbb{G}_{(r)}$-modules. The operator $\Theta_{\mathbb{G},r}$ acting on $k[V_{r}(\mathbb{G})]\otimes M$ determines the Jordan type function $JT^{exp}_{\mathfrak{g},r,M}(-)$. The categorically inclined reader might want to categorify the functor $JT_{\mathbb{G},r}$ by replacing $Hom_{cont}(\mathbb{P}V_{r}(\mathbb{G}),\mathcal{Y})$ by the category of graded $k[V_{r}(\mathbb{G})]$-modules equipped with a $p$-nilpotent homogeneous operator.
In order to make the Jordan type function more explicit and computationally accessible, we restrict our attention to affine group schemes $\mathbb{G}$ which are of exponential type of height $r$. By definition, such groups $\mathbb{G}$ are equipped with an exponential map with the property that height $r$ 1-parameter subgroups of $\mathbb{G}$ are given by applying the exponential map to $r$-tuples of $p$-nilpotent, pair-wise commuting elements of the Lie algebra $\mathfrak{g}$ of $\mathbb{G}$. Thus, the “moduli space” for height $r$ 1-parameter subgroups of $\mathbb{G}$ (of exponential type $r$) is the variety $\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))$ of such $r$-tuples of elements in $\mathfrak{g}$. For $\mathbb{G}=GL_{N}$, the associated exponential map is the usual (truncated) exponential sending a $p$-nilpotent element of $A\otimes\mathfrak{g}l_{N}$ to its exponential in $GL_{N}(A)$ for any commutative $k$-algebra AA.
For a given finite dimensional $\mathbb{G}_{(r)}$-module $M$ with $\mathbb{G}$ of exponential type of height $r$, the Jordan type function becomes a continuous function $JT_{\mathfrak{g},r,M}(-):\mathbb{P}\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))\to\mathcal{Y}$ associated to the “projectivization” of the universal $p$-nilpotent operator $\Theta_{\mathfrak{g},r}\in k[\mathcal{C}_{r}(\mathcal{N}_{p}\mathfrak{g}))]\otimes k\mathbb{G}_{(r)}$ corresponding to $\Theta_{\mathbb{G},r}$. Our exponential Jordan type function $JT_{\mathfrak{g},r,M}^{exp}(-):\mathbb{P}\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))\to\mathcal{Y}$ is defined by replacing $\Theta_{\mathfrak{g},r}$ by its linear part (denoted $\Theta_{\mathfrak{g},r}^{exp}$) with respect to a natural grading on the group algebra $k\mathbb{G}_{a(r)}$ of $\mathbb{G}_{a(r)}$. In the very special case that $\mathbb{G}$ equals $\mathbb{G}_{a(r)}$, this simplification is related to J. Carlson’s use of cyclic shifted subgroups to identify support varieties for $(\mathbb{Z}/p)^{\times r}$-modules (see [5]).
The suitability of $\Theta_{\mathfrak{g},r}^{exp}$ follows from work of Pevtsova, Suslin, and the author [17] and subsequent work of P. Sobaje [24]. Namely, for a $\mathbb{G}_{(r)}$-module $M$ of dimension $m$ and a point of $x\in\mathbb{P}\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g})$), $JT_{\mathfrak{g},r,M}(x)\ =\ m/p\cdot[p]$ if and only if $JT^{exp}_{\mathfrak{g},r,M}(x)\ =\ m/p\cdot[p]$. Thus, the replacement of $JT_{\mathbb{G},r,M}(-)$ by $JT_{\mathfrak{g},r,M}^{exp}(-):\mathbb{P}\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))\to\mathcal{Y}$ does not affect (refined) support varieties considered in previous work of Pevtsova and the author in [15] and [16]. We provide sample computations which demonstrate the accessibility of $JT_{\mathfrak{g},r,M}^{exp}(-)$ for certain $\mathbb{G}_{(r)}$-modules $M$.
For $\mathbb{G}$ of exponential type of height $r$, we verify that vector bundles on $\mathbb{P}\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))$ associated to $\mathbb{G}_{(r)}$-modules $M$ of constant Jordan type using $\Theta_{\mathfrak{g},r}$ (naturally corresponding to vector bundles previously constructed in [16]) are $\mathbb{A}^{1}$-homotopy equivalent to vector bundles using $\Theta_{\mathfrak{g},r}^{exp}$. We extend these constructions of vector bundles to apply to arbitrary finite dimensional $\mathbb{G}_{(r)}$-modules $M$, yielding coherent sheaves on $\mathbb{P}\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))$ whose restrictions on strata associated to the Jordan type function for $M$ are locally free. Unlike the situation for modules of constant Jordan type, these stratified coherent sheaves associated to a given finite dimensional $\mathbb{G}_{(r)}$-module will typically depend upon whether we use the Jordan type function associated to $\Theta_{\mathbb{G},r}$ or associated to $\Theta_{\mathfrak{g},r}^{exp}$.
As we observe in the last section, the issue of “stabilization with respect to rr” of Jordan type functions of height $\leq r$ is closely related to the Jordan type function for finite dimensional rational $\mathbb{G}$-modules introduced in [12] if $\mathbb{G}$ is of exponential type of infinite height. This corrects a minor error of [12, Prop 4.8] which incorrectly formulated such a stabilization result.
We thank Julia Pevtsova and Paul Sobaje for their insights into the mathematics we present.
1. Jordan types and specialization
We begin by recalling some basic properties of Jordan forms of $p$-nilpotent endomorphisms.
Definition 1.1.
Let $M$ be a $k[t]/t^{p}$-module of dimension $m$ given by the map $\rho_{M}:k[t]/t^{p}\to M_{m,m}(k)$ of $k$-algebras. Let $\underline{a}=(a_{1},\ldots,a_{p})\in\mathbb{N}^{\times p}$ (i.e., a $p$-tuple of non-negative integers) be defined by setting $a_{i}$ to be the number of blocks of size $i$ of the Jordan canonical form of the matrix $\rho_{M}(t)$. We say that the Jordan type of $M$ is $\underline{a}$, expressing this as $JT(M)=\underline{a}$ or as $JT(M)=a_{p}[p]+a_{p-1}[p-1]+\cdots+a_{1}[1]$.
The following proposition recalls familiar interpretations of Jordan types.
Proposition 1.2.
Let $M$ be a $k[t]/t^{p}$-module of dimension $m$ such that $JT(M)=\underline{a}$.
  1. (1)
    As a $k[t]/t^{p}$-module, $M$ is isomorphic to $\bigoplus_{i=1}^{p}(k[t]/t^{i})^{\oplus a_{i}}$.
  2. (2)
    The partition of $m=\sum_{i=1}^{p}a_{i}\cdot i$ with $a_{i}$ subsets of size $i$ is naturally associated to $\underline{a}=(a_{1},\ldots,a_{p})$.
  3. (3)
    The usual partial ordering of partitions determines the partial ordering on $\mathbb{N}^{\times p}$ given by
    $\underline{a}\ \leq\ \underline{b}\quad\iff\quad\forall s,1\leq s\leq p,\quad(\sum_{i=s}^{p}a_{i}\cdot i)\leq(\sum_{i=s}^{p}b_{i}\cdot i).$
  4. (4)
    Uniquely associated to $\underline{a}$ is a Young diagram with $p$-columns such that the diagram has $a_{i}$ rows with $i$ boxes and the rows of the diagram are left-adjusted with descending length. Thus, the partial ordering $\underline{a}\ \leq\ \underline{b}$ requires for each $\ell$ that the number of boxes in the top $\ell$ rows of $\underline{a}$ is less or equal to the number of boxes in the top $\ell$ rows of $\underline{b}$.
  5. (5)
    The Jordan type of $t^{j}$ acting on $M\simeq\oplus_{i=1}^{p}(k[t]/t^{i})^{\oplus a_{i}}$ (denoted by $\underline{a}^{j}\in\mathbb{N}^{\times p}$) has associated Young diagram obtained from the Young diagram for $\underline{a}$ by removing the first (i.e., left-most) j1j-1 columns of $\underline{a}$ and adding rows with each having only 1 box so that the number of boxes of $\underline{a}^{j}$ equals $m$.
The lemma below recalls some properties of the partial ordering on Jordan types. We denote by $rk(t^{j}:M\to M)$ the dimension of the subspace $t^{j}(M)\subset M$, the dimension of the image of multiplication by $t^{j}$ on the vector space $M$.
Proposition 1.3.
Consider $m$-dimensional $k[t]/t^{p}$-modules M,NM,\ N isomorphic to $\bigoplus_{i=1}^{p}(k[t]/t^{i})^{\oplus a_{i}},\ \bigoplus_{i=1}^{p}(k[t]/t^{i})^{\oplus b_{i}}$. As above, let $\underline{a}^{s}$ (respectively, $\underline{b}^{s}$) denotes the Jordan type of $t^{s}$ acting on $M$ (resp, $N$) for any $s,\ 1\leq s<p$.
  1. (1)
    $rk(t^{s}:M\to M)\ =\ \sum_{i=s+1}^{p}a_{i}(i-1-s)$.
  2. (2)
    For all $s,\ 1\leq s<p$, $\underline{a}^{s}\ =\ \sum_{i=s+2}^{p}a_{i}[i-s]+(m-rk(t^{s}:M\to M))[1]$.
  3. (3)
    $\underline{a}\leq\underline{b}\ \iff\underline{a}^{s}\leq\underline{b}^{s},\ \forall s,1\leq s<p$.
  4. (4)
    For all $s,\ 1\leq s<p$, $\underline{a}^{s}\ \leq\underline{b}^{s}\ \iff rk(t^{s}:M\to M)\leq rk(t^{s}:N\to N).$
Proof.
Assertions (1) and (2) are verified by an easy computation. Assertion (3) follows from assertion (2) and Proposition 1.2(3). Assertion (4) follows easily from assertion (2). ∎
Since a surjective map $f:M\to N$ of $k[t]/t^{p}$-modules induces a surjective map $f:t^{j}(M)\to t^{j}(N)$ for any $j$ with $1\leq j<p$, Proposition 1.3(4) implies the following additional corollary.
Corollary 1.4.
Let $f:M\to N$ be a surjective map of finite dimensional $k[t]/t^{p}$-modules. Then $JT(N)\leq JT(M)$, with equality if and only if $f$ is an isomorphism.
We formulate the partially ordered abelian monoid $\mathcal{Y}_{\leq}$ of “Young diagrams” equipped with a (tensor) product.
Definition 1.5.
Denote by $\mathcal{Y}_{\leq}$ the monoid $\mathbb{N}^{\times p}$ provided with the partial ordering of Proposition 1.2(3), and denote by $\mathcal{Y}_{\leq,n}$ the partially ordered subset of $\mathcal{Y}_{\leq}$ whose objects are $p$-tuples $\underline{a}$ with $\sum_{i=1}^{p}a_{i}\cdot\ \leq\ n$.
We equip $\mathcal{Y}_{\leq}$ with the bi-additive pairing
$(-)\otimes(-):\ \mathcal{Y}_{\leq}\times\mathcal{Y}_{\leq}\quad\to\quad\mathcal{Y}_{\leq}\ ,$
where $\underline{a}\otimes\underline{b}$ is the $p$-tuple $\underline{c}$ defined by the isomorphism
$\bigoplus_{i=1}^{p}(k[t]/t^{i})^{\oplus c_{i}})\quad\simeq\quad(\bigoplus_{i=1}^{p}(k[t]/t^{i})^{\oplus a_{i}})\ \bigotimes\ (\bigoplus_{i=1}^{p}(k[t]/t^{i})^{\oplus b_{i}})$
of $k[t]t^{p}$-modules (see [7, Cor10.3]).
We recall the lower semi-continuity of Jordan type, an easy consequence of Nakayama’s Lemma (see [17, Prop 3.2]).
Proposition 1.6.
Consider a noetherian, local domain $R$ with field of fractions $L$ and residue field $K$. Let $M_{R}$ be a free $R$-module of finite rank equipped with a $p$-nilpotent endomorphism $\theta_{R}\in End_{R}(M_{R},M_{R})$. Denote $K\otimes_{R}M$ by $M_{K}$ and $L\otimes_{R}M$ by $M_{L}$, and set
$\theta_{K}=K\otimes_{R}\theta_{R}\in End_{K}(M_{K},M_{K}),\quad\theta_{L}=L\otimes_{R}\theta_{R}\in End_{L}(M_{L},M_{L}).$
So defined, $M_{K}$ is a $K[t]/t^{p}$-module with the action of $t$ given by $\theta_{K}$ and $M_{L}$ is an $L[t]/t^{p}$-module with the action of $t$ given by $\theta_{L}$. We say that the $L[t]/t^{p}$-module $M_{L}$ specializes along $R$ to the $K[t]/t^{p}$-module $M_{K}$.
For such a specialization, $JT(M_{K})\ \leq\ JT(M_{L})$.
Proof.
By applying Nakayma’s Lemma [22, Lem 4.3] to the cokernels of the maps $(\theta_{R})^{i}:M_{R}\to M_{R}$, we conclude that $dim(coker(\theta_{L}^{i}))\leq dim(coker(\theta_{K}^{i}))$ so that $rk(\theta_{K}^{i})\leq rk(\theta_{L}^{i})$ for each $i,\ 1\leq i<p$. Thus, the proposition follows from Proposition 1.3. ∎
Definition 1.7.
Let $X$ be a scheme of finite type over $k$. For two scheme-theoretic points $x,y$ of $X$, we write $x\leq y$ if $x$ is in the (Zariski) closure of $y$. This endows the set of scheme-theoretic points of $X$ with a natural partial ordering.
If $\operatorname{Spec}\nolimits R\subset X$ is an affine open subset of $X$ containing $x$, then the scheme-theoretic point $x\in\operatorname{Spec}\nolimits R$ corresponds to the prime ideal $\mathfrak{p}_{x}\subset R$. The structure sheaf $\mathcal{O}_{X}$ of $X$ has stalk the local ring $R_{(\mathfrak{p}_{x})}$ with residue field $k(x)=R_{(\mathfrak{p}_{x})}/\mathfrak{p}_{x}R_{(\mathfrak{p}_{x})}$. We identify the scheme-theoretic point $x\in X$ with the associated map of schemes $\chi_{x}:\operatorname{Spec}\nolimits k(x)\to X$.
Lemma 1.8.
Let $x,y$ be two scheme-theoretic points of $X$. Then $x\leq y$ if and only if there exists a local domain $R$ with field of fractions $k(y)$ and residue field $k(x)$ and a morphism $\operatorname{Spec}\nolimits R\to X$ whose restriction to $\operatorname{Spec}\nolimits k(y)$ is $\chi_{y}:\operatorname{Spec}\nolimits k(y)\to X$ and whose restriction to $\operatorname{Spec}\nolimits k(x)$ is $\chi_{x}:\operatorname{Spec}\nolimits k(x)\to X$.
Proof.
We may assume $x\not=y$. Let $\operatorname{Spec}\nolimits A$ be a Zariski open subset of $X$ containing $x$ with corresponding prime $\mathfrak{p}_{x}\subset A$. Since $x$ is in the closure of $y$, the prime ideal $\mathfrak{p}_{y}$ is contained in $\mathfrak{p}_{x}$. Set $S$ equal to the quotient $A/\mathfrak{p}_{y}$ and let $R$ be the localization of $S$ at (the image of ) $\mathfrak{p}_{x}$. So defined, $R$ is a local domain with residue field $k(x)$ and field of fractions $k(y)$. Moreover, $\chi_{x}:\operatorname{Spec}\nolimits k(x)\to X$ and $\chi_{y}:\operatorname{Spec}\nolimits k(y)\to X$ factor through $\operatorname{Spec}\nolimits R\subset\operatorname{Spec}\nolimits S\subset\operatorname{Spec}\nolimits A\subset X$.
Conversely, assume given a local domain $R$ with field of fractions $k(y)$ and residue field $k(x)$ and a morphism $\operatorname{Spec}\nolimits R\to X$ whose restriction to $\operatorname{Spec}\nolimits k(x)$ is $\chi_{x}:\operatorname{Spec}\nolimits k(x)\to X$ and whose restriction to $\operatorname{Spec}\nolimits k(y)$ is $\chi_{y}:\operatorname{Spec}\nolimits k(y)\to X$. Then the kernel of $R\to k(x)$ equals the prime ideal $\mathfrak{p}_{x}$ and contains $\mathfrak{p}_{y}$, the kernel of $R\to k(y)$. Thus, $x\leq y$. ∎
The following proposition will be employed with $A=k[V_{r}(\mathbb{G})]$ as in Definition 2.1 and with the action of $t$ given by various powers of the universal $p$-nilpotent operator of Theorem 2.15. We point out that in this proposition we consider the tensor product over the commutative algebra $A$ of $A$-modules; in discussion below, we shall frequently consider tensor product over $k$ of $\mathbb{G}$-modules using the Hopf algebra structure of $k[\mathbb{G}]$.
Proposition 1.9.
Let $A$ be a Noetherian commutative $k$-algebra and consider a finitely generated, projective $A$-module $M_{A}$ equipped with a $p$-nilpotent endomorphism $\theta_{A}\in Hom_{A}(M_{A},M_{A})$. Then sending a scheme-theoretic point $x\in\operatorname{Spec}\nolimits A$ (i.e., a prime ideal $\mathfrak{p}_{x}\subset A$) to the Jordan type of $k(x)\otimes_{A}M_{A}$ as a $k(x)[t]/t^{p}$-module with $t$ acting as $k(x)\otimes\Theta_{A}\in Hom_{k(x)}(M\otimes k(x),M\otimes k(x))$ determines a map of partially ordered sets
(1.9.1) $JT_{A,M_{A}}(-):(\operatorname{Spec}\nolimits A)\quad\to\quad\mathcal{Y}_{\leq}\ ,\quad x\mapsto JT(k(x)\otimes_{A}M_{A}).$
If $M_{A},N_{A}$ are finitely generated $A[t]/t^{p}$-modules, then for all $x\in\operatorname{Spec}\nolimits A$
$JT_{A,M_{A}}(x)\oplus JT_{A,N_{A}}(x)\quad=\quad JT_{A,M_{A}\oplus N_{A}}(x),$
$JT_{A,M_{A}}(x)\otimes JT_{A,N_{A}}(x)\quad=\quad JT_{A,M_{A}\otimes_{A}N_{S}}(x).$
Proof.
The assertion that $JT_{A,M_{A}}(-)$ is a map of posets is the assertion that $x\leq y\in\operatorname{Spec}\nolimits A$ implies that $JT_{A,M_{A}}(x)\leq JT_{N,N_{A}}(y)$. As in Lemma 1.8, we may restrict $M_{A}$ along a map from $A$ to a local domain $R$ with field of fractions $k(y)$ and residue field $k(x)$. Thus, we reduce to the case that $A=R$ is a local domain with field of fractions $k(y)$ and residue field $k(x)$; in this case, $M_{A}$ is a free $A$-module of finite rank. Thus, Proposition 1.6 applies to prove $JT_{A,M_{A}}(x)\leq JT_{A,M_{A}}(y)$.
Observe that specialization along a scheme-theoretic point of $\operatorname{Spec}\nolimits A$ commutes with direct sums and tensor products. Thus, the fact that the function $JT_{A,-}(x)$ (sending a finitely generated $A[t]/t^{p}$-module $M_{A}$ to $JT_{A,M_{A}}(x)\in\mathcal{Y}_{\leq}$) respects both addition and tensor product is immediate from the definitions of addition and tensor product for $\mathcal{Y}_{\leq}$. ∎
Proposition 1.9 does not immediately imply that $JT_{A,M_{A}}(-):\operatorname{Spec}\nolimits A\ \to\ \mathcal{Y}_{\leq}$ is continuous when $\operatorname{Spec}\nolimits A$ is given the Zariski topology. This is because arbitrary unions of closed subsets (i.e., subsets closed under specialization) are often not closed in the Zariski topology. We supplement Proposition 1.9 with another consequence of Nakayama’s Lemma in commutative algebra. (It seems more convenient for this lemma to refer to the scheme-theoretic points of $\operatorname{Spec}\nolimits A$ as prime ideals $\mathfrak{p}\subset A$.)
Lemma 1.10.
Let $A$ be a Noetherian commutative $k$-algebra and consider a projective $A$-module $M_{A}$ of rank $m$ equipped with a $p$-nilpotent endomorphism $\theta_{A}:M_{A}\to M_{A}$. Then for any $d>0$,
(1.10.1) $C_{d}\ \equiv\ \{\mathfrak{p}\in\operatorname{Spec}\nolimits A:\ rk(\theta_{\mathfrak{p}}:k(\mathfrak{p})\otimes_{A}M_{A}\to k(\mathfrak{p})\otimes_{A}M_{A})\leq d\}$
is a closed subset of $\operatorname{Spec}\nolimits R$.
Proof.
By considering the intersection of $C_{d}$ with each member of an affine open covering of $\operatorname{Spec}\nolimits A$ restricted to which $M_{A}$ is free, it suffices to prove the lemma assuming $M_{A}$ is free. Moreover, if $\mathfrak{q}_{1},\ldots,\mathfrak{q}_{r}$ are the minimal primes of $\operatorname{Spec}\nolimits A$, then the closures of $\{\mathfrak{q}_{i}\}$ are the (closed) irreducible components of $\operatorname{Spec}\nolimits A$. Thus, by considering the intersection of $C_{d}$ with each of these closures, it suffices to assume that $A$ is a domain.
We view $\theta_{A}$ as an element of $M_{m,m}(A)$. Then $C_{d}$ is the closed subset defined as the zero locus of the determinants of all $(d+1\times d+1)$-submatrices of $\theta_{A}$. ∎
We remind the reader that an Alexandrov discrete topology on a $T_{0}$ space is a topology in which the arbitrary union of closed subsets is closed. (See [1] and [21].) The topology on $\mathcal{Y}_{\leq}$ which we utilize is such a topology.
Definition 1.11.
We give $\mathcal{Y}_{\leq}$ the topology whose closed subsets are those subsets $S\subset\mathcal{Y}_{\leq}$ satisfying the condition that if $s\in S$ and $s^{\prime}\leq s$ then $s^{\prime}\in S$.
In particular, $S\subset\mathcal{Y}_{\leq}$ is closed if and only if $S\cap\mathcal{Y}_{\leq,m}$ is closed in $\mathcal{Y}_{\leq,m}$ for all $m>0$.
Theorem 1.12.
Let $A$ be a Noetherian commutative $k$-algebra and consider a projective $A$-module $M_{A}$ of rank $m$ equipped with a $p$-nilpotent endomorphism $\theta_{A}$. As usual, we give $\operatorname{Spec}\nolimits A$ the Zariski topology. Then
$JT_{A,M_{A}}(-):\operatorname{Spec}\nolimits A\quad\to\quad\mathcal{Y}_{\leq,m}\quad\hookrightarrow\quad\mathcal{Y}_{\leq}$
is continuous.
Proof.
A subset $C\subset\mathcal{Y}_{\leq,m}$ is closed if and only
$C\quad=\quad\bigcup_{\underline{a}\in C}\overline{\{\underline{a}\}},$
where $\overline{\{\underline{a}\}}$ is the closure of the singleton subset $\{\underline{a}\}$. Thus, it suffices to prove that $JT_{A,M_{A}}^{-1}(\overline{\{\underline{a}\}})$ is closed for all $\underline{a}\in\mathcal{Y}_{\leq,m}$ (since $\mathcal{Y}_{\leq,m}$ is finite). By definition, $JT_{A,M_{A}}^{-1}(\overline{\{\underline{a}\}})$ consists of those $x\in\operatorname{Spec}\nolimits A$ such that $JT_{A,M_{A}}(x)\ \leq\ \underline{a}$ in $\mathcal{Y}_{\leq,m}.$ By Proposition 1.3 (4), $JT_{A,M_{A}}(x)\ \leq\ \underline{a}$ if and only if the rank of $(\theta_{x})^{s}:k(x)\otimes_{A}M_{A}\to k((x)\otimes_{A}M_{A}$ is less than or equal to the rank of $t^{s}$ on $\oplus_{i=1}^{p}(k[t]/t^{i})^{\oplus a_{i}}$ (which equals $\sum_{i=s+1}^{p}a_{i}\cdot(i-s)$) for all $s$, $1\leq s<p$.
Applying Lemma 1.10 to each of $\theta^{s},1\leq s<p$, we conclude the set of those $x\in\operatorname{Spec}\nolimits A$ such that $JT_{A,M_{A}}(x)\ \leq\ \underline{a}$ is closed in $\operatorname{Spec}\nolimits A$. ∎
Remark 1.13.
Considering stable Jordan types (i.e., disregarding blocks of size $p$) can be useful, but has the disadvantage that the stable version of $JT_{A,M_{A}}(-)$ is not continuous.
Notation 1.14.
In the remainder of this paper, we shall leave implict the subscript $\leq$ on $\mathcal{Y}_{\leq,m}$ and $\mathcal{Y}_{\leq}$.
2. Jordan type functions for $\mathbb{G}$-modules
Consider an affine group scheme $\mathbb{G}$ of finite type over $k$. We are interested in formulating invariants for finite dimensional $\mathbb{G}_{(r)}$-modules. Here, $\mathbb{G}_{(r)}$ is the kernel of the $r$-th iterate of the Frobenius map $F:\mathbb{G}\to\mathbb{G}^{(1)}$, where $k[\mathbb{G}^{(1)}]$ is the base change of $k[\mathbb{G}]$ along the $p$-th power map $(-)^{p}:k\to k$. (See [18, §1].) Thus,
$\mathbb{G}_{(r)}\ \equiv\ ker\{F^{r}:\mathbb{G}\to\mathbb{G}^{(r)}\}$
is an infinitesimal group scheme of height $\leq r$ with coordinate algebra $k[\mathbb{G}_{(r)}]$. We denote by $k\mathbb{G}_{(r)}$ the $k$-linear dual of the finite dimensional Hopf algebra $k[\mathbb{G}_{(r)}]$.
For any commutative $k$-algebra $R$, we shall employ various notations (depending upon context) for
$\mathbb{G}_{R}\quad\equiv\quad\mathbb{G}\times\operatorname{Spec}\nolimits R\quad\equiv\quad R\otimes\mathbb{G}$
for the group scheme defined as the base change along $k\to R$ of the affine group scheme $\mathbb{G}$ over $k$. Thus, the coordinate algebra of $\mathbb{G}_{R}$ equals $R\otimes k[\mathbb{G}]$. If $\mathbb{G}$ is a finite group scheme over $k$, then the “group algebra” of $\mathbb{G}_{R}$ (namely, $Hom_{R}(R\otimes k[\mathbb{G}],R))$ is identified with $R\otimes k\mathbb{G}$ by sending $1\otimes\phi\in R\otimes k\mathbb{G}$ to $id_{R}\cdot\phi(-):R\otimes k[\mathbb{G}]\to R$.
Definition 2.1.
Let $\mathbb{G}$ be an affine group scheme of finite type over $k$ and let $r$ be a positive integer. As proved in [26, Thm1.5], the functor of commutative $k$-algebras sending an algebra $A$ to the set of maps of group schemes $\mu:\mathbb{G}_{a(r),A}\to\mathbb{G}_{A}$ over $A$ (i.e., the set of height $r$ 1-parameter subgroups of $\mathbb{G}_{A}$) is representable. The commutative $k$-algebra representing this functor is denoted $k[V_{r}(\mathbb{G})]$, its spectrum by $V_{r}(\mathbb{G})$.
Given a $\mathbb{G}_{(r)}$-module $M$ and a scheme theoretic point $x\in V_{r}(\mathbb{G})$, base change provides a $\mathbb{G}_{(r),k(x)}$-module structure on $k(x)\otimes M$. Namely, if $\mu_{x}:\mathbb{G}_{a(r),k(x)}\to\mathbb{G}_{(r),k(x)}$ is the 1-parameter subgroup over $k(x)$ parametrized by $\chi_{x}:\operatorname{Spec}\nolimits k(x)\to V_{r}(G)$, then restricting along $\mu_{x}$ determines a $\mathbb{G}_{a(r),k(x)}$-module structure on $k(x)\otimes M$.
In other words, restriction along the map of “group algebras” $\mu_{x,*}:k(x)\otimes k\mathbb{G}_{a(r)}\to k(x)\otimes k\mathbb{G}_{(r)}$ determines the $k(x)\otimes k\mathbb{G}_{a(r)}$-module structure on $k(x)\otimes M$.
The following definition makes explicit how the map of group algebras $\mu_{x,*}:k(x)\otimes k\mathbb{G}_{a(r)}\to k(x)\otimes k\mathbb{G}_{(r)}$ naturally determines by “restricting along $\epsilon_{r}:k[u]/u^{p}\to k\mathbb{G}_{a(r)}$” a map of $k(x)$-algebras $(\mu_{x})_{*}\circ\epsilon_{r}:k(x)\otimes k[u]/u^{p}\to\ k(x)\otimes k\mathbb{G}_{(r)}$. We remind the reader that the group algebra $k\mathbb{G}_{a(r)}$ (i.e., the dual of the coordinate algebra $k[\mathbb{G}_{a(r)}]=k[t]/t^{p^{r}}$) can be described as the commutative $k$-algebra $k[u_{0},\ldots,u_{r-1}]/(u_{i}^{p})$ where $u_{i}$ is the $k$-linear dual to $t^{p^{i}}$.
Definition 2.2.
Let $\mathbb{G}$ be an affine group scheme of finite type over $k$. For $r\geq 1$, we denote by $\epsilon_{r}:k[u]/u^{p}\to k\mathbb{G}_{a(r)}$ the map of $k$-algebras (but not of Hopf algebras if $r>1$) which sends $u$ to the element $u_{r-1}\in k\mathbb{G}_{a(r)}$. A scheme-theoretic point $x\in V_{r}(\mathbb{G})$ corresponding to the 1-parameter subgroup $\mu_{x}:\mathbb{G}_{a(r),k(x)}\to\mathbb{G}_{(r),k(x)}$ determines the map of $k(x)$-algebras
(2.2.1) $(\mu_{x})_{*}\circ\epsilon_{r}:k(x)\otimes k[u]/u^{p}\to k(x)\otimes k\mathbb{G}_{a(r)}\to k(x)\otimes k\mathbb{G}_{(r)},\ u\mapsto(\mu_{x})_{*}(1\otimes u_{r-1}).$
Thus, if $M$ is a $\mathbb{G}$-module, then $(\mu_{x})_{*}\circ\epsilon_{r})^{*}(k(x)\otimes M)$ is equipped with a $k(x)\otimes k[u]/u^{p}$-module structure such that $u$ acts as $(\mu_{x})_{*}(1\otimes u_{r-1})$ on $k(x)\otimes M$.
To unpack $(\mu_{x})_{*}(u_{r-1})\in k(x)\otimes k\mathbb{G}_{(r)}$, we observe that this is the $k(x)$-linear function sending $f\in k[\mathbb{G}]$ to the coefficient of $t^{p^{r-1}}$ in the expansion of $\mu_{x}^{*}(f)\in k(x)\otimes k[\mathbb{G}_{a(r)}]=k(x)[t]/t^{p^{r}}$.
We recall the local Jordan type of a finite dimensional $\mathbb{G}_{(r)}$-module as first formulated in [15] and [16] by Pevtsova and the author.
Definition 2.3.
(See [16, Cor 3.8].) Let $\mathbb{G}$ be an affine group scheme of finite type over $k$, $r\geq 1$, and $M$ a finite dimensional $\mathbb{G}_{(r)}$-module. Then the local Jordan type of $M$ at a scheme-theoretic point $x\in V_{r}(\mathbb{G})$, $JT_{\mathbb{G},r,M}(x)$, is defined to be the Jordan type of the $k(x)[u]/u^{p}$-module $((\mu_{x})_{*}\circ\epsilon_{r})^{*}(k(x)\otimes M)$.
In the following definition, we recall the natural grading on $V_{r}(\mathbb{G})$ as discussed in [26, §1]. This is particularly relevant because there is a natural morphism $\operatorname{Proj}\nolimits k[V_{r}(\mathbb{G})]\to\mathbb{P}H^{*}(\mathbb{G}_{(r)},k)\simeq\Pi(\mathbb{G}_{(r)})$ which is a homeomorphism of Zariski spaces to the $\pi$-point scheme $\Pi(\mathbb{G})$ (see [17, Prop 3.1]).
Definition 2.4.
The natural right action $V_{r}(\mathbb{G}_{a})\times\mathbb{A}^{1}\to V_{r}(\mathbb{G}_{a})$ sends
$(\nu:\mathbb{G}_{a(r),A}\to\mathbb{G}_{a(r),A},\alpha\in A)$ to $\alpha\cdot\nu:\mathbb{G}_{a(r),A}\to\mathbb{G}_{a(r),A}$ defined by $(\alpha\cdot\nu)^{*}(t)=\nu^{*}(\alpha\cdot t)$. Composing this with 1-parameter subgroups $\mu:\mathbb{G}_{a(r),A}\to\mathbb{G}_{(r),A}$ defines a right action $V_{r}(\mathbb{G})\times\mathbb{A}^{1}\to V_{r}(\mathbb{G})$. This action corresponds to a grading on the commutative algebra $k[V_{r}(\mathbb{G})]$.
We set
$\mathbb{P}V_{r}(\mathbb{G})\quad\equiv\quad\operatorname{Proj}\nolimits k[V_{r}(\mathbb{G})].$
If $\mu:\mathbb{G}_{(r),R}\to\mathbb{G}_{R}$ is an $R$-point of $V_{r}(\mathbb{G})$ for some commutative $k$-algebra $R$ and $s$ an $R$-point of $A^{1}$ (i.e., an element $s\in R$), then $(s\circ\mu)_{*}(u_{j})=s^{p^{i}}\mu_{*}(u_{j})$ for $u_{j}\in k\mathbb{G}_{a(r)}=k[u_{0},\cdots,u_{r-1}]/(u_{i}^{p}).$ (See the proof of [27, Prop 6.5]).)
We rephrase Lemma 1.12 of [26] in order to elaborate upon this grading of $V_{r}(\mathbb{G})$.
Proposition 2.5.
The commutative algebra $k[V_{r}(GL_{N})]$ is a graded quotient of the symmetric algebra over $k$ generated with generators $\{X_{i,j}^{\ell};1\leq i,j,\leq N,0\leq\ell<r\}$ provided that $X_{i,j}^{\ell}$ is given degree $p^{\ell}$. Moreover, if $\mathbb{G}\subset GL_{N}$ is a closed immersion, then $k[V_{r}(\mathbb{G})]$ is a graded quotient of $k[V_{r}(GL_{N})]$.
Consequently, for any affine group scheme $\mathbb{G}$ of finite type over $k$, $k[V_{r}(\mathbb{G})]$ is a graded commutative algebra generated by homogeneous elements of degrees $p^{i},\ 0\leq i<r$.
The following theorem summarizes fundamental results of Suslin, Bendel, and the author which relate 1-parameter subgroups of $\mathbb{G}$ of height $r$ to cohomological support varieties of $\mathbb{G}_{(r)}$.
Theorem 2.6.
[26, Thm 5.2], [27, Thm 6.7], [27, Cor 6.8] Let $\mathbb{G}$ be an affine group scheme of finite type over $k$, $r$ a positive integer, and $M$ a finite dimensional $\mathbb{G}_{(r)}$-module. There is a natural morphism of commutative graded $k$-algebras $\psi:H^{ev}(\mathbb{G}_{(r)},k)\ \to\ k[V_{r}(\mathbb{G})]$ of degree $\frac{p^{r}}{2}$ with associated morphism $\Psi:V_{r}(\mathbb{G})\ \to\ \operatorname{Spec}\nolimits H^{ev}(\mathbb{G}_{(r)},k)$ of schemes which is a homeomorphism on scheme-theoretic points with the Zariski topology. This morphism restricts to a homeomorphism of closed subspaces
$\Psi:V_{r}(\mathbb{G})_{M}\quad\stackrel{{\scriptstyle\approx}}{{\to}}\quad\operatorname{Spec}\nolimits H^{ev}(\mathbb{G}_{(r)},k)_{M},$
where
$V_{r}(\mathbb{G})_{M}\quad=\quad\{x\in V_{r}(\mathbb{G}):\ J_{\mathbb{G},r,M}(x)\not=\frac{dimM}{p}\cdot[p]\}$
and $\operatorname{Spec}\nolimits H^{ev}(\mathbb{G}_{(r)},k)_{M}$ is the (affine) cohomological support variety of $M$.
Moreover, $\psi$ determines the “projectivized morphism”
$\Psi:\operatorname{Proj}\nolimits V_{r}(\mathbb{G})\ \to\ \operatorname{Proj}\nolimits H^{ev}(\mathbb{G}_{(r)},k)$ which restricts to a homeomorphism
$\Psi:\operatorname{Proj}\nolimits V_{r}(\mathbb{G})_{M}\ \stackrel{{\scriptstyle\approx}}{{\to}}\ \operatorname{Proj}\nolimits H^{ev}(\mathbb{G}_{(r)},k)_{M}.$
Remark 2.7.
We refer the reader to [23] for some explorations of $V_{r}(\mathbb{G})$, including the reducibility and lack of equi-dimensionality in general, and some bounds on dimensions.
Remark 2.8.
The map $(\mu_{x})_{*}\circ\epsilon_{r}$ of (2.2.1) is a $\pi$-point of $\mathbb{G}_{(r)}$ whose equivalence class $(\mu_{x})_{*}\circ\epsilon_{r}]$ is a scheme-theoretic point of the $\pi$-point space $\Pi(\mathbb{G}_{(r)})$. This determines the map of points induced by the homeomorphism
(2.8.1) $\mathbb{P}V_{r}(\mathbb{G})\quad\stackrel{{\scriptstyle\approx}}{{\to}}\quad\mathbb{P}H^{*}(\mathbb{G}_{(r)},k)\quad\stackrel{{\scriptstyle\sim}}{{\to}}\quad\Pi(\mathbb{G}_{(r)})$
obtained by composing $\Psi$ of Theorem 2.6 with the natural isomorphism $\mathbb{P}H^{*}(\mathbb{G}_{(r)},k)\stackrel{{\scriptstyle\sim}}{{\to}}\Pi(\mathbb{G}_{(r)})$ of [14, Thm 7.5].
Observe that sending the 1-parameter subgroup $\mu_{x}:\mathbb{G}_{a(r),k(x)}\to\mathbb{G}_{k(x)}$ to the equivalence class of the $\pi$-point $(\mu_{x})_{*}\circ\epsilon_{r}:k(x)\otimes k[u]/u^{p}\to k(x)\otimes k\mathbb{G}_{a(r)}\to k(x)\otimes k\mathbb{G}$ determines a natural projection
(2.8.2) $V_{r}(\mathbb{G}){\text{-}}\{0\}\quad\to\quad\mathbb{P}V_{r}(\mathbb{G}).$
For any finite dimensional $\mathbb{G}_{(r)}$-module $M$, the maps of (2.8.1) restrict to homeomorphisms of corresponding support varieties for $M$
$\mathbb{P}V_{r}(\mathbb{G})_{M}\ \stackrel{{\scriptstyle\approx}}{{\to}}\ \mathbb{P}H^{*}(\mathbb{G}_{(r)},k)_{M}\ \stackrel{{\scriptstyle\sim}}{{\to}}\ \Pi(\mathbb{G}_{(r)})_{M}.$
Remark 2.9.
The more general theory of $\pi$-points supports developed by Pevtsova and the author in [14], $M\to\Pi(H)_{M}$, applies to $H$-modules for an arbitrary finite group scheme $H$. A disadvantage of considering $M\to\Pi(H)_{M}$ rather than $M\mapsto V_{r}(\mathbb{G})_{M}$ is that points of $\Pi(H)$ are equivalence classes of certain flat maps of $K$-algebras of the form $\alpha_{K}:K[u]/u^{p}\to KH$ for field extensions $K/k$. For maximal or generic points of $\Pi(H)$, the Jordan type of a finite dimensional $H$-module $M$ is independent of the choice of representative of such an an equivalence class $\alpha_{K}]$, but for other points of $\Pi(H)$ the Jordan type of $M$ depends upon a choice of representative of the equivalence class. See [17]).
To emphasize the subtlety of the geometry of $\mathbb{P}V_{r}(\mathbb{G})$, we analyze the very special example of $\mathbb{P}V_{r}(\mathbb{G}_{a})$.
Example 2.10.
We identify the graded coordinate algebra of $V_{r}(\mathbb{G}_{a})=V_{r}(\mathbb{G}_{a(r)})$ with the graded polynomial algebra $k[T_{0},T_{1},\ldots,T_{r-1}]$, where $T_{i}$ is given degree $p^{i}$. Then $\operatorname{Proj}\nolimits$ applied to the graded algebra is the weighted projective space $w\mathbb{P}(1,p,\ldots,p^{r-1})$. Let $k[t_{0},t_{1},\ldots,t_{r-1}]$ denote the graded polynomial algebra with each $t_{i}$ of degree 1. The graded map of $k$-algebras
$\phi^{*}:k[T_{0},T_{1},\ldots,T_{r-1}]\ \to\ k[t_{0},t_{1},\ldots,t_{r-1}],\quad T_{i}\mapsto t_{i}^{p^{i}}$
determines a ramified covering map
$\Phi:\mathbb{P}^{r-1}\quad\to\quad w\mathbb{P}(1,p,\ldots,p^{r-1}).$
Similarly, the map of $k$-algebras
$\psi^{*}:k[t_{0},t_{1},\ldots,t_{r-1}]\ \to\ k[T_{0},T_{1},\ldots,T_{r-1}],\quad t_{i}\mapsto T_{i}^{p^{r-i-1}}$
multiplies all degrees by $p^{r-1}$ and thus induces a morphism
$\Psi:w\mathbb{P}(1,p,\ldots,p^{r-1})\quad\to\quad\mathbb{P}^{r-1}.$
Moreover,
$\Psi\circ\Phi=F^{r-1}:\mathbb{P}^{r-1}\ \to\mathbb{P}^{r-1},\quad\Phi\circ\Psi=F^{r-1}:w\mathbb{P}(1,p,\ldots,p^{r-1})\to w\mathbb{P}(1,p,\ldots,p^{r-1}).$
This will enable us to compare Jordan type functions for $(\mathbb{G}_{a}^{\times r})_{1}$ and $\mathbb{G}_{a(r)}$ in Example 5.2.
Example 2.11.
Consider the case $r=3$, so that $\mathbb{P}V_{3}(\mathbb{G}_{a})\ =\ w\mathbb{P}(1,p,p^{2})$ is 2-dimensional, covered by three affine open subsets:
$U_{0}\ =\ \operatorname{Spec}\nolimits k[T_{1}/T_{0}^{p},T_{2}/T_{0}^{p^{2}}]\ \simeq\ \mathbb{A}^{2},$
$U_{1}\ =\ \operatorname{Spec}\nolimits k[T_{0}^{p}/T_{1},T_{2}/T_{1}^{p}]\ \simeq\ \mathbb{A}^{2},$
$U_{2}\ =\ \operatorname{Spec}\nolimits k[T_{0}^{ip}T_{1}^{p-i}/T_{2};\ 0\leq i\leq p].$
Observe that $U_{2}\ \subset\ w\mathbb{P}(1,p,p^{2})$ has an isolated singularity at $T_{0}=0=T_{1}$; the maximal ideal at this singular point is generated by the $p+1$ elements
$T_{0}^{p^{2}}/T_{2},T_{0}^{p^{2}-p}T_{1}/T_{2},\ldots,T_{0}^{p}T_{1}^{p-1}/T_{2},T_{1}^{p}/T_{2}$.
The representability by $k[V_{r}(\mathbb{G})]$ of the functor sending a commutative $k$-algebra AA to the set of all height $r$ 1-parameter subgroups $\mathbb{G}_{a(r),A}\to\mathbb{G}_{A}$ (over $\operatorname{Spec}\nolimits A$) leads to the universal 1-parameter subgroup $\mathcal{U}_{\mathbb{G},r}$ as recalled in the following definition.
Definition 2.12.
([16, §2]) Let $\mathbb{G}$ be an affine group scheme of finite type over $k$ and let $r$ be a chosen positive integer. Then the universal 1-parameter subgroup of height $r$ for $\mathbb{G}$,
($2.1$) $\mathcal{U}_{\mathbb{G},r}:\mathbb{G}_{a(r),k[V_{r}(\mathbb{G})]}\ \to\mathbb{G}_{(r),k[V_{r}(\mathbb{G})]},$
is the 1-parameter subgroup for the group scheme $\mathbb{G}_{(r),k[V_{r}(\mathbb{G})]}$ over $k[V_{r}(\mathbb{G})]$ corresponding to the identity map $V_{r}(\mathbb{G})\to V_{r}(\mathbb{G})$. By the universality of $\mathcal{U}_{\mathbb{G},r}$, for any scheme-theoretic point $x\in V_{r}(\mathbb{G})$ the restriction along $\chi_{x}:\operatorname{Spec}\nolimits k(x)\to V_{r}(\mathbb{G})$ of $\mathcal{U}_{\mathbb{G},r}$ equals the map $\mu_{x}:\mathbb{G}_{a(r),k(x)}\ \to\mathbb{G}_{(r),k(x)}$ of Definition 2.1.
We denote by
(2.12.2) $(\mathcal{U}_{\mathbb{G},r})_{*}:k[V_{r}(\mathbb{G})]\otimes k\mathbb{G}_{a(r)}\quad\to\quad k[V_{r}(\mathbb{G})]\otimes k\mathbb{G}_{(r)}$
the map of group algebras over $k[V_{r}(\mathbb{G})]$ induced by $\mathcal{U}_{\mathbb{G},r}$. In other words, $(\mathcal{U}_{\mathbb{G},r})_{*}$ is the $k[V_{r}(\mathbb{G})]$-dual of the map on coordinate algebras $\mathcal{U}_{\mathbb{G},r}^{*}:k[V_{r}(\mathbb{G})]\otimes k[\mathbb{G}_{(r)}]\to k[V_{r}(\mathbb{G})]\otimes k[\mathbb{G}_{a(r)}]$ given by $\mathcal{U}_{\mathbb{G},r}$.
By [16, Prop 2.9], a closed embedding $i:\mathbb{H}\hookrightarrow\mathbb{G}$ of affine group schemes of finite type over $k$ determines the commutative square
(2.12.3) ${\mathbb{G}_{a(r),k[V_{r}(\mathbb{G})]}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{(\mathcal{U}_{\mathbb{G},r})_{*}}$${\mathbb{G}_{(r),k[V_{r}(\mathbb{G})]}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathbb{G}_{a(r),k[V_{r}(\mathbb{H})]}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{(\mathcal{U}_{\mathbb{H},r})_{*}}$${\mathbb{H}_{(r),k[V_{r}(\mathbb{H})]}].}$
We recall and reformulate the “universal $p$-nilpotent operator” $\Theta_{\mathbb{G},r}$ for $\mathbb{G}_{(r)}$-modules introduced in [16, Defn 2.1].
Definition 2.13.
Let $\mathbb{G}$ be an affine group scheme of finite type over $k$ and let $r$ be a chosen positive integer. We define
(2.13.1) $\Theta_{\mathbb{G},r}\quad\equiv\quad(\mathcal{U}_{\mathbb{G},r})_{*}(1\otimes u_{r-1})\quad\in\quad k[V_{r}(\mathbb{G})]\otimes k\mathbb{G}_{(r)}\ \equiv\ k\mathbb{G}_{(r),k[V_{r}(\mathbb{G})]},$
where $(\mathcal{U}_{\mathbb{G},r})_{*}$ is defined in (2.12.2).
In [16, Ex 2.6], the reader will find a few explicit computations of $\Theta_{\mathbb{G},r}$. We mention two such examples consistent with the indexing used in this paper.
Example 2.14.
Consider $\mathbb{G}=\mathbb{G}_{a}^{\times s}$ whose Lie algebra is the dimension $s$ Lie algebra $\mathfrak{g}_{a}^{\oplus s}$ (which is commutative, with $p$-th power map trivial). Identify $V_{1}(\mathbb{G})$ with $\mathbb{A}^{r}=\operatorname{Spec}\nolimits k[t_{0},\ldots,t_{s-1}]$ and identify the group algebra $k\mathbb{G}_{(1)}$ with $k[x_{0},\ldots,x_{s-1}]/(x_{i}^{p})\\ =\ \mathfrak{u}(\mathfrak{g}_{a}^{\oplus s})$, the restricted enveloping algebra of the restricted Lie algebra $\mathfrak{g}_{a}^{\oplus s}$. Then
$\Theta_{\mathbb{G},1}\ =\ \sum_{i=0}^{r-1}t_{i}\otimes x_{i}\ \in\ k[V_{1}(\mathbb{G}_{a}]\otimes\mathfrak{u}(\mathfrak{g}_{a}^{\oplus s}).$
Consider $\mathbb{G}=\mathbb{G}_{a}$ whose Lie algebra is $\mathfrak{g}_{a}$ and some $r>0$. Identify $V_{r}(\mathbb{G}_{a})$ with $\operatorname{Spec}\nolimits k[T_{0},\ldots,T_{r-1}]$ with $T^{i}$ given grading $p^{i}$ and identify the group algebra $k\mathbb{G}_{(r)}$ with $k[u_{0},\ldots,u_{r-1}]/(u_{i}^{p})$. Then
$\Theta_{\mathbb{G}.r}\ =\ \sum_{i=0}^{r-1}T_{i}^{p^{r-1-i}}\otimes u_{i}\ +\ \cdots\quad\in\ k[V_{r}(\mathbb{G}_{a})]\otimes k\mathbb{G}_{a(r)},$
where the non-explicit, remaining terms of $\Theta_{\mathbb{G}_{a},r}$ each consist of a polynomial homogeneous of weighted degree $p^{r-1}$ in the $T_{i}$’s tensor a monomial in the $u_{i}$’s of degree at least 2. (See [27, Eqn 6.5.1].)
The somewhat subtle aspect of Theorem 2.15 is its last assertion of homogeneity of degree $p^{r-1}$ whose proof is given in [16] (reflected in the examples above).
Theorem 2.15.
([16]) Let $\mathbb{G}$ be an affine group scheme of finite type over $k$ and let $r$ be a chosen positive integer. Then $\Theta_{\mathbb{G},r}$ as given in Definition 2.13 satisfies the following properties.
  1. (1)
    For any scheme-theoretic point $x\in V_{r}(\mathbb{G})$, the specialization along $\chi_{x}:\operatorname{Spec}\nolimits k(x)\to V_{r}(\mathbb{G})$ of $\Theta_{\mathbb{G},r}\in k[V_{r}(\mathbb{G})]\otimes k\mathbb{G}_{(r)}$ equals $(\mu_{x})_{*}(1\otimes u_{r-1})\quad\in\ k(x)\otimes k\mathbb{G}_{(r)}$, where $\mu_{x}$ is the 1-parameter subgroup of Definition 2.1.
  2. (2)
    Let $\phi:\mathbb{H}\hookrightarrow\mathbb{G}$ be a closed subgroup scheme determining $\phi^{*}:k[V_{r}(\mathbb{G})]\twoheadrightarrow k[V_{r}(\mathbb{H})]$ and $\phi_{(r)*}:k\mathbb{H}_{(r)}\hookrightarrow k\mathbb{G}_{(r)}$. Then the restriction of $\Theta_{\mathbb{G},r}$ along $\phi^{*}$, $(\phi^{*}\otimes 1)(\Theta_{\mathbb{G},r})\in k[V_{r}(\mathbb{H})]\otimes k\mathbb{G}_{(r)}$, equals the image of $(1\otimes\phi_{(r)*})(\Theta_{\mathbb{H},r})$.
  3. (3)
    For any $\mathbb{G}_{(r)}$-module $M$, $\Theta_{\mathbb{G},r}$ determines a $p$-nilpotent, $k[V_{r}(\mathbb{G})]$-linear operator $\Theta_{\mathbb{G},r,M}:k[V_{r}(\mathbb{G})]\otimes M\ \to\ k[V_{r}(\mathbb{G})]\otimes M$.
  4. (4)
    The base change along a scheme theoretic point $x\in V_{r}(\mathbb{G})$ of $\Theta_{\mathbb{G},r,M}$ equals the action of $(\mu_{x})_{*}(1\otimes u_{r-1})$ on $k(x)\otimes M$.
  5. (5)
    $\Theta_{\mathbb{G},r}$ is of the form $\sum f_{i}\otimes\phi_{i}\in k[V_{r}(\mathbb{G})]\otimes k\mathbb{G}_{(r)}$ with each $f_{i}$ homogeneous of degree $p^{r-1}$ in $k[V_{r}(\mathbb{G})]$ and each $\phi_{i}\in k\mathbb{G}_{(r)}$ an element of the augmentation ideal $I_{\mathbb{G}_{(r)}}$ of $k\mathbb{G}_{(r)}$.
Proof.
Since $\mu_{x}$ is the pull-back along $\chi_{x}$ of $\mathcal{U}_{\mathbb{G},r}$, the following diagram commutes:
(2.15.1) ${k[V_{r}(\mathbb{G})]\otimes k\mathbb{G}_{a(r)}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\chi^{*}}$$\scriptstyle{(\mathcal{U}_{\mathbb{G},r})_{*}}$${k[V_{r}(\mathbb{G})]\otimes k\mathbb{G}_{(r)}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\chi^{*}}$${k(x)\otimes k\mathbb{G}_{a(r)}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{(\mu_{x})_{*}}$${k(x)\otimes k\mathbb{G}_{(r)}.}$
Consequently, the restriction of $\Theta_{\mathbb{G},r}=(\mathcal{U}_{\mathbb{G},r})_{*}(1\otimes u_{r-1})$ equals $(\mu_{x})_{*}(1\otimes u_{r-1})$.
The naturality of $\mathbb{G}\mapsto\Theta_{\mathbb{G},r}$ with respect to closed embeddings $\phi:\mathbb{H}\hookrightarrow\mathbb{G}$ as in assertion (2) follows from the commutativity of (2.12.3).
The action by $\Theta_{\mathbb{G},r,M}$ is given by the restriction to $\{\Theta_{\mathbb{G},r}\}\otimes(k[V_{r}(\mathbb{G})]\otimes M)$ of the tensor product pairing
(2.15.2) $(k[V_{r}(\mathbb{G})]\otimes k\mathbb{G}_{(r)})\otimes(k[V_{r}(\mathbb{G})]\otimes M)\quad\to\quad k[V_{r}(\mathbb{G})]\otimes M.$
The fact that this action is $p$-nilpotent follows from the observation that the $i$-th power of $\Theta_{\mathbb{G},r}$ equals $(\mathcal{U}_{\mathbb{G},r})_{*}((1\otimes u_{r-1})^{i})$. Assertion (4) follows directly from assertion (1).
The last assertion is stated and proved in [16, Prop 2.11]. The proof proceeds by first observing that it suffices to verify that the restriction of the $k[V_{r}(\mathbb{G})]$-linear map $\mathcal{U}_{\mathbb{G},r}^{*}:k[V_{r}(\mathbb{G})]\otimes k[\mathbb{G}_{(r)}]\to k[V_{r}(\mathbb{G})]\otimes k[\mathbb{G}_{a(r)}]$ to $1\otimes k[\mathbb{G}_{(r)}]$ reads off the coefficient of $t^{p^{r-1}}$ in $k[V_{r}(\mathbb{G})]\otimes k\mathbb{G}_{a(r)}\simeq k[V_{r}(\mathbb{G})]\otimes k[t]/t^{p^{r}}$. The authors then verify that the actions of $\mathbb{A}^{1}$ on $\mathbb{G}_{a(r)}$ and $V_{r}(\mathbb{G})$ are appropriately compatible. ∎
Definition 2.16.
Let $\mathbb{G}$ be an affine group scheme of finite type over $k$ and let $r$ be a positive integer. For any finite dimensional $\mathbb{G}_{(r)}$-module $M$, we define the function
(2.16.1) $JT_{\mathbb{G},r,M}(-):V_{r}(\mathbb{G})\quad\to\quad\mathcal{Y}$
from the set of scheme-theoretic points of $V_{r}(\mathbb{G})$ to the set of points of the partially orders set $\mathcal{Y}$ by sending $x\in V_{r}(\mathbb{G})$ to
$JT_{\mathbb{G},r,M}(x)\quad\equiv\quad JT((\mu_{x})_{*}\circ\epsilon_{r})^{*}(k[V_{r}(\mathbb{G})]\otimes M)),$
the Jordan type of the $k(x)[u]/u^{p}$-module obtained by restricting $k[V_{r}(\mathbb{G})]\otimes M$ along $(\mu_{x})_{*}\circ\epsilon_{r}:k(x)[u]/u^{p}\to k(x)\otimes k\mathbb{G}_{a(r)}\to k[V_{r}(\mathbb{G})]\otimes k\mathbb{G}_{(r)}$; equivalently, $JT_{\mathbb{G},r,M}(x)$ is the Jordan type of $(\mu_{x})_{*}(1\otimes u_{r-1})$ acting on $k(x)\otimes M$.
By Theorem 2.15(1), $JT_{\mathbb{G},r,M}(x)$ equals the Jordan type of $k(x)\otimes_{k[V_{r}(\mathbb{G})]}\Theta_{\mathbb{G},r,M}\ \in\ End_{k(x)}(k(x)\otimes M)$.
Let $R$ be a graded, commutative $k$-algebra; for example, $R\ =\ k[V_{r}(\mathbb{G})]$. For any graded $R$-module $M_{*}$, we denote by $(M_{*})^{\sim}$ the associated quasi-coherent sheaf on $\operatorname{Proj}\nolimits(R)$. We shall often consider graded $R$-modules of the form $R\otimes V$ where $V$ is a $k$-vector space and $R\otimes V$ inherits the grading of $R$.
If $M_{*}$ is a graded $R$-module, we denote by $M_{*}(n)$ the graded module obtained from $M_{*}$ by “shifting down by nn” the degrees of $M_{*}$; thus, $(M_{*}(n))_{d}=M_{n+d}$. We use $(M_{*})^{\sim}(n)$ to denote $(M_{*}(n))^{\sim}$. The operator $\Theta_{\mathbb{G},r,M}$ of Theorem 2.15(3) multiplies degrees by $p^{r-1}$, so that viewed as an operator on graded modules one should write $\Theta_{\mathbb{G},r,M}$ as
$\Theta_{\mathbb{G},r,M}:k[V_{r}(\mathbb{G})]\otimes M\quad\to\quad(k[V_{r}(\mathbb{G})]\otimes M)(p^{r-1}).$
Because $k$ has positive characteristic, the informative survey by M. Beltrametti and L. Robbiano in [2] does not apply to our consideration of $\mathbb{P}V_{r}(\mathbb{G})$. Because $V_{r}(\mathbb{G})$ is usually not generated by homogeneous elements of degree 1, we supplement the clear exposition of R. Hartshorne in [19, §II.5] with [10, Ex 9.5] when considering the relationship between graded modules over a commutative, graded $k$-algebra $R$ and sheaves of modules on $\operatorname{Proj}\nolimits(R)$.
Proposition 2.17.
Consider a polynomial algebra $k[T_{0},\ldots,T_{N}]$ with each $T_{i}$ given a “weighted” degree $d_{i}$. Assume that $I_{Y}\subset k[T_{0},\ldots,T_{N}]$ is an ideal generated by polynomials $F_{j}(\underline{T})$ which are (weighted) homogeneous. Then, the zero locus $Y=Z(I_{Y})\ \subset\operatorname{Proj}\nolimits k[T_{0},\ldots,T_{N}]$ is a weighted projective variety associated to the commutative, graded algebra $k[T_{0},\ldots,T_{N}]/I_{Y}$.
Let $d$ be a positive integer divisible by each $d_{i}$. The graded $k[T_{0},\ldots,T_{N}]/I_{Y}$-module $(k[T_{0},\ldots,T_{N}]/I_{Y})(d)$ is a locally free, coherent rank 1 sheaf of $\mathcal{O}_{Y}$-modules on $Y$. We denote this sheaf by $\mathcal{O}_{Y}(d)$. For each $i$ such that $T_{i}$ is not in the radical of $I_{Y}$, the restriction $(\mathcal{O}_{Y}(d))_{|V_{i}}$ of $\mathcal{O}_{Y}(d)$ to open subvariety
$V_{i}\ \equiv(\operatorname{Proj}\nolimits k[T_{0},\ldots,T_{N}])-Z(T_{i}))\cap Y\ \subset\operatorname{Proj}\nolimits k[T_{0},\ldots,T_{N}]$
is generated as a free, rank 1 $\mathcal{O}_{V_{i}}$-module by the image of $T_{i}^{d/d_{i}}\in\mathcal{O}_{Y}(d)(Y)$ in $\mathcal{O}_{Y}(d)(V_{i})$.
Proof.
It suffices to assume that $I_{Y}$ is a radical ideal . Moreover, by considering each irreducible component of $Y$, we may assume that $Y$ is irreducible. Then [10, Ex 9.5] applies to verify that the variety $Z(I_{Y})$ is isomorphic to its image in $\operatorname{Proj}\nolimits S$, where $S\ =k[T_{0},\ldots,T_{N}]_{(d)}\subset k[T_{0},\ldots,T_{N}]$ is the graded subalgebra generated by (weighted) homogeneous polynomials each of whose weights is divisible by $d$ and the image of $Z(Y)$ is given by the homogeneous ideal $I_{Y}\cap S\subset S$ which is generated by elements of degree $d$. Thus, replacing $k[T_{0},\ldots,T_{N}]$ by $S$ with grading modified by dividing degrees by $d$, we may apply [19, Prop II.5.12] (which is stated for graded submodules of a commutative, graded algebra generated in degree 1). ∎
In particular, Proposition 2.17 applies to the graded commutative algebra $k[V_{r}(\mathbb{G})]$ which is generated by elements homogeneous of degree $p^{i},1\leq i<r$. (See Proposition 2.5.)
We next present the projectivization of the universal operator $\Theta_{\mathbb{G},r}$. This operator has been used in [16] to produce global invariants on $\mathbb{P}V_{r}(\mathbb{G})$ for $\mathbb{G}$-modules (especially, non-trivial vector bundles on $\mathbb{P}V_{r}(\mathbb{G})$) in contrast to the Jordan type function $JT_{\mathbb{G},r,M}(-)$ which is a “point-wise invariant” on $\mathbb{P}V_{r}(\mathbb{G})$ invariant for a $\mathbb{G}$-module $M$. See Section 6 below.
Theorem 2.18.
Let $\mathbb{G}$ be an affine group scheme of finite type over $k$ and let $r$ be a chosen positive integer. The universal operator $\Theta_{\mathbb{G},r}$ of Theorem 2.15 determines a global section
$\tilde{\Theta}_{\mathbb{G},r}\quad\in\quad\Gamma(\mathbb{P}V_{r}(\mathbb{G}),\mathcal{O}_{\mathbb{P}V_{r}(\mathbb{G})}(p^{r-1})\otimes k\mathbb{G}_{(r)})$
of the coherent, locally free sheaf $\mathcal{O}_{\mathbb{P}V_{r}(\mathbb{G})}(p^{r-1})\otimes k\mathbb{G}_{(r)})$ on $\mathbb{P}V_{r}(\mathbb{G})$ satisfying the following properties.
  1. (1)
    For any $\mathbb{G}_{(r)}$-module $M$, tensor product with $\tilde{\Theta}_{\mathbb{G},r}$ determines the map of $\mathcal{O}_{\mathbb{P}V_{r}(\mathbb{G})}$-modules
    $\tilde{\Theta}_{\mathbb{G},r,M}:(k[V_{r}(\mathbb{G})]\otimes M)^{\sim}\quad\to\quad(k[V_{r}(\mathbb{G})]\otimes M)^{\sim}(p^{r-1}).$
  2. (2)
    For any non-zero divisior $h\in k[V_{r}(\mathbb{G})]_{p^{r-1}}$, let $U_{h}$ denote the affine open subset $Spec(k[V_{r}(\mathbb{G})][h^{-1}])_{0}\ \subset\mathbb{P}V_{r}(\mathbb{G})$. The natural isomorphism of $(\mathcal{O}_{\mathbb{P}V_{r}(\mathbb{G})})_{U_{h}}$-modules
    $h^{-1}\otimes 1:(\mathcal{O}_{\mathbb{P}V_{r}(\mathbb{G})}(p^{r-1})\otimes k\mathbb{G}_{(r)})_{|U_{h}}\quad\stackrel{{\scriptstyle\sim}}{{\to}}\quad(\mathcal{O}_{\mathbb{P}V_{r}(\mathbb{G})}\otimes k\mathbb{G}_{(r)})_{|U_{h}}$
    sends
    $(\tilde{\Theta}_{\mathbb{G},r})_{|U_{h}}\ \in\ \Gamma(U_{h},\mathcal{O}_{\mathbb{P}V_{r}(\mathbb{G})}(p^{r-1})\otimes k\mathbb{G}_{(r)})\ =\ (k[V_{r}(\mathbb{G})][1/h])_{p^{r-1}}\otimes k\mathbb{G}_{(r)}$
    to
    $(h^{-1}\otimes 1)\cdot\Theta_{\mathbb{G},r}\in k[V_{r}(\mathbb{G})][1/h]\otimes k\mathbb{G}_{(r)}.$
  3. (3)
    If $\overline{x}$ is a scheme-theoretic point of $\mathbb{P}V_{r}(\mathbb{G})$ satisfying the condition that $h(\overline{x})\not=0$ and if $x$ is a scheme theoretic point of $V_{r}(\mathbb{G}){\text{-}}\{0\}$ projecting to $\overline{x}$, then the restriction $k(x)\otimes_{\mathcal{O}_{\mathbb{P}V_{r}(\mathbb{G})}}\tilde{\Theta}_{\mathbb{G},r}$ along $\chi_{\overline{x}}:\operatorname{Spec}\nolimits k(x)\to\mathbb{P}V_{r}(\mathbb{G})$ of $\tilde{\Theta}_{\mathbb{G},r}$ is given by
    (2.18.1) $k(x)\otimes_{\mathcal{O}_{\mathbb{P}V_{r}(\mathbb{G})}}\tilde{\Theta}_{\mathbb{G},r}\quad=\quad c\cdot(\mu_{x})_{*}(1\otimes u_{r-1})\in k(x)\otimes k\mathbb{G}_{(r)}$
    for some $0\not=c\in k(x)$.
  4. (4)
    For any $\mathbb{G}_{(r)}$-module $M$, the restriction of $\tilde{\Theta}_{\mathbb{G},r,M}$ along $\chi_{\overline{x}}:\operatorname{Spec}\nolimits k(x)\to\mathbb{P}V_{r}(\mathbb{G})$ is a non-zero scalar multiple of the image of $(\mu_{x})_{*}(1\otimes u_{r-1})$ in
    $End_{k(x)}(k(x)\otimes M$).
Proof.
The global sections of the sheaf $\mathcal{O}_{\mathbb{P}V_{r}(\mathbb{G})}(p^{r-1})\otimes k\mathbb{G}_{(r)}$ on $\mathbb{P}V_{r}(\mathbb{G})$ are those elements of degree 0 of the graded $k[V_{r}(\mathbb{G})]$-module $k[V_{r}(\mathbb{G})](p^{r-1})\otimes k\mathbb{G}_{(r)}$. Since $\Theta_{\mathbb{G},r}\in k[V_{r}(\mathbb{G})]\otimes k\mathbb{G}_{(r)}$ is homogeneous of degree $p^{r-1}$, it determines the global section $\Theta_{\mathbb{G},r}$ of $\mathcal{O}_{\mathbb{P}V_{r}(\mathbb{G})}(p^{r-1})\otimes k\mathbb{G}_{(r)}$.
The first assertion of the theorem follows from Theorem 2.15(3) and the correspondence between graded $k[V_{r}(\mathbb{G})]$-modules and quasi-coherent sheaves on $\mathbb{P}V_{r}(\mathbb{G})$. The second assertion follows from the standard description of principal open subsets of $\operatorname{Proj}\nolimits R$ associated to homogenous elements $h\in R$ as $\operatorname{Spec}\nolimits R[h^{-1}]$.
The equality (2.18.1) follows from Theorem 2.15(1) and the identification of assertion (2). Assertion (4) follows immediately from assertion (3) (see Theorem 2.15(4)). ∎
Remark 2.19.
For any $\mathbb{G}_{(r)}$-module $M$ of dimension $m$ with $k$-linear dual $M^{*}$,
$JT_{\mathbb{G},r,M}(-)\quad=\quad JT_{\mathbb{G},r,M^{*}}(-):\mathbb{P}V_{r}(\mathbb{G})\to\mathcal{Y}_{m}.$
Moreover, if $I$ and $J$ restrict to injective $\mathbb{G}_{(r)}$ modules of the same dimension, then $JT_{\mathbb{G},r,I}(-)=JT_{\mathbb{G},r,J}(-)$. Thus, $M\mapsto JT_{\mathbb{G},r,M}(-)$ is far from faithful.
The next proposition verifies that $JT_{\mathbb{G},r,M}$ is a “function on $\mathbb{G}$-conjugacy classes”. Fix the convention that the action $(-)^{g}:G\to G$ for a group $G$ with $g\in G$ sends $h\in G$ to $g^{-1}\cdot h\cdot g\in G$. This easily leads to the adjoint action of $\mathbb{G}$ on itself (and on its normal subgroups $\mathbb{G}_{(r)}$). Composition of a 1-parameter group $\mathbb{G}_{a(r)}\to\mathbb{G}$ with this adjoint action determines an “adjoint action” of $\mathbb{G}$ on $V_{r}(\mathbb{G})$.
Proposition 2.20.
Consider an affine group scheme $\mathbb{G}$ of finite type over $k$, $r$ a positive integer, $M$ a $\mathbb{G}$-module of dimension $m$, and $\chi_{x}:\operatorname{Spec}\nolimits k(x)\to V_{r}(\mathbb{G})$ a scheme-theoretic point of $V_{r}(\mathbb{G})$ corresponding to the 1-parameter subgroup $\mu_{x}:\mathbb{G}_{a(r),k(x)}\to\mathbb{G}_{k(x)}$. For $g\in\mathbb{G}(k(x))$, denote by $\chi_{x^{g}}:\operatorname{Spec}\nolimits k(x)\to V_{r}(\mathbb{G})$ the the scheme-theoretic point corresponding to the 1-parameter subgroup $(-)^{g}\circ\mu_{x}:\mathbb{G}_{a(r),k(x)}\to\mathbb{G}_{k(x)}$. Then
$JT_{\mathbb{G},r,M}(x^{g})\quad=\quad JT_{\mathbb{G},r,M}(x).$
Proof.
Using base change from $k$ to $k(x)$, we reduce to the notationally simpler case that $k=k(x)$. Let $M^{g}$ denote the $\mathbb{G}$-module with the same underlying vector space structure as that of $M$ and with $h\in\mathbb{G}(k)$ sending $n\in M^{g}$ to $(g^{-1}hg)\cdot n\in M^{g}$ where $\cdot$ refers to the given action of $\mathbb{G}(k)$ on $M$. Observe that the map $g:M^{g}\to M,\ n\mapsto g\cdot n$ is $\mathbb{G}(k)$-equivariant.
This leads to a commutative diagram (viewing group schemes as functors from commutative $k$-algebras to groups)
(2.20.1) ${k[V_{r}(\mathbb{G})]\otimes M^{g}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{1\otimes g\bullet}$$\scriptstyle{\Theta_{\mathbb{G},r,M^{g}}}$${k[V_{r}(\mathbb{G})]\otimes M\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\Theta_{\mathbb{G},r,M}}$${k[V_{r}(\mathbb{G})]\otimes M^{g}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{1\otimes g\bullet}$$\scriptstyle{(\mu_{x^{g}})^{*}\otimes 1}$${k[V_{r}(\mathbb{G})]\otimes M\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{(\mu_{x})^{*}\otimes 1}$${k[V_{r}(\mathbb{G}_{a})]\otimes M^{g}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{1\otimes g\bullet}$${k[V_{r}(\mathbb{G}_{a})]\otimes M.}$
Consequently, the action of $(\mu_{x^{g}})_{*}(1\otimes u_{r-1})=(\Theta_{\mathbb{G},r,x^{g}})_{*}(1\otimes u_{r-1})$ on $M^{g}$ is isomorphic to the action of $(\mu_{x})_{*}(1\otimes u_{r-1})=(\Theta_{\mathbb{G},r,x})_{*}(1\otimes u_{r-1})$ on $M$, where $\Theta_{\mathbb{G},r,x^{g}}$ and $\Theta_{\mathbb{G},r,x}$ are the specializations of $\Theta_{\mathbb{G},r}$ at $x^{g}$ and $x$. ∎
In the theorem below, we equip (the scheme-theoretic points of) $\mathbb{P}V_{r}(\mathbb{G})$ with the Zariski topology and equip $\mathcal{Y}_{m}$ with the Alexandrov discrete topology (see Theorem 1.12). We refer to [15, Prop 2.8.3] for an earlier discussion of properties of the Jordan type function.
Theorem 2.21.
Let $\mathbb{G}$ be an affine group scheme of finite type over $k$ and let $r$ be a positive integer.
  1. (1)
    For any finite dimensional $\mathbb{G}_{(r)}$-module $M$, the function $JT_{\mathbb{G},r,M}(-)$ of (2.16.1) induces a continuous map
    $JT_{\mathbb{G},r,M}(-):\mathbb{P}V_{r}(\mathbb{G})\quad\to\quad\mathcal{Y}$
    sending a scheme-theoretic point $\overline{x}\in\mathbb{P}V_{r}(\mathbb{G})$ to $JT_{\mathbb{G},r,M}(x)$ for any choice of scheme-theoretic point $x\in V_{r}(\mathbb{G})$ projecting to $\overline{x}$.
  2. (2)
    The finite dimensional $\mathbb{G}_{(r)}$-module $M$ is injective if and only if the dimension $m$ of $M$ is divisible by $p$ and $JT_{\mathbb{G},r,M}(x)\ =\ \frac{m}{p}\cdot[p]$ for all $x\in\mathbb{P}V_{r}(\mathbb{G})$.
  3. (3)
    For any $\overline{x}\in\mathbb{P}V_{r}(\mathbb{G})$, $M\in mod(\mathbb{G}_{(r)},m)$, and $N\in mod(\mathbb{G}_{(r)},n)$,
    $JT_{\mathbb{G},r,M\oplus N}(\overline{x})\ =\ JT_{\mathbb{G},r,M}(\overline{x})+JT_{\mathbb{G},r,N}(\overline{x})\ \in\ \mathcal{Y}_{m+n}.$
  4. (4)
    If $r=1$ or if $\overline{x}\in\mathbb{P}V_{r}(\mathbb{G})$ is a generic point, if $M\in mod(\mathbb{G}_{(r)},m)$, and if $N\in mod(\mathbb{G}_{(r)},n)$, then
    $JT_{\mathbb{G},r,M\otimes N}(\overline{x})\ =\ JT_{\mathbb{G},r,M}(\overline{x})\otimes JT_{\mathbb{G},r,N}(\overline{x})\ \in\ \mathcal{Y}_{m\cdot n}.$
  5. (5)
    If $f:M\to N$ is a surjective map of finite dimensional $\mathbb{G}_{(r)}$-modules, then $JT_{\mathbb{G},r,N}(\overline{x})\ \leq\ JT_{\mathbb{G},r,M}(\overline{x})$ for any $\overline{x}\in\mathbb{P}V_{r}(\mathbb{G})$; in other words, $JT_{\mathbb{G},r,N}(-)\leq JT_{\mathbb{G},r,M}(-)$, where the partial ordering of functions is that inherited by the partial ordering of $\mathcal{Y}$.
  6. (6)
    Consider a short exact sequence $\mathcal{E}:0\to N\ \to\ M\ \to\ Q\to 0$ of finite dimensional $\mathbb{G}_{(r)}$-modules. Then $JT_{\mathbb{G},r,M}(-)=JT_{\mathbb{G},r,N}(-)+JT_{\mathbb{G},r,Q}(-)$ if and only if the short exact sequence $\mu_{x}^{*}(\mathcal{E})$ of $\mathbb{G}_{k(x)}$-modules splits for all $x\in V_{r}(\mathbb{G})$.
  7. (7)
    If $\phi:\mathbb{H}\to\mathbb{G}$ is a morphism of affine group schemes of finite type over $k$, then for any finite dimensional $\mathbb{G}$-module $M$
    $JT_{\mathbb{H},r,\phi^{*}(M)}(-):\mathbb{P}V_{r}(\mathbb{H})\ \to\ \mathcal{Y}$
    equals
    $JT_{\mathbb{G},r,M}(-)\circ\phi_{*}:\mathbb{P}V_{r}(\mathbb{H})\ \to\ \mathbb{P}V_{r}(\mathbb{G})\ \to\ \mathcal{Y}.$
Proof.
The fact that $JT_{\mathbb{G},r,M}(\overline{x})$ is well defined (namely, independent of the choice of $x\in V_{r}(\mathbb{G})$ projecting to $\overline{x}\in\mathbb{P}V_{r}(\mathbb{G})$), follows from the equality (2.18.1) of Theorem 2.18. To prove the asserted continuity of assertion (1) it suffices to prove continuity of $JT_{\mathbb{G},r,M}(-)$ when restricted to each affine open subset of the form $\operatorname{Spec}\nolimits(k[V_{r}(\mathbb{G})][1/h])_{0}\subset\mathbb{P}V_{r}(\mathbb{G})$, where $h\in k[V_{r}(\mathbb{G})]_{p^{r-1}}$ is an irreducible homogeneous generator of $k[V_{r}(\mathbb{G})]$ raised to the power $p^{r-1}/deg(h)$. This follows from the continuity proved in Theorem 1.12 for $A=(k[V_{r}(\mathbb{G})][1/h])_{0}$.
Assertion (2) follows from the detection theorem for nilpotence for $\mathbb{G}_{(r)}$-modules given in [27, Thm 4.3, Cor 4.4].
The fact that $JT_{\mathbb{G},r,M\oplus N}(-)=JT_{\mathbb{G},r,M}(-)+JT_{\mathbb{G},r,N}(-)$ is self evident. Since $(\mu_{x})_{*}:k(x)\otimes k\mathbb{G}_{a(r)}\to k(x)\otimes k\mathbb{G}_{(r)}$ is a map of Hopf algebras (over $k(x)$), the restriction map $\mu_{x}^{*}:mod(k(x)\otimes k\mathbb{G}_{(r)})\to mod(k(x)\otimes k\mathbb{G}_{a(r)})$ commutes with tensor products. Since $\epsilon_{1}$ is the identity, this implies assertion (4) whenever $r=1$. On the other hand, if $x$ is a generic point of $V_{r}(\mathbb{G})$ for any $r>0$, the fact that $M\ \mapsto\ JT_{\mathbb{G},r,M}(x)$ commutes with tensor products is given by [17, Prop 4.7].
A surjective map $f:M\to N$ of finite dimensional $\mathbb{G}_{(r)}$-modules induces a surjective map $k(x)\otimes M\ \to\ k(x)\otimes N$ of $k(x)[u]/t^{p}$-modules. Consequently, assertion (5) follows from Corollary 1.4. Assertion (6) follows from the definition of a locally split exact sequence (see [6]) and the fact for any field $K$ and any short exact sequence $0\to N_{K}\to M_{K}\to Q_{K}\to 0$ that the Jordan type of $M_{K}$ equals the sum of the Jordan types of $N_{K}$ and $Q_{K}$ if and only if the short exact sequence splits.
Assertion (7) is equivalent to the assertion that $JT_{\mathbb{G},r,M}(\phi_{*}(y))=JT_{\mathbb{H},r,\phi^{*}(M)}(y)$ for every $y\in\mathbb{P}V_{r}(\mathbb{H})$. This is immediate from the identification (with $x=\phi_{*}(y)$)
$JT((\mu_{x})_{*}\circ\epsilon_{r})^{*}(k(x)\otimes M)\quad=\quad JT((\mu_{y})_{*}\circ\epsilon_{r})^{*}(k(y)\otimes\phi^{*}(M)).$
The following corollary is rephrasing of some of the properties of Theorem 2.21.
Corollary 2.22.
We denote by $Hom_{cont}(\mathbb{P}V_{r}(\mathbb{G}),\mathcal{Y})$ the set of continuous maps from $\mathbb{P}V_{r}(\mathbb{G})$ with the Zariski topology to $\mathcal{Y}$ with the Alexandrov discrete topology. Give (the scheme-theoretic points of ) $\mathbb{P}V_{r}(\mathbb{G})$ the partial ordering of specialization and give $Hom_{cont}(\mathbb{P}V_{r}(\mathbb{G}),\mathcal{Y})$ the partial ordering inherited from that of $\mathcal{Y}$.
Then sending a finite dimension $\mathbb{G}_{(r)}$-module $M$ to $JT_{\mathbb{G},r,M}(-)$ determines
$JT_{\mathbb{G},r}:mod(\mathbb{G}_{(r)})\quad\to\quad Hom_{cont}(\mathbb{P}V_{r}(\mathbb{G}),\mathcal{Y})$
commuting with sums and the partial orderings.
The following theorem associates to a finite dimensional $\mathbb{G}$-module $M$ the configuration $\{(\mathbb{P}V_{r}(\mathbb{G}))_{M,\leq\underline{a}}\}$ which gives considerably more information about $M$ than does the support variety $\mathbb{P}V_{r}(\mathbb{G})_{M}\subset\mathbb{P}V_{r}(\mathbb{G})$ of $M$. The fact that $(V_{r}(\mathbb{G}))_{M,\leq\underline{a}}\subset V_{r}(\mathbb{G})$ is closed is given with a sketch of proof in [15, Prop 2.8.3].
Theorem 2.23.
(see Proposition 2.8.3 of [15]) Let $\mathbb{G}$ be an affine group scheme of finite type over $k$ and let $r$ be a positive integer. Consider a $\mathbb{G}_{(r)}$-module $M$ of dimension $m$. For each $\underline{a}\in\mathcal{Y}_{m}$, we define the subset
$(\mathbb{P}V_{r}(\mathbb{G}))_{M,\leq\underline{a}}\ \equiv\ \{\overline{x}:JT_{\mathbb{G},r,M}(\overline{x})\leq\underline{a}\}\ \subset\ \mathbb{P}V_{r}(\mathbb{G}).$
Then $\{(\mathbb{P}V_{r}(\mathbb{G}))_{M,\leq\underline{a}};\ \underline{a}\in\mathcal{Y}_{m}\}$ is a partially ordered configuration of closed subspaces of $\mathbb{P}V_{r}(\mathbb{G})$.
With this notation, the support variety $\mathbb{P}V_{r}(\mathbb{G})_{M}$ of [27] equals the union of those subspaces $(\mathbb{P}V_{r}(\mathbb{G}))_{M,\underline{a}\leq}$ indexed by $\underline{a}\in\mathcal{Y}_{m}$ which satisfy the condition that $a_{i}\not=0$ for some $i<p$. (See Theorem 2.6.)
The non-maximal support variety $\Gamma(\mathbb{G}_{(r)})_{M}$ of [15] equals the union of those subspaces $(\mathbb{P}V_{r}(\mathbb{G}))_{M,\underline{a}\leq}$ indexed by those $\underline{a}\in\mathcal{Y}_{m}$ which satisfy the condition that there exists some $\underline{b}\in\mathcal{Y}_{m}$ with $\underline{a}\leq\underline{b}$ such that $(\mathbb{P}V_{r}(\mathbb{G}))_{M,\leq\underline{a}}$ is strictly contained in $(\mathbb{P}V_{r}(\mathbb{G}))_{M,\leq\underline{b}}$
Proof.
The continuity of $JT_{\mathbb{G},r,M}(-)$ together with the evident fact for any $\underline{a}\in\mathcal{Y}$ that $\{\underline{c}\in\mathcal{Y}:\ \underline{c}\leq a\}$ is closed in $\mathcal{Y}$ implies $(\mathbb{P}V_{r}(\mathbb{G}))_{M,\leq\underline{a}}$ is closed in $\mathbb{P}V_{r}(\mathbb{G})$. The identification of the subspaces $\mathbb{P}V_{r}(\mathbb{G})_{M}=\cup_{\underline{a}\not=n[p]}(\mathbb{P}V_{r}(\mathbb{G})_{M,\leq\underline{a}})$ and $\Gamma(\mathbb{G}_{(r)})_{M}$ (as asserted above) inside $\mathbb{P}V_{r}(\mathbb{G})$ follows immediately from the definitions of these subspaces. ∎
We make explicit the following corollary first stated and proved in [15, Cor 2.15].
Corollary 2.24.
Let $M$ be a $\mathbb{G}_{(r)}$-module of dimension $m$, and let $\underline{a}$ be an object on $\mathcal{Y}_{\leq,m}$, The subset
$(\mathbb{P}V_{r}(\mathbb{G}))_{M,=\underline{a}}\ \equiv\ \{x:JT_{\mathbb{G},r,M}(x)=\underline{a}\}\ \subset\ \mathbb{P}V_{r}(\mathbb{G})$
is locally closed, the union of those $(\mathbb{P}V_{r}(\mathbb{G}))_{M,\leq\underline{b}}$ such that $\underline{b}\leq\underline{a}$ and $\underline{b}\not=\underline{a}$,
$(\mathbb{P}V_{r}(\mathbb{G}))_{M,=\underline{a}}\quad=\quad(\mathbb{P}V_{r}(\mathbb{G}))_{M,\leq\underline{a}}-(\cup_{\underline{a}\not=\underline{b}\leq\underline{a}}(\mathbb{P}V_{r}(\mathbb{G}))_{M,\leq\underline{b}}).$
Remark 2.25.
We can interpret Corollary 2.24 as asserted that the Jordan type function $JT_{\mathbb{G},r,M}(-)$ for a finite dimensional $\mathbb{G}_{(r)}$-module $M$ determines a stratification of $\mathbb{P}V_{r}(\mathbb{G})$ with locally closed strata, each colored by a Young diagram such that if one stratum is strictly contained in the closure of another stratum then the Young diagram of the smaller stratum is less than the Young diagram of the bigger stratum.
Remark 2.26.
In [15, Defn 4.1], generalized support varieties $\Gamma^{j}(\mathbb{G})_{M}$ were introduced for finite dimensional $\mathbb{G}$-modules $M$ for an arbitrary finite group scheme $\mathbb{G}$ over $k$. Namely, $\Gamma^{j}(\mathbb{G})_{M}$ is the subspace of those $\pi$-points $\alpha_{K}]\in\Pi(\mathbb{G})$ with the property that the rank of $\alpha_{K}(t^{j})$ on $M_{K}$ is not maximal. For $\mathbb{G}$ infinitesimal, one can recover $\Gamma^{j}(\mathbb{G})_{M}$ as in Theorem 2.23 by replacing $\Theta_{\mathbb{G},r,M}$ in Theorem 2.15 by its $j$-th power, then replacing $JT((\mu_{x})_{*}\circ\epsilon_{r})^{*}(k[V_{r}(\mathbb{G})]\otimes M)$ in Theorem 2.21 by the Jordan type of the $j$-th power of $(\mu_{x})_{*}(u_{r-1})$ acting on $k(x)\otimes M$.
3. First examples
In this section, we consider a few general examples of computations of $JT_{\mathbb{G},r,M}(-)$ (some found in [15, §2]). More specific computations for the simpler $JT^{exp}_{\mathbb{G},r,M}(-)$ are given in Section 5.
Example 3.1.
Let $\mathbb{G}=\mathbb{G}_{(r)}$ be an infinitesimal group scheme of height $\leq r$. A $\mathbb{G}$-module $M$ of dimension $m$ is of constant Jordan type $\underline{a}$ if and only if $JT_{\mathbb{G},M,r}(-)$ is the constant function sending each $\overline{x}\in\mathbb{P}V_{r}(\mathbb{G})$ to $\underline{a}\in\mathcal{Y}_{m}$. (See [7] for numerous examples.)
In particular, if $I$ is an injective $\mathbb{G}$-module of dimension $m$, then $p$ divides $m$ and $JT_{\mathbb{G},I,r}(-)$ sends each $\overline{x}\in\mathbb{P}V_{r}(\mathbb{G})$ to $\frac{m}{p}[p]\in\mathcal{Y}_{m}$ (i.e., $JT_{\mathbb{G},I-r}(\overline{x})=\frac{m}{p}[p]+0[p-1]+\cdots+0[1]$).
Example 3.2.
Let $\mathbb{G}=\mathbb{G}_{(r)}$ be an infinitesimal group scheme of height $\leq r$ and $\overline{x}\in\mathbb{P}V_{r}(\mathbb{G})$. Consider $0\not=\zeta\in H^{2d}(\mathbb{G},k)$ and take $M$ to be the Carlson $L_{\zeta}$-module as first constructed in [5]; namely, $M$ is the kernel of a map $\tilde{\zeta}:\Omega^{2d}(k)\to k$ representing $\zeta$.
Then the dimension $n$ of $M$ is divisible by $p$. Moreover,
$J_{\mathbb{G},M,r}(\overline{x})=\frac{n}{p}[p]\quad{\text{i}f}\quad((\mu_{x})_{*}\circ\epsilon_{r})^{*}(\zeta))\not=0\in H^{2d}(k(x)[u]/u^{p},k),$
$J_{\mathbb{G},M,r}(\overline{x})=(\frac{n}{p}-1)[p]+[p-1]+[1]\quad{\text{i}f}\quad((\mu_{x})_{*}\circ\epsilon_{r})^{*}(\zeta))=0\in H^{2d}(k(x)[u]/u^{p},k).$
Example 3.3.
(See [15, Prop 2.9]) Let $\mathbb{G}=\mathbb{G}_{(r)}$ be an infinitesimal group scheme of height $\leq r$. Let $M$ be a finite dimensional $\mathbb{G}$-module and $P\twoheadrightarrow M$ a finite dimensional projective cover of $M$ with kernel $\Omega(M)$. If $JT_{\mathbb{G},M,r}(\overline{x})\ =\underline{a}$ for some $\overline{x}\in\mathbb{P}V_{r}(\mathbb{G})$, then $JT_{\mathbb{G},\Omega(M),r}(\overline{x})\ =\underline{a}^{\perp}+n[p]$ for some integer $n$, where $\underline{a}^{\perp}=\sum_{i=1}^{p-1}a_{p-i}[i]$.
Example 3.4.
(See [20, §I.3.8].) Consider $\mathbb{G}_{1},\ \mathbb{G}_{2}$, infinitesimal group schemes of height $\leq r$. Let $\mathbb{G}=\mathbb{G}_{1}\times\mathbb{G}_{2}$ and $M_{1}$ be a a finite dimensional $\mathbb{G}_{1}$-module. Set $M$ equal to $ind_{\mathbb{G}_{1}}^{\mathbb{G}}M_{1}\simeq M_{1}\otimes k[\mathbb{G}_{2}]$. Let $\overline{(x,y)}$ be a scheme-theoretic point in $\mathbb{P}V_{r}(\mathbb{G})$ represented by the scheme-theoretic point $(x,y)\in V_{r}(\mathbb{G}_{1})\times V_{r}(\mathbb{G}_{2})$. Then
$JT_{\mathbb{G},M,r}(\overline{(x,0}))\ =\ dim(k[\mathbb{G}_{2}])\cdot JT_{\mathbb{G}_{1},M_{1}}((\overline{x})),$
$JT_{\mathbb{G},M,r}(\overline{(x,y}))\ =\ dim(M)/p\cdot[p],\quad y\not=0.$
We generalize the previous example as follows.
Example 3.5.
Consider $\mathbb{G}_{1},\ \mathbb{G}_{2}$, infinitesimal group schemes of height $\leq r$, a finite dimensional $\mathbb{G}_{1}$-module $M_{1}$, and a finite dimensional $\mathbb{G}_{2}$-module $M_{2}$. Let $(x,y)$ be a scheme-theoretic point in $V_{r}(\mathbb{G}_{1})\times V_{r}(\mathbb{G}_{2})$. Then
$JT_{\mathbb{G}_{1}\times\mathbb{G}_{2},M\otimes N,r}((x,y))\quad=\quad JT_{\mathbb{G}_{1},M,r}(x)\otimes JT_{\mathbb{G}_{2},N,r}(y).$
Example 3.6.
Let $\mathbb{G}$ be an infinitesimal group scheme of height 1 with Lie algebra $\mathfrak{g}$. Thus, $k\mathbb{G}$ is naturally identified with $\mathfrak{u}(\mathfrak{g})$, the restricted enveloping algebra of $\mathfrak{g}$ and $V_{1}(\mathbb{G})$ is naturally identified with the $p$-nilpotent cone $\mathcal{N}_{p}(\mathfrak{g})\subset\mathfrak{g}$. (See [26, Lem 1.6].)
Let $V$ be a faithful representation of $\mathbb{G}$ of dimension $N$, and associate to $V$ a closed embedding $\phi:\mathbb{G}\hookrightarrow GL_{N(1)}\hookrightarrow GL_{N}$. For any $\alpha\in\mathcal{N}_{p}(\mathfrak{g})(A)$, the associated 1-parameter subgroup of $\mu_{\alpha}:\mathbb{G}_{a(1),A}\to\mathbb{G}_{A}$ sends $s\in\mathbb{G}_{a(1)}(A)$ (i.e., $s$ a $p$-nilpotent element of $A$) to
$\mu_{\alpha}(s)\ =1+s\cdot\alpha+\frac{s^{2}}{2}\alpha^{2}+\cdots+\frac{s^{p-1}}{(p-1)!}\alpha^{p-1}\ \in\ \mathbb{G}_{(1)}(A)\ \subset\ GL_{N}(A),$
where the powers of $\alpha$ are taken in $\mathfrak{g}l_{N}$. We associate to $\mu_{\alpha}$ the linear operator $(\mu_{\alpha})_{*}:A\otimes k\mathbb{G}_{a(1)}\to\mathcal{A}\otimes k\mathbb{G}$.
Let $M$ be a finite dimensional $\mathbb{G}$-module of dimension $m$ and choose a scheme-theoretic point $\alpha\in\mathcal{N}_{p}(\mathfrak{g})$, given by $\alpha\in k(\alpha)\otimes\mathcal{N}_{p}(\mathfrak{g})$. Then $JT_{\mathbb{G},M,1}(\alpha)$ is the Jordan type of $(\mu_{\alpha})_{*}(1\otimes u)\in k(\alpha)\otimes k\mathbb{G}$ acting on $k(\alpha)\otimes M$. This action is the action of $\alpha$ as an element of the Lie algebra $k(\alpha)\otimes\mathfrak{g}$ action on $k(\alpha)\otimes M$.
Example 3.7.
Assume $\mathbb{G}=\mathbb{G}_{a(r)}$ for some $r>0$. Then $\mathbb{G}$ admits an embedding in $GL_{2}$ as the $r$-th Frobenius kernel of the unipotent subgroup $U_{2}$ of strictly upper triangular matrices. As in Example 2.10, $\mathbb{P}V_{r}(\mathbb{G})$ is a weighted projective space with weights $(1,p,\cdots,p^{r-1})$ associated to the graded polynomial algebra $k[T_{0},T_{1},\ldots,T_{r-1}]$ where $T_{i}$ has weight $p^{i}$. Consider a non-zero $r$-tuple $\underline{a}=(a_{0},\ldots,a_{r-1})$ of elements of $k(x)$ determining a scheme-theoretic point $x\in V_{r}(X)$ projecting to the scheme-theoretic point $\overline{x}\in\mathbb{P}V_{r}(\mathbb{G})$, and consider the associated 1-parameter subgroup $\mu_{x}:\mathbb{G}_{a(r),k(x)}\to\mathbb{G}_{a(r),k(x)}.$ An explicit (but quite complicated) formula for $\mu_{x*}(1\otimes u_{r-1})\in k(x)\otimes k\mathbb{G}_{a(r)}\ =\ k(x)\otimes k[u_{0},\ldots,u_{r-1}]/(u_{i}^{p})$ is given in [27, (6.5.12)].
The sum of terms linear in the $u_{i}$’s in the expression for $\mu_{x*}(1\otimes u_{r-1})$ is $\sum_{i=0}^{r-1}a_{r-1-i}^{p^{i}}u_{i}\in k(x)\otimes k\mathbb{G}$. Replacing $\mu_{x*}(1\otimes u_{r-1})$ by $\sum_{i=0}^{r-1}a_{r-1-i}^{p^{i}}u_{i}$ is a special case of replacing $\Theta_{\mathbb{G},r}$ by $\Theta^{exp}_{\mathfrak{g},r}$ as discussed in Section 4.
4. Exponential Jordan type
In this section, we consider affine group schemes whose 1-parameter subgroups are given by exponentiating $r$-tuples of $p$-nilpotent, pair-wise commuting elements of $\mathfrak{g}\equiv Lie(\mathbb{G})$. The more explicit nature of these 1-parameter subgroups allows us to simplify our Jordan type functions, not necessarily preserving their values but preserving their close relationship with support varieties.
For an affine group scheme $\mathbb{G}$ equipped with an exponential map $\mathcal{E}_{(r)}:\mathbb{G}_{a(r)}\times\mathcal{N}_{p}(\mathfrak{g})\to\mathbb{G}_{(r)}$ of height $r$, there is a natural morphism $\Phi_{\mathfrak{g},r}:\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))\ \to\ V_{r}(\mathbb{G})$ where $\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))$ is the affine variety of $r$-tuples of $p$-nilpotent, pair-wise commuting elements of $\mathfrak{g}$ made explicit in Definition 4.2. We say that $\mathbb{G}$ is of exponential type of height $r$ if $\Phi_{\mathfrak{g},r}$ is a homeomorphism of scheme-theoretic points. Examples include Frobenius kernels of various classical groups, their parabolic subgroups, and unipotent subgroups of these parabolic subgroups. (See, for example, [25].) For $\mathbb{G}$ of exponential type of height $r$, Theorem 4.8 introduces $\Theta_{\mathfrak{g},r}\in k[\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))]\otimes k\mathbb{G}_{(r)}$ which very closely models $\Theta_{\mathbb{G},r}$.
This leads to our simpler $p$-nilpotent operator $\Theta_{\mathfrak{g},r}^{exp}\in k[\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))]\ \otimes k\mathbb{G}_{(r)}$ for $\mathbb{G}$ of exponential type of height $r$, a linearization of $\Theta_{g,r}$. A fundamental computation of [17, Thm 1.12] enables us to conclude that $\Theta_{\mathfrak{g},r}$ and $\Theta_{\mathfrak{g},r}^{exp}$ determine the same support varieties for finite dimensional $\mathbb{G}_{(r)}$-modules. By adopting the formalism we have discussed for $JT_{\mathbb{G},r,M}(-)$, we establish the Jordan type functions $JT_{\mathfrak{g},r,M}(-)$ and $JT_{\mathfrak{g},r,M}^{exp}(-)$ which satisfy similar properties.
We begin by defining the class of affine group schemes $\mathbb{G}$ for which $r$-tuples of $p$-nilpotent, pair-wise commuting elements of the Lie algebra of $\mathbb{G}$ determine height $r$ 1-parameter subgroups of $\mathbb{G}$.
Definition 4.1.
Let $\mathbb{G}$ be an affine group scheme of finite type over $k$ with Lie algebra $\mathfrak{g}$. Fix a positive integer $r$. We say that a morphism
$\mathcal{E}_{(r)}:\mathbb{G}_{a(r)}\times\mathcal{N}_{p}(\mathfrak{g})\quad\to\quad\mathbb{G}_{(r)}$
equips $\mathbb{G}$ with the structure of an exponentiation of height $r$ if $\mathcal{E}_{(r)}$ satisfies the following conditions for any commutative $k$-algebra AA, $p$-nilpotent elements $B,B^{\prime}\ \in A\otimes\mathfrak{g}$, and elements $\alpha,\beta\in\mathbb{G}_{a(r)}(A)$:
  1. (1)
    The specialization of $\mathcal{E}_{(r)}$ along $B:k[\mathcal{N}_{p}(\mathfrak{g})]\to A$, $\mathcal{E}_{(r),B}:\mathbb{G}_{a(r),A}\to\mathbb{G}_{(r),A}$, is a 1-parameter subgroup of $\mathbb{G}_{(r),A}$.
  2. (2)
    If $B^{\prime}$ commutes with $B$, then
    $\mathcal{E}_{(r),B}(\alpha)\cdot\mathcal{E}_{(r),B^{\prime}}(\beta)=\mathcal{E}_{(r),B^{\prime}}(\beta)\cdot\mathcal{E}_{(r),B}(\alpha)\in\mathbb{G}_{(r)}(A).$
  3. (3)
    $\mathcal{E}_{(r),\alpha\cdot B}(-)\ =\ \mathcal{E}_{(r),B}(\alpha\cdot-):\mathbb{G}_{a(r),A}\to\mathbb{G}_{(r),A}$.
As exhibited in Construction 4.3, the structure $\mathcal{E}_{(r)}$ of exponentiation of height $r$ for $\mathbb{G}$ enables us to “exponentiate” $r$-tuples of $p$-nilpotent, pair-wise commuting elements of $\mathfrak{g}$. We first define the variety of such $r$-tuples, $\mathcal{C}_{r}(\mathfrak{g})$.
Definition 4.2.
Let $\mathbb{G}$ be an affine group scheme of finite type over $k$ with Lie algebra $\mathfrak{g}$. Consider the functor on commutative $k$ algebras sending $A$ to an $r$-tuple $\underline{B}=(B_{0},\ldots,B_{r-1})$ of elements in $A\otimes\mathfrak{g}$ which satisfies the conditions that $B_{s}^{[p]}=0$ for all $s,\ 0\leq s<r$ and that $B_{s},B_{s^{\prime}}]=0$ for all $s,s^{\prime}$ with $0\leq s,s^{\prime}<r$. This functor is represented by an affine scheme $\mathcal{C}_{r}(\mathfrak{g})$. In contrast to [13, 1.11.2], we give $\mathcal{C}_{r}(\mathfrak{g})$ the monoid action $\mathbb{A}^{1}\times\mathcal{C}_{r}(\mathfrak{g})\to\mathcal{C}_{r}(\mathfrak{g})$ by sending $(\alpha;B_{0},\ldots,B_{r-1})$ to $(\alpha\cdot B_{0},\alpha^{p}\cdot B_{1},\ldots,\alpha^{p^{r-1}}\cdot B_{r-1})$.
We now assume the $\mathbb{G}$ is equipped with an exponential $\mathcal{E}_{(r)}$ of height $r$. The following construction of $\mathcal{E}_{\mathfrak{g},r}:\mathbb{G}_{a(r)}\times\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))\ \to\ \mathbb{G}_{(r)}$ given in (4.3.2) extends the exponential structure $\mathcal{E}_{(r)}:\mathbb{G}_{a(r)}\times\mathcal{N}_{p}(\mathfrak{g})\ \to\ \mathbb{G}_{(r)}$. This leads to the “Lie algebra analogue”
$\mathcal{U}_{\mathfrak{g},r}:\mathbb{G}_{a(r)}\times\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))\quad\to\quad\mathbb{G}_{(r)}\times\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))$
of the universal 1-parameter subgroup $\mathcal{U}_{\mathbb{G},r}:\mathbb{G}_{a(r)}\times V_{r}(\mathbb{G})\to\mathbb{G}_{(r)}\times V_{r}(\mathbb{G})$ of Definition 2.12.
Construction 4.3.
Let $\mathbb{G}$ be an affine group scheme of finite type over $k$ equipped with an exponentiation $\mathcal{E}_{(r)}:\mathbb{G}_{a(r)}\times\mathcal{N}_{p}(\mathfrak{g})\to\mathbb{G}_{(r)}$ of height $r$. For any $s,\ 0\leq s<r$, we define
(4.3.1) $\mathcal{E}_{r|s}\quad\equiv\quad\mathcal{E}_{(r)}\circ(1\times pr_{s})\circ(F^{s}\times 1):\mathbb{G}_{a(r)}\times\mathcal{C}_{r}(\mathfrak{g})\ \to\ \mathbb{G}_{(r)}$
where $pr_{s}$ is projection onto the $s$-th factor and $F^{s}$ is the $s$-th iterate of the Frobenius morphism. We denote by $\mathcal{E}_{r|s,B}$ the restriction of $\mathcal{E}_{r|s}$ along a (geometric) point $B:k(B)\to\mathcal{N}_{p}(\mathfrak{g})$.
We define
(4.3.2) $\mathcal{E}_{\mathfrak{g},r}\quad\equiv\quad\prod\circ\ (\times_{s=0}^{r-1}\ \mathcal{E}_{r|s})\circ Diag$
to be the composition
$\mathbb{G}_{a(r)}\times\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))\ \to(\mathbb{G}_{a(r)}\times\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g})))^{\times r}\ \to\ (\mathbb{G}_{(r)})^{\times r}\ \to\mathbb{G}_{(r)}$
where the product is taken with respect to the group structure of $\mathbb{G}_{(r)}$.
We define
(4.3.3) $\mathcal{U}_{\mathfrak{g},r|s}\equiv\mathcal{E}_{r|s}\times pr_{\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))},\ \mathcal{U}_{\mathfrak{g},r}\equiv\mathcal{E}_{\mathfrak{g},r}\times pr_{\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))}:\mathbb{G}_{a(r)}\times\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))\ \to\ \mathbb{G}_{(r)}\times\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g})),$
given by maps on coordinate agebras as the the $k[\mathcal{C}_{r}(\mathfrak{g})]$-linear extensions of $\mathcal{E}_{r|s}^{*},\ \mathcal{E}_{\mathfrak{g},r}^{*}$ along $k\to k[\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))]$.
The following proposition identifies the specializations of the maps $\mathcal{U}_{\mathfrak{g},r|s}$ and $\mathcal{U}_{\mathfrak{g},r}$ at a geometric point of $\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))$. In particular, the specialization of $\mathcal{U}_{\mathfrak{g},r}$ at a geometric point $\underline{B}=(B_{0},\ldots,B_{r-1})$ of $\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))$ is expressed as a product of Frobenius twists of exponentiations of the $B_{i}$’s.
Proposition 4.4.
Retain the hypotheses and notation of Definition 4.1. For any geometric point $\underline{B}=(B_{0},\ldots,B_{r-1})$ of $\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))$, the specialization of $\mathcal{U}_{\mathfrak{g},r|s}$ at $\underline{B}$ is denoted by $\mathcal{U}_{\mathfrak{g},r|s,B_{s}}$. Then the specialization of $\mathcal{U}_{\mathfrak{g},r}$ at $\underline{B}$ is given by
(4.4.1) $\mathcal{U}_{\mathfrak{g},r,\underline{B}}\ =\ \prod_{s=0}^{r-1}\mathcal{U}_{\mathfrak{g},r|s,B_{s}}\circ F^{s}:\mathbb{G}_{a(r),k(\underline{B})}\ \to\mathbb{G}_{a(r),k(\underline{B})}\ \to\ \mathbb{G}_{(r),k(\underline{B})}.$
In particular,
(4.4.2) $(\mathcal{U}_{\mathfrak{g},r,\underline{B}})_{*}:k(\underline{B})\otimes k\mathbb{G}_{a(r)}\to k(\underline{B})\otimes k\mathbb{G}_{(r)},\quad 1\otimes u_{r-1}\ \mapsto\ \prod_{s=0}^{r-1}(\mathcal{E}_{r|s,B_{s}})_{*}(u_{r-s-1}).$
Proof.
Equations (4.4.1) and (4.4.1) follow from equations (4.3.1) and (4.3.2) together with the definitions of $\mathcal{U}_{\mathfrak{g},r|s,\underline{B}}$ and $\mathcal{U}_{\mathfrak{g},r,\underline{B}}$ as $k[\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))]$-linear extensions of $\mathcal{E}_{r|s},\ \mathcal{E}_{\mathfrak{g},r}$.
Equation (4.4.2) follows immediately from (4.4.1). ∎
We proceed to investigate the close relationship between $\mathcal{U}_{\mathbb{G},r}$ and $\mathcal{U}_{\mathfrak{g},r}$.
Proposition 4.5.
Let $\mathbb{G}$ and $\mathcal{E}_{(r)}$ satisfy the conditions of Definition 4.1.
  1. (1)
    The morphisms
    $\mathcal{U}_{\mathfrak{g},r|s},\ \mathcal{U}_{\mathfrak{g},r}:\mathbb{G}_{a(r),k[\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))]}\quad\to\quad\mathbb{G}_{(r),k[\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))]}$
    are 1-parameter subgroups.
  2. (2)
    These 1-parameter subgroups determine morphisms
    $\Phi_{\mathfrak{g},r|s},\ \Phi_{\mathfrak{g},r}:\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))\quad\to\quad V_{r}(\mathbb{G})$
    with the property that their pull-backs of the universal 1-parameter group $\mathcal{U}_{\mathbb{G},r}:\mathbb{G}_{a(r),k[V_{r}(\mathbb{G})]}\to\ \mathbb{G}_{(r),k[V_{r}(\mathbb{G})]}$ equal the 1-parameter subgroups $\mathcal{U}_{\mathfrak{g},r|s},\ \mathcal{U}_{\mathfrak{g},r}$.
  3. (3)
    The morphisms $\Phi_{\mathfrak{g},r|s},\Phi_{\mathfrak{g},r}$ commute with the monoid actions of $\mathbb{A}^{1}$ on $\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))$ and $V_{r}(\mathbb{G})$ in the sense that each of the following squares commute:
    (4.5.1) ${\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))\times\mathbb{A}^{1}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\Phi_{\mathfrak{g},r|s}\times 1}$${\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\Phi_{r,s}}$${V_{r}(\mathbb{G})\times\mathbb{A}^{1}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${V_{r}(\mathbb{G}).}$
Proof.
As can be seen using Proposition 4.4 and the condition of Definition 4.1(2), $\mathcal{U}_{\mathfrak{g},r|s,\underline{B}}$ and $\mathcal{U}_{\mathfrak{g},r,\underline{B}}$ are morphisms of group schemes over $\underline{B}$. This implies assertion (1). Assertion 2 follows from the universal property of $\mathcal{U}_{\mathbb{G},r}:\mathbb{G}_{a(r),k[V_{r}(\mathbb{G})]}\quad\to\quad\mathbb{G}_{(r),k[V_{r}(\mathbb{G})]}$.
To check the commutativity of (4.5.1), it suffices to observe for any $\alpha\in A$ and for any $A$-point $\mathcal{C}_{r}(\mathbb{G})$ of the form $(0,\ldots,B_{s},\ldots 0)$ (and thus of weight $p^{s}$) that $\mathcal{E}_{(r),\alpha\cdot\underline{B}}(t)\ =\mathcal{E}_{(r),(0,\ldots,\alpha^{p^{s}}B_{s},\ldots 0)}(t)=\mathcal{E}_{(r),\underline{B}}(\alpha^{p^{s}}t)=(\alpha\cdot\mathcal{E}_{(r),\underline{B}})(t)$. ∎
In [26], Suslin, Bendel, and the author introduced the condition that an embedding $\mathbb{G}\hookrightarrow GL_{N}$ of linear algebraic groups be of exponential type; they observe that various algebraic groups satisfy this condition, including many classical simple groups, their parabolic subgroups, and unipotent radicals of these parabolic subgroups. Subsequent generalizations of this condition on $\mathbb{G}$ occur in [12] and [13]. A more definitive account of such linear algebraic groups can be found in P. Sobaje’s article [25]. The following definition is a natural generalization of those earlier considerations of groups of exponential type. The motivating special case is the exponential function of $GL_{N}$ applied to a $p$-nilpotent element of $\mathfrak{g}l_{N}$.
Definition 4.6.
Let $\mathbb{G}$ be an affine group scheme of finite type over $k$ equipped with an exponentiation $\mathcal{E}_{(r)}:\mathbb{G}_{a(r)}\times\mathcal{N}_{p}(\mathfrak{g})\ \to\ \mathbb{G}_{(r)}$ of height $\leq r$. We say that $\mathbb{G}$ equipped with $\mathcal{E}_{(r)}$ is of exponential type of height $r$ if the morphism $\Phi_{\mathfrak{g},r}:\mathcal{C}_{r}(\mathfrak{g})\ \to\ V_{r}(\mathbb{G})$ of Theorem 4.5(1) induces a bijection on geometric points.
In [11, Defn 1.6], the notions of a “continuous algebraic map” and a “bicontinuous morphism” were introduced in order to study rational maps between schemes which induce well defined maps on geometric points. In that work, the topologies of interest were either the ètale topology or the analytic topology (in the case that the ground field was the complex numbers). We are now working with varieties over $k$, a field of characteristic $p>0$.
Lemma 4.7.
Let $(\mathbb{G},\mathcal{E}_{(r)})$ be an affine group scheme of exponential type of height $r$. The bicontinuous morphism
(4.7.1) $\Phi_{\mathfrak{g},r}:\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))\quad\to\quad V_{r}(\mathbb{G})$
is a homeomorphism of Zariski topological spaces which commutes with $\mathbb{A}^{1}$-actions.
Proof.
The assumption of exponential type of height $r$ implies that $\Phi_{\mathfrak{g},r}$ is a bijection on scheme-theoretic points. Since $\Phi_{\mathfrak{g},r}$ is a morphism, it is continuous. Since $\mathbb{P}\Phi_{\mathfrak{g},r}:\mathbb{P}\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))\to\mathbb{P}V_{r}(\mathbb{G})$ is proper, it is a closed map, thus induces a bijection on closed subsets, thus a homeomorphism. This implies that the pull-back/restriction $\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))-\{0\}\to V_{r}(\mathbb{G})-\{0\}$ is a homeomorphism. Since the cone point 0 is closed in both $\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))$ and $V_{r}(\mathbb{G})$, this implies that $\Phi_{\mathfrak{g},r}$ is also a homeomorphism. ∎
We adapt the construction of $\Theta_{\mathbb{G},r}$ in Definition 2.13 and its use in Theorem 2.15 by replacing $\mathcal{U}_{\mathbb{G},r}$ by $\mathcal{U}_{\mathfrak{g},r}$.
Theorem 4.8.
Let $(\mathbb{G},\mathcal{E}_{(r)})$ be an affine group scheme of exponential type of height $r$. We define the $p$-nilpotent operators
(4.8.1) $\Theta_{\mathfrak{g},r|s}\ \equiv\ (\mathcal{U}_{\mathfrak{g},r|s})_{*}(1\otimes u_{r-1}),\ \Theta_{\mathfrak{g},r}\ \equiv\ (\mathcal{U}_{\mathfrak{g},r})_{*}(1\otimes u_{r-1})\ \in\ k[\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g})))]\otimes k\mathbb{G}_{(r)},$
and denote by $\Theta_{\mathfrak{g},r|s,\underline{B}},\ \Theta_{\mathfrak{g},r,\underline{B}}$ their specializations at a scheme-theoretic point $\underline{B}\in\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))$.
  1. (1)
    $\Theta_{\mathfrak{g},r}\ =\ (\Phi^{*}_{\mathfrak{g},r}\otimes 1)(\Theta_{\mathbb{G},r})$.
  2. (2)
    $\Theta_{\mathfrak{g},r|s,\underline{B}}\ =\ (\mathcal{E}_{r|s,B_{s}})_{*}(u_{r-s-1}),\ \Theta_{\mathfrak{g},r,\underline{B}}\ =\prod_{s=0}^{r-1}\Theta_{\mathfrak{g},r|s,\underline{B}}\quad\in\ k(\underline{B})\otimes k\mathbb{G}_{(r)},$ elements of the group algebra of $\mathbb{G}_{(r),k(\underline{B})}$.
  3. (3)
    Both $\Theta_{\mathfrak{g},r|s}$ and $\Theta_{\mathfrak{g},r}$ are homogeneous of degree $p^{r-1}$ with respect to the grading on $k[\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))]$.
  4. (4)
    For any $\mathbb{G}_{(r)}$-module $M$, tensor product with $\Theta_{\mathfrak{g},r|s}$ and with $\Theta_{\mathfrak{g},r}$ determines $p$-nilpotent, $k[\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))]$-linear operators
    $\Theta_{\mathfrak{g},r|s,M},\ \Theta_{\mathfrak{g},r,M}:k[\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))]\otimes M\ \to\ k[\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))]\otimes M.$
  5. (5)
    The base changes along a scheme theoretic point $\underline{B}\in\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))$ of the operators $\Theta_{\mathfrak{g},r|s,M},\ \Theta_{\mathfrak{g},r,M}$ are given by the actions of $\Theta_{\mathfrak{g},r|s,\underline{B}},\ \Theta_{\mathfrak{g},r,\underline{B}}$ on $k(\underline{B})\otimes M$.
Proof.
Observe that the elements $\Theta_{\mathfrak{g},r|s},\ \Theta_{\mathfrak{g},r}$ are $p$-nilpotent because $1\otimes u_{r-1}\in\mathbb{G}_{a(r),\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))]}$ is $p$-nilpotent. Assertion (1) follows from Proposition 4.5(2), the definition of $\Theta_{\mathbb{G},r}$ in Definition 2.13, and the definition of $\Theta_{\mathfrak{g},r}$ given above. The first part of assertion (2) identifying $\Theta_{\mathfrak{g},r|s,\underline{B}}$ follows from Proposition 4.4, so that the second part is verified using the fact that specialization along $k[\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))]\to k(\underline{B})$, $k[\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))]\otimes\mathbb{G}_{(r)}\to k(\underline{B})\otimes\mathbb{G}_{(r)}$, commutes with the product map of $\mathbb{G}_{(r)}$.
Assertion (3) follows from Theorem 2.15(5). Assertions (4) and (5) and their proofs are direct analogues of Theorem 2.15(3),(4). ∎
As an immediate corollary of Theorem 4.8(1), we have the following.
Corollary 4.9.
Let $(\mathbb{G},\mathcal{E}_{(r)})$ be an affine group scheme of exponential type of height $r$. Consider a geometric point $\underline{B}$ of $\mathcal{C}_{r}(\mathfrak{g})$ with image $x=\Phi_{\mathfrak{g},r}(\underline{B})$. For any finite dimensional $\mathbb{G}_{(r)}$-module $M$,
$JT(\Theta_{\mathfrak{g},r,M})(\underline{B}))\quad=\quad JT(\Theta_{\mathbb{G},r,M})(x).$
In other words,
$JT(\Theta_{\mathfrak{g},r,M})(-)\ =\ JT(\Theta_{\mathbb{G},rM}(-)\circ\Phi_{\mathfrak{g},r}(-):\mathbb{P}\mathcal{C}_{r}(\mathfrak{g})\ \to\ \mathcal{Y},$
for any finite dimensional $\mathbb{G}_{(r)}$-module $M$.
We denote by $\mathbb{P}\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))$ the weighted projective variety $\operatorname{Proj}\nolimits k[\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))]$, parallel to the notation of $\mathbb{P}V_{r}(\mathbb{G})$ for $\operatorname{Proj}\nolimits k[V_{r}(\mathbb{G})]$. We omit the proof of the following elementary lemma.
Lemma 4.10.
A geometric point of $\operatorname{Spec}\nolimits K\to\mathbb{P}\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))$ corresponds to an equivalence class $\langle\underline{B}\rangle$ of non-zero $r$-tuples $\underline{B}=(B_{0},\ldots,B_{r-1})$ of $p$-nilpotent, pair-wise commuting elements of $K\otimes\mathfrak{g}$, where $(B_{0},\ldots,B_{r-1})$ is equivalent to $(B_{0}^{\prime},\ldots,B_{r-1}^{\prime})$ if and only it there exists some $0\not=\alpha\in K$ such that $B_{i}^{\prime}=\alpha^{p^{i}}B_{i}$ for each $i\ 0\leq i<r$.
Under the bicontinuous homeomorphism $\Phi_{\mathfrak{g},r}$ of Lemma 4.7, the projection
$V_{r}(\mathbb{G}){\text{-}}\{0\}\to\mathbb{P}V_{r}(\mathfrak{g})$ corresponds to $\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g})){\text{-}}\{0\}\to\mathbb{P}\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))$ sending $\underline{B}$ to $\langle\underline{B}\rangle$.
We next formulate the projectivization $\tilde{\Theta}_{\mathfrak{g},r}$ of $\Theta_{\mathfrak{g},r}$ and give the analogue of Theorem 2.18 for $\tilde{\Theta}_{\mathfrak{g},r}$. As in Theorem 2.18, we retain the notation and conventions introduced prior to Proposition 2.17 relating graded modules for a graded commutative ring $R$ and sheaves on $\operatorname{Proj}\nolimits R$.
Proposition 4.11.
Let $(\mathbb{G},\mathcal{E}_{(r)})$ be an affine group scheme of exponential type of height $r$. The operator $\Theta_{\mathfrak{g},r}$ of Theorem 4.8 determines a global section
$\tilde{\Theta}_{\mathfrak{g},r}\quad\in\quad\Gamma(\mathbb{P}\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g})),\mathcal{O}_{\mathbb{P}\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))}(p^{r-1})\otimes k\mathbb{G}_{(r)})$
of the coherent, locally free sheaf $\mathcal{O}_{\mathbb{P}\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))}(p^{r-1})\otimes k\mathbb{G}_{(r)}$ of $\mathcal{O}_{\mathbb{P}\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))}$-modules which satisfies the following properties.
  1. (1)
    For any $\mathbb{G}_{(r)}$-module $M$, tensor product with $\tilde{\Theta}_{\mathfrak{g},r}$ determines the map of quasi-coherent $\mathcal{O}_{\mathbb{P}\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))}$-modules
    $\tilde{\Theta}_{\mathfrak{g},r,M}:(k[\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))]\otimes M)^{\sim}\quad\to\quad(k[\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))]\otimes M)^{\sim}(p^{r-1}).$
  2. (2)
    For any non-zero divisior $h\in k[\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))]_{p^{r-1}}$, let $U_{h}$ denote the affine open subset $Spec(k[\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))][h^{-1}])_{0}\ \subset\mathbb{P}\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))$. The natural isomorphism of $(\mathcal{O}_{\mathbb{P}\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))})_{U_{h}}$-modules
    $h^{-1}\otimes 1:(\mathcal{O}_{\mathbb{P}\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))}(p^{r-1})\otimes k\mathbb{G}_{(r)})_{|U_{h}}\quad\stackrel{{\scriptstyle\sim}}{{\to}}\quad(\mathcal{O}_{\mathbb{P}\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))}\otimes k\mathbb{G}_{(r)})_{|U_{h}}$
    sends
    $(\tilde{\Theta}_{\mathfrak{g},r})_{|U_{h}}\ \in\ \Gamma(U_{h},\mathcal{O}_{\mathbb{P}\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))}(p^{r-1})\otimes k\mathbb{G}_{(r)})\ =\ (k[\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))][1/h])_{p^{r-1}}\otimes k\mathbb{G}_{(r)}$
    to
    $(h^{-1}\otimes 1)\cdot\Theta_{\mathfrak{g},r}\ \in\ k[\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))][1/h]\otimes k\mathbb{G}_{(r)}.$
  3. (3)
    If $\underline{B}$ is a scheme theoretic point of $\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))$ projecting to $\langle\underline{B}\rangle$ such that $h(\langle\underline{B}\rangle)\not=0$, then the restriction along $\operatorname{Spec}\nolimits k(\underline{B})\to\mathbb{P}\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))$ of $\tilde{\Theta}_{\mathfrak{g},r}$ is given by
    (4.11.1) $k(\underline{B})\otimes_{\mathcal{O}_{\mathbb{P}\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))}}\tilde{\Theta}_{\mathfrak{g},r}\quad=\quad c\cdot\prod_{s=0}^{r-1}(\mathcal{E}_{r|s,B_{s}})_{*}(1\otimes u_{r-s-1})\ \in\ k(\underline{B})\otimes k\mathbb{G}_{(r)}$
    for some $0\not=c\in k(\underline{B})$.
Proof.
By Proposition 2.17, $\mathcal{O}_{\mathbb{P}\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))}(p^{r-1})\otimes k\mathbb{G}_{(r)}$ is a coherent, locally free sheaf of $\mathcal{O}_{\mathbb{P}\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))}$-modules on $\mathbb{P}\mathcal{C}_{r}$. Proposition 4.5(3) implies that $(\Phi_{r|s})^{*}:k[V_{r}(\mathbb{G})]\to k[\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))]$ is a map of graded $k$-algebras. Thus, Theorem 2.15(5) and Theorem 4.8(1) tell us that $\Theta_{\mathfrak{g},r}\in k[\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))]\otimes k\mathbb{G}_{(r)}$ is homogenous of degree $p^{r-1}$. Therefore, $\Theta_{\mathfrak{g},r}$ determines $\tilde{\Theta}_{\mathfrak{g},r}$ in $\Gamma(\mathcal{O}_{\mathbb{P}\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))},\mathcal{O}_{\mathbb{P}\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))}(p^{r-1})\otimes k\mathbb{G}_{(r)})$.
The proofs of assertions (1), (2), and (3) are verified by evident analogues for $\tilde{\Theta}_{\mathfrak{g},r,}$ of the proofs given in the proof of Theorem 2.18 for the corresponding statements for $\tilde{\Theta}_{\mathbb{G},r,}$
We proceed to simplify $\Theta_{\mathfrak{g},r}$ by replacing the product in (4.11.1) by a sum.
Definition 4.12.
Let $(\mathbb{G},\mathcal{E}_{(r)})$ be an affine group scheme of exponential type of height $r$. We define
$\Theta^{exp}_{\mathfrak{g},r}\ \equiv\ \sum_{s=0}^{r-1}\Theta_{\mathfrak{g},r|s}\quad\in\quad k[\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))]\otimes k\mathbb{G}_{(r)}.$
Thus, for each $\underline{B}\in\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))$
$\Theta^{exp}_{\mathfrak{g},r,\underline{B}}\ \equiv\ \sum_{s=0}^{r-1}\Theta_{\mathfrak{g},r|s,\underline{B}}\ =\ \sum_{s=0}^{r-1}(\mathcal{E}_{r|s,B_{s}})_{*}(1\otimes u_{r-s-1})\quad\in\quad k(\underline{B})\otimes k\mathbb{G}_{(r)}.$
The justification of replacing $\Theta_{\mathfrak{g},r}$ by $\Theta_{\mathfrak{g},r}^{exp}$ is implicit in the following proposition, a reformulation of [12, Prop 4.3] (following a result of Sobaje [24, Prop 2.3]).
Proposition 4.13.
Let $(\mathbb{G},\mathcal{E}_{(r)})$ be an affine group scheme of exponential type of height $r$, let $M$ be a finite dimensional $\mathbb{G}_{(r)}$-module, and let $\underline{a}=\sum_{i=1}^{p}a_{i}[i]$ be a Jordan type which is maximal among the set of Jordan types $\{JT(\Theta_{\mathfrak{g},r,\underline{B}},k(\underline{B})\otimes M);\ \underline{B}\in\mathcal{C}_{r}(\mathbb{G})\}$ for $M$. For any scheme-theoretic point $\underline{B}\in\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))$,
$JT(\Theta^{exp}_{\mathfrak{g},r,\underline{B}},k(\underline{B})\otimes M)\ =\ \underline{a}\quad{\text{i}f\ and\ only\ if}\quad JT(\Theta_{\mathfrak{g},r,\underline{B}},k(\underline{B})\otimes M)\ =\ \underline{a}.$
Proof.
In [24, Prop 2.3], Sobaje regards $\Theta_{\mathfrak{g},r,\underline{B}}\ =\prod_{s=0}^{r-1}\Theta_{\mathfrak{g},r|s,\underline{B}}$ as the composition
$(\prod_{s=0}^{r-1}\circ(\otimes_{s=0}^{r-1}(\mathcal{E}_{r|s,B_{s}})_{*})\circ(1\otimes\Delta):\ k(\underline{B})\otimes k\mathbb{G}_{a(r)}\ \to\ k(\underline{B})\otimes k\mathbb{G}_{a(r)}^{\otimes r}$
$\quad\to\ k(\underline{B})\otimes k\mathbb{G}_{(r)}^{\otimes r}\to k(\underline{B})\otimes k\mathbb{G}_{(r)}$
applied to $1\otimes u_{r-1}\in k(\underline{B})\otimes k\mathbb{G}_{a(r)}$. Using this composition, he verifies that $\Theta_{\mathfrak{g},r,\underline{B}}$ equals $\sum_{s=0}^{r-1}\Theta_{\mathfrak{g},r|s,\underline{B}}$ plus other terms which are scalar multiples of products of at least two factors of this sum, factors which are $p$-nilpotent and pair-wise commuting. The proposition then follows from [17, Thm 1.12] (see also [8, Prop 8] and [27, Lem 6.4]). ∎
As shown in Theorem 2.23 and remarked in Remark 2.26, the support variety and its generalizations for a finite dimensional $\mathbb{G}_{(r)}$-module $M$ are determined by $JT_{\mathbb{G},r,M}(-)$ and thus by $JT_{\mathfrak{g},r,M}(-)$. (See Corollary 4.9.)
Since these support varieties are formulated in terms of maximality of Jordan types, Proposition 4.13 has the following corollary.
Corollary 4.14.
Let $(\mathbb{G},\mathcal{E}_{(r)})$ be an affine group scheme of exponential type of height $r$ and let $M$ be a finite dimensional $\mathbb{G}_{(r)}$-module. Then the support variety and generalized support varieties of $M$ interpreted as closed subvarieties of $\mathcal{C}_{r}(\mathfrak{g})$ using $JT_{\mathbb{G},r,M}(-)$ are left unchanged by replacing $JT_{\mathbb{G},r,M}(-)$ by $JT^{exp}_{\mathbb{G},r,M}(-)$.
Remark 4.15.
By definition, $\Theta^{exp}_{\mathfrak{g},1}\ =\ \Theta_{\mathfrak{g},1}$. As observed in [24, Prop 2.3] (see also the proof of [27, Prop 6.5]), $\Theta^{exp}_{\mathfrak{g},2}\ =\ \Theta_{\mathfrak{g},2}$. On the other hand, for $r>2$, $\Theta^{exp}_{\mathfrak{g},r}\ \not=\ \Theta_{\mathfrak{g},r}$.
We next give the evident definition of the “exponential Jordan type function” of a $\mathbb{G}_{(r)}$-module, parallel to the definition of $JT_{\mathbb{G},r,M}(-)$ given in Definition 2.16.
Definition 4.16.
Let $(\mathbb{G},\mathcal{E}_{(r)})$ be an affine group scheme of exponential type of height $r$ and let $M$ be a finite dimensional $\mathbb{G}_{(r)}$-module. We define
$JT^{exp}_{\mathfrak{g},r,M}(-):\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))\quad\to\quad\mathcal{Y}$
by sending a geometric point $\underline{B}:\operatorname{Spec}\nolimits k(\underline{B})\to\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))$ to
(4.16.1) $JT(\Theta^{exp}_{\mathfrak{g},r,M})(\underline{B})\quad=\quad JT(\sum_{s=0}^{r-1}(\mathcal{E}_{r|s,B_{s}})_{*}(1\otimes u_{r-s-1}),k(\underline{B})\otimes M))\ ,$
the Jordan type of the operator $\Theta^{exp}_{\mathfrak{g},r,\underline{B}}$ on $k(\underline{B})\otimes M$.
Observe that the operator $(\mathcal{E}_{r|s,B_{s}})_{*}(1\otimes u_{r-s-1})$ on $k(\underline{B})\otimes M$ can be identified with the operator $1\otimes u_{r-s-1}\in k(\underline{B})\otimes k\mathbb{G}_{a(r)}$ on $(\mathcal{E}_{r|s,B_{s}})^{*}(k(\underline{B})\otimes M)$.
We next formulate the projectivization $\tilde{\Theta}_{\mathfrak{g},r}^{exp}$ of $\Theta_{\mathfrak{g},r}^{exp}$ and give the analogues of Theorem 2.18 for $\tilde{\Theta}_{\mathbb{G},r}$ and Proposition 4.11 for $\tilde{\Theta}_{\mathfrak{g},r}$.
Proposition 4.17.
Let $(\mathbb{G},\mathcal{E}_{(r)})$ be an affine group scheme of exponential type of height $r$. The operator $\Theta_{\mathfrak{g},r}^{exp}$ of Definition 4.12 determines a global section
$\tilde{\Theta}_{\mathfrak{g},r}^{exp}\quad\in\quad\Gamma(\mathcal{O}_{\mathbb{P}\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))},\mathcal{O}_{\mathbb{P}\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))}(p^{r-1})\otimes k\mathbb{G}_{(r)})$
of the coherent, locally free sheaf $\mathcal{O}_{\mathbb{P}\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))}(p^{r-1})\otimes k\mathbb{G}_{(r)}$ of $\mathcal{O}_{\mathbb{P}\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))}$-modules which satisfies the following properties.
  1. (1)
    For any $\mathbb{G}_{(r)}$-module $M$, tensor product with $\tilde{\Theta}_{\mathfrak{g},r}^{exp}$ determines the map of quasi-coherent $\mathcal{O}_{\mathbb{P}\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))}$-modules
    $\tilde{\Theta}_{\mathfrak{g},r,M}^{exp}:(k[\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))]\otimes M)^{\sim}\quad\to\quad(k[\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))]\otimes M)^{\sim}(p^{r-1}).$
  2. (2)
    For any non-zero divisior $h\in k[\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))]_{p^{r-1}}$, let $U_{h}$ denote the affine open subset $Spec(k[\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))][h^{-1}])_{0}\ \subset\mathbb{P}\mathcal{C}_{r}(\mathfrak{g})$. The natural isomorphism of $(\mathcal{O}_{\mathbb{P}\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))})_{U_{h}}$-modules
    $h^{-1}\otimes 1:(\mathcal{O}_{\mathbb{P}\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))}(p^{r-1})\otimes k\mathbb{G}_{(r)})_{|U_{h}}\quad\stackrel{{\scriptstyle\sim}}{{\to}}\quad(\mathcal{O}_{\mathbb{P}\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))}\otimes k\mathbb{G}_{(r)})_{|U_{h}}$
    sends
    $(\tilde{\Theta}_{\mathfrak{g},r}^{exp})_{|U_{h}}\ \in\ \Gamma(U_{h},\mathcal{O}_{\mathbb{P}\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))}(p^{r-1})\otimes k\mathbb{G}_{(r)})\ =\ (k[\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))][1/h])_{p^{r-1}}\otimes k\mathbb{G}_{(r)}$
    to
    $(h^{-1}\otimes 1)\cdot\Theta_{\mathfrak{g},r}^{exp}\ \in\ k[\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))][1/h]\otimes k\mathbb{G}_{(r)}.$
  3. (3)
    If $\underline{B}$ is a scheme theoretic point of $\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))$ projecting to $\langle\underline{B}\rangle$ such that $h(\langle\underline{B}\rangle)\not=0$, then the restriction along $\operatorname{Spec}\nolimits k(\underline{B})\to\mathbb{P}\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))$ of $\tilde{\Theta}_{\mathfrak{g},r}^{exp}$ is given by
    (4.17.1) $k(\underline{B})\otimes_{\mathcal{O}_{\mathbb{P}\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))}}\tilde{\Theta}_{\mathfrak{g},r}^{exp}\quad=\quad c\cdot\sum_{s=0}^{r-1}(\mathcal{E}_{(r),B_{s}})_{*}(1\otimes u_{r-s-1})\ \in\ k(\underline{B})\otimes k\mathbb{G}_{(r)}$
    for some $0\not=c\in k(\underline{B})$.
  4. (4)
    For any $\mathbb{G}_{(r)}$-module $M$, the restriction of $\tilde{\Theta}_{\mathfrak{g},r,M}^{exp}$ along $\operatorname{Spec}\nolimits k(\underline{B})\to\mathbb{P}\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))$ as in (3) is a non-zero scalar multiple of the image of
    $\sum_{s=0}^{r-1}(\mathcal{E}_{r|s,B_{s}})_{*}(1\otimes u_{r-s-1})$ in $End_{k(\underline{B})}(k(\underline{B})\otimes M)$.
Proof.
By Proposition 2.17, $\mathcal{O}_{\mathbb{P}\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))}(p^{r-1})\otimes k\mathbb{G}_{(r)}$ is a coherent, locally free sheaf of $\mathcal{O}_{\mathbb{P}\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))}$-modules on $\mathbb{P}\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))$. Proposition 4.5(3) implies that $(\Phi_{r|s})^{*}:k[V_{r}(\mathbb{G})]\to k[\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))]$ is a map of graded $k$-algebras so that $\Theta_{\mathfrak{g},r}^{exp}$ is homogenous by Theorem 2.15(5) and Theorem 4.8(1). Thus, $\Theta_{\mathfrak{g},r}^{exp}$ determines $\tilde{\Theta}_{\mathfrak{g},r,M}^{exp}$ in
$\Gamma(\mathcal{O}_{\mathbb{P}\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))},\mathcal{O}_{\mathbb{P}\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))}(p^{r-1})\otimes k\mathbb{G}_{(r)})$.
The proofs of assertions (1) - (4) are verified by evident analogues for $\tilde{\Theta}^{exp}_{\mathfrak{g},r,}$ of the proofs given in the proof of Theorem 2.18 for the corresponding statements for $\tilde{\Theta}_{\mathbb{G},r,}$
[…]
Theorem 4.18.
Let $(\mathbb{G},\mathcal{E}_{(r)})$ be an affine group scheme of exponential type of height $r$. For any finite dimensional $\mathbb{G}_{(r)}$-module $M$ and any scheme-theoretic point $\underline{B}$ of $\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))$ projecting to $\langle\underline{B}\rangle\in\mathbb{P}\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))$, we define
$JT_{\mathfrak{g},r,M}^{exp}(\langle\underline{B}\rangle)\quad\equiv\quad JT(\sum_{s=0}^{p-1}(\mathcal{E}_{r|s,B_{s}})_{*}(1\otimes u_{r-s-1}),k(\underline{B})\otimes M),$
the Jordan type of the $p$-nilpotent element $k(\underline{B})\otimes_{\mathcal{O}_{\mathbb{P}\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))}}\tilde{\Theta}_{\mathfrak{g},r}^{exp}\in k(\underline{B})\otimes\mathbb{G}_{(r)}$ acting on $k(\underline{B})\otimes M$.
  1. (1)
    For any finite dimensional $\mathbb{G}_{(r)}$-module $M$, the association $\langle\underline{B}\rangle\ \mapsto JT_{\mathfrak{g},r,M}^{exp}(\langle\underline{B}\rangle)$ determines a continuous map
    $JT_{\mathfrak{g},r,M}^{exp}(-):\mathbb{P}\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))\quad\to\quad\mathcal{Y}.$
  2. (2)
    For $\mathbb{G}_{(r)}$-modules $M,N$ of dimensions $m,n$ and any scheme-theoretic point $\underline{B}\in\mathbb{P}\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))$,
    $JT_{\mathfrak{g},r,M\oplus N}^{exp}(\langle\underline{B}\rangle)\ =\ JT_{\mathfrak{g},r,M}^{exp}(\langle\underline{B}\rangle)+JT_{\mathfrak{g},r,N}^{exp}(\langle\underline{B}\rangle)\ \in\ \mathcal{Y}_{m+n}.$
  3. (3)
    Assume either that $r=1$ or that $\langle\underline{B}\rangle\in\mathbb{P}\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))$ is a generic point. For $\mathbb{G}_{(r)}$-modules $M,N$ of dimensions $m,n$,
    $JT_{\mathfrak{g},r,M\otimes N}^{exp}(\langle\underline{B}\rangle)\ =\ JT_{\mathfrak{g},r,M}^{exp}(\langle\underline{B}\rangle)\otimes JT_{\mathfrak{g},r,N}^{exp}(\langle\underline{B}\rangle)\ \in\ \mathcal{Y}_{m\cdot n}.$
  4. (4)
    If $f:M\to N$ is a surjective map of $\mathbb{G}_{(r)}$-modules, then $JT_{\mathfrak{g},r,N}^{exp}(\langle\underline{B}\rangle)\ \leq JT_{\mathfrak{g},r,M}^{exp}(\langle\underline{B}\rangle)$ for any $\langle\underline{B}\rangle\in\mathbb{P}\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))$; in other words, $JT_{\mathfrak{g},r,N}^{exp}(-)\leq JT_{\mathfrak{g},r,M}^{exp}(-)$.
  5. (5)
    A short exact sequence $0\to N\ \to\ M\ \to\ Q\to 0$ of finite dimensional $\mathbb{G}_{(r)}$-modules is locally split if and only if $JT_{\mathfrak{g},r,M}^{exp}(-)=JT_{\mathfrak{g},r,N}^{exp}(-)+JT_{\mathfrak{g},r,Q}^{exp}(-).$
  6. (6)
    If $\phi:\mathbb{H}\to\mathbb{G}$ is a morphism of affine group schemes of exponential type of height $r$, then for any finite dimensional $\mathbb{G}_{(r)}$-module $M$
    $JT_{\mathfrak{h},r,\phi*(M)}^{exp}(-):\mathbb{P}\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{h}))\ \to\ \mathcal{Y}$
    equals
    $JT_{\mathfrak{g},r,M}^{exp}(-)\circ\phi_{*}:\mathbb{P}\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{h}\ \to\ \mathbb{P}\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))\ \to\ \mathcal{Y}.$
5. Examples of exponential Jordan types
We provide some examples of computations of the simplified Jordan type function $JT^{exp}_{\mathfrak{g},r,M}(-)$ of Theorem 4.18. One challenge to computing Jordan types is the behavior of Jordan types with respect to the sum of two commuting, $p$-nilpotent operators. Another challenge is the fact that for $r>1$ the Jordan type functions $JT_{\mathbb{G},r,M}(-),\ JT_{\mathfrak{g},r,M}^{exp}(-)$ rarely commute with tensor products of $\mathbb{G}_{(r)}$-modules at all points of $\mathbb{P}\mathcal{C}_{r}(\mathfrak{g})$.
Proposition 5.1.
By Proposition 4.13, Examples 3.1 and 3.2 remain valid if $(\mathbb{G},\mathcal{E}_{(r)})$ is an affine group scheme of exponential type of height $r$ and if $JT_{\mathbb{G},r,M}(-)$ is replaced by $JT^{exp}_{\mathfrak{g},r,M}(-)$. The justification of Example 3.3 also justifies its validity when $JT_{\mathbb{G},r,M}(-)$ is replaced by $JT^{exp}_{\mathfrak{g},r,M}(-)$.
Example 3.6 remains unchanged, since $JT_{\mathbb{G},1,M}(-)\ =\ JT^{exp}_{\mathfrak{g},1,M}(-)$.
On the other hand, as we see in the following examples, some computations simplify if we replace $JT_{\mathbb{G}_{(r)},M}(-)$ by $JT^{exp}_{\mathfrak{g},M}(-)$.
Example 5.2.
Adopt the notation of Examples 2.10 and 2.14. Let $V$ be an $n$-dimensional $k$-vector space, and let $\alpha_{0},\ldots\alpha_{r-1}$ be be pair-wise commuting $p$-nilpotent endomorphisms of $V$. Denote by $M$ the $(\mathbb{G}_{a}^{\times r})_{1}$-module such that $x_{i}\in k(\mathbb{G}_{a}^{r})_{1}$ acts as $\alpha_{i}$ and denote by $N$ the $\mathbb{G}_{a(r)}$-module such that $u_{i}\in k\mathbb{G}_{a(r)}$ also acts as $\alpha_{i}$.
Recall from Example 2.10 the morphism $\Phi:\mathbb{P}^{r-1}\to w\mathbb{P}(1,p,\ldots,p^{r-1})$. Since $\Theta_{\mathbb{G}_{a,r}}^{exp}\ =\ \sum_{i=0}^{r-1}T_{i}^{p^{r-1-i}}\otimes u_{i}\ \in\ k[V_{r}(\mathbb{G}_{a(r)})]\otimes k\mathbb{G}_{a(r)}$,
$\Phi^{*}(\Theta_{\mathbb{G}_{a},r}^{exp})\quad=\quad\sum_{i=0}^{r-1}t_{i}^{p^{r-1}}\otimes x_{i}\ =\ F^{r-1*}(\Theta_{\mathbb{G}_{a}^{\times r},1}).$
Thus,
$JT_{\mathfrak{g}_{a},r,N}^{exp}(-)\circ\Phi\ =\ JT_{\mathbb{G}_{a}^{\times r},1,M}\circ F^{r-1}:\ \mathbb{P}^{r-1}\ \to\ \mathcal{Y}.$
In the following examples, we use (4.16.1) of Definition 4.16 which becomes
$JT^{exp}_{\mathfrak{g},r,M}(\underline{B})\quad=\quad JT(\sum_{s=0}^{r-1}(exp_{B_{s}})_{*}(1\otimes u_{r-s-1}),k(\underline{B}))$
provided that $(\mathbb{G},\mathcal{E}_{r})$ is associated with an embedding $i:\mathbb{G}\hookrightarrow GL_{N}$ of exponential type of height $r$.
Example 5.3.
Set $\mathbb{G}=SL_{2(2)}$, so that $\mathfrak{g}=sl_{2}$. We consider the irreducible $\mathbb{G}$-module
$M\ =\ S(\lambda_{0})\otimes S(\lambda_{1})^{(1)},\quad 0\leq\lambda_{0},\lambda_{1}<p,\ \lambda=\lambda_{0}+p\lambda_{1}.$
We proceed to compute $JT^{exp}_{sl_{2},2,M}(-):\mathcal{C}_{2}(\mathcal{N}_{p}(sl_{2}))\to\mathcal{Y}_{\leq}$ as in Definition 4.16. To compute $JT_{SL_{2},2,M}(\underline{B})$ for $B_{0},B_{1}$ commuting elements in $sl_{2}$, we appeal to Proposition 2.20 to reduce to the case that $\underline{B}=(a_{0}E,a_{1}E)$ where $E$ is the $2\times 2$ matrix with $E_{1,2}=1$ and $E_{1,1}=E_{2,1}=E_{2,2}=0$. We may then view $\mathcal{E}_{(2),B_{s}}$ as factoring as $\mathbb{G}_{a(2),k(\underline{B})}\to\mathbb{G}_{a(2),k(\underline{B})}\subset SL_{2(2),k(\underline{B})}$.
To compute $JT_{sl_{2},2,S(\lambda)}^{exp}(\underline{B})$, we identify the Jordan type of
$(\mathcal{E}_{(2),B_{0}})_{*}(u_{1})+(\mathcal{E}_{(2),B_{1}})_{*}(u_{0})=a_{0}E\cdot u_{1}+a_{1}^{p}E\cdot u_{0}\in k(\underline{B})\otimes kSL_{2(2)}$
on $S(\lambda_{0})\otimes S(\lambda_{1})^{(1)}$. Here, the action of $a_{0}E\cdot u_{1}$ on $S(\lambda_{0})$ is zero, whereas the action of $a_{1}^{p}E\cdot u_{0}$ on $S(\lambda_{0})$ is $a_{1}^{p}$ times the action of $E\in sl_{2}$ on $S(\lambda_{0})$; the action of $a_{0}E\cdot u_{1}$ on $S(\lambda_{1})^{(1)}$ is $a_{0}$ times the action of $E\in sl_{2}$ on $S(\lambda_{1})$ and the action of $a_{1}^{p}E\cdot u_{0}$ on $S(\lambda_{1})^{(1)}$ is zero.
Observe that the Jordan type of $E$ on $S(\lambda_{0})=[m]$ where $m=\lambda_{0}+1$ and on $S(\lambda_{1})=[n]$ where $n=\lambda_{1}+1$. Thus, if $a_{0}\not=0,\ a_{1}=0$, then the Jordan type $S(\lambda)$ is $m\cdot[n]$; and if $a_{0}=0,\ a_{1}\not=0$, then the Jordan type $S(\lambda)$ is $n\cdot[m]$. More generally, $S(\lambda_{0})\otimes S(\lambda_{1})^{(1)}$ as a $k(\underline{B})\otimes k[u]/u^{p}$-module with $1\otimes u$ acting as $a_{0}E\cdot u_{1}+a_{1}^{p}E\cdot u_{0}$ can be identified with a tensor product of two $k[u]/u^{p}$-modules: the first with with $u$ acting as $a_{1}^{p}E\cdot u_{0}$ on $S(\lambda_{0})$, the second with $u$ acting as $a_{0}E\cdot u_{1}$ on $S(\lambda_{1})^{(1)}$. Consequently, if both $a_{0},\ a_{1}$ are non-zero, then the the Jordan type of $a_{1}^{p}E\cdot u_{0}+a_{0}E\cdot u_{1}$ on $S(\lambda)$ equals $m]\otimes[n]$.
Example 5.4.
Let $V$ be a polynomial representation of $GL_{N}$ of degree <p<p and consider the $i$-th Frobenius twist $V^{(i)}$ for some $i\geq 0$. We fix some $r>i$ and proceed to determine $JT_{\mathfrak{g}l_{N},r,V^{(i)}}(\underline{B})$ for $\underline{B}=(B_{0},\ldots,B_{r-1})$, an $r$-tuple of $p$-nilpotent, pairwise commuting elements of $k(\underline{B})\otimes\mathfrak{g}l_{N}$. Then the action of $(\exp_{B_{s}})_{*}(u_{r-s-1})$ on $V^{(i)}$ is non-zero if and only if i=rs1i=r-s-1, in which case this action equals the action of $B_{s}\in k(\underline{B})\otimes\mathfrak{g}l_{N}$ acting on $k(\underline{B})\otimes V$. Consequently, $JT_{\mathfrak{g}l_{N},r,V^{(i)}}^{exp}(\underline{B})=JT(B_{s},V)$ where s=ri1s=r-i-1.
More generally, we can take a polynomial representation $V_{s}$ of $GL_{N}$ for each $s,\ 0\leq s<r$ with each $V_{s}$ of degree <p<p and consider $M=\otimes_{s=1}^{r-1}V_{s}^{(s)}$. Let $\underline{B}$ be a geometric point of $\mathcal{C}_{r}(\mathfrak{g}l_{N})$. The action of $u=\sum_{s=0}^{r-1}(exp_{B_{s}})_{*}(1\otimes u_{r-s-1})$ on $k(\underline{B})\otimes M$ is the tensor product of the actions of $B_{s}\in k(\underline{B})\otimes\mathfrak{g}l_{N}$ acting on $k(\underline{B})\otimes V_{s}$.
Thus, $JT_{\mathfrak{g}l_{N},r,M}^{exp}(\underline{B})\ =\ \bigotimes_{s=0}^{r-1}JT(B_{r-s-1},V_{s}).$
Example 5.5.
We employ Theorem 3.7 of [16] to consider the general situation of $(\mathbb{G},\mathcal{E}_{(r)})$ an affine group scheme of exponential type of height $r$ and $M$ a $\mathbb{G}$ module of dimension $m$. Namely, if $\rho:\mathbb{G}\to GL_{m}$ determines $M$, then $(\rho\circ\mathcal{E}_{(r),B_{s}})_{*}(1\otimes u_{r-s-1})$ acting on $k(\underline{B})\otimes M$ is given by the $m\times m$ matrix whose $(i,j)$-entry equals the coefficient of $t^{p^{r-s-1}}\in k(\underline{B})\otimes\mathbb{G}_{a(r)}$ of $(\mathcal{E}_{r|s,B_{s}})_{*}(1\otimes u_{r-s-1})$ applied to $\rho^{*}(X_{i,j}).$
Thus $JT^{exp}_{\mathfrak{g},r,M}(\underline{B})$ is the Jordan type of the sum (indexed by $s,0\leq s<r$) of these $m\times m$ matrices.
6. Vector bundles and stratifications
In this section, we revisit the construction of vector bundles associated to $\mathbb{G}_{(r)}$-modules of constant Jordan type (and more general constant $j$-type). We refer to [7] for an introduction to modules of constant Jordan type and for numerous examples of these interesting $\mathbb{G}$-modules. In [16], such vector bundles were associated to $\mathbb{G}_{(r)}$-modules of constant Jordan type in order to provide invariants which can distinguish some $\mathbb{G}_{(r)}$-modules whose Jordan type functions are equal.
As we point out, the constructions of [16] when applied to an arbitrary $\mathbb{G}_{(r)}$-module $M$ yield coherent sheaves on $\mathbb{P}V_{r}(\mathbb{G})$ whose restrictions to each stratum associated to $JT_{\mathbb{G},r,M}(-)$ are locally free.
Definition 6.1.
Let $\mathbb{G}$ be an affine group scheme of finite type over $k$, let $r$ be a chosen positive integer, $j$ and integer with $1\leq j<r$, and let $M$ be a $\mathbb{G}_{(r)}$-module. We define the following coherent sheaves on $\mathbb{P}V_{r}(\mathbb{G})$
$\mathcal{K}\text{\it er}\{\tilde{\Theta}_{\mathbb{G},r,M}^{j}\},\quad\mathcal{C}\text{\it oker}\{\tilde{\Theta}_{\mathbb{G},r,M}^{j}\},\quad\mathcal{I}\text{\it m}\{\tilde{\Theta}_{\mathbb{G},r,M}^{j}\}$
as the sheaves associated to kernel, cokernel, and image of the $j$-th iterate of the map of graded $k[V_{r}(\mathbb{G})]$-modules
$\tilde{\Theta}_{\mathbb{G},r,M}:(k[V_{r}(\mathbb{G})]\otimes M)^{\sim}\quad\to\quad(k[V_{r}(\mathbb{G})]\otimes M)^{\sim}(p^{r-1})$
given in Theorem 2.18 (1).
Definition 6.2.
Let $(\mathbb{G},\mathcal{E}_{(r)})$ be an affine group scheme of exponential type of height $r$, $j$ and integer with $1\leq j<r$, and let $M$ be a $\mathbb{G}_{(r)}$-module. We define the following coherent sheaves on $\mathbb{P}\mathcal{C}_{r}(\mathfrak{g})$
$\mathcal{K}\text{\it er}\{\tilde{\Theta}_{\mathfrak{g},r,M}^{j}\},\quad\mathcal{C}\text{\it oker}\{\tilde{\Theta}_{\mathfrak{g},r,M}^{j}\},\quad\mathcal{I}\text{\it m}\{\tilde{\Theta}_{\mathfrak{g},r,M}^{j}\}$
$\mathcal{K}\text{\it er}\{(\tilde{\Theta}_{\mathfrak{g},r,M}^{exp})^{j}\},\quad\mathcal{C}\text{\it oker}\{(\tilde{\Theta}_{\mathbb{G},r,M}^{exp})^{j}\},\quad\mathcal{I}\text{\it m}\{(\tilde{\Theta}_{\mathbb{G},r,M}^{exp})^{j}\}$
as the sheaves associated to kernel, cokernel, and image of the $j$-th iterate of the maps $\tilde{\Theta}_{\mathfrak{g},r,M}$ and $\tilde{\Theta}_{\mathfrak{g},r,M}^{exp}$ of graded $k[\mathcal{C}_{r}(\mathfrak{g})]$-modules given in Proposition 4.11 (1) and Proposition 4.17 (1).
We next verify that the coherent sheaves of Definition 6.2 are locally constant on each stratum.
Theorem 6.3.
Let $\mathbb{G}$ be an affine group scheme of finite type of $k$ and let $M$ be a finite dimensional $\mathbb{G}_{(r)}$-module. Then for each Jordan type $\underline{a}\ =\ \sum_{i=1}^{p}a_{i}[i]$, the restrictions of the coherent sheaves $\mathcal{K}\text{\it er}\{\tilde{\Theta}_{\mathbb{G},r,M}^{j}\},\ \mathcal{C}\text{\it oker}\{\tilde{\Theta}_{\mathbb{G},r,M}^{j}\},\ \mathcal{I}\text{\it m}\{\tilde{\Theta}_{\mathbb{G},r,M}^{j}\}$ to the locally closed subvariety $JT_{\mathbb{G},r,M}^{-1}(\underline{a})\ \subset\ \mathbb{P}V_{r}(\mathbb{G})$ are locally free coherent sheaves on $JT_{\mathbb{G},r,M}^{-1}(\underline{a})$.
If $(\mathbb{G},\mathcal{E}_{(r)})$ is an affine group scheme of exponential type of height $r$, then the pull-backs along $\Phi_{\mathfrak{g},r}:JT_{\mathfrak{g},r,M}^{-1}(\underline{a})\to JT_{\mathbb{G},r,M}^{-1}(\underline{a})$ of the sheaves $ker\{\tilde{\Theta}_{\mathbb{G},r,M}^{j}\},\\ coker\{\tilde{\Theta}_{\mathbb{G},r,M}^{j}\},\ im\{\tilde{\Theta}_{\mathbb{G},r,M}^{j}\}$ are locally free sheaves on $JT_{\mathfrak{g},r,M}^{-1}(\underline{a})\ \subset\ \mathbb{P}\mathcal{C}_{r}(\mathfrak{g})$.
Moreover, if $(\mathbb{G},\mathcal{E}_{(r)})$ is an affine group scheme of exponential type of height $r$, then the restrictions of the coherent sheaves $\mathcal{K}\text{\it er}\{(\tilde{\Theta}_{\mathfrak{g},r,M}^{exp})^{j}\},\ \mathcal{C}\text{\it oker}\{(\tilde{\Theta}_{\mathfrak{g},r,M}^{exp})^{j}\},\\ \mathcal{I}\text{\it m}\{(\tilde{\Theta}_{\mathfrak{g},r,M}^{exp})^{j}\}$ to $(JT_{\mathfrak{g},r,M}^{exp})^{-1}(\underline{a})\ \subset\ \mathbb{P}\mathcal{C}_{r}(\mathfrak{g})$ are locally free sheaves for any Jordan type $\underline{a}\ =\ \sum_{i=1}^{p}a_{i}[i]$.
Proof.
Let $Y_{\underline{a}}$ denote $(JT_{\mathfrak{g},r,M}^{exp})^{-1}(\underline{a})\ \subset\ \mathbb{P}V_{r}(\mathfrak{g})$ and consider its closure
$\overline{Y_{\underline{a}}}\ \equiv\ JT_{\mathfrak{g},r,M}^{-1}(\overline{\{\underline{a}\}})\ =\ \bigcup_{\underline{b}\leq\underline{a}}JT_{\mathfrak{g},r,M}^{-1}(\underline{b})\quad\hookrightarrow\quad\mathbb{P}V_{r}(\mathbb{G})$
corresponding to the homogeneous ideal $I_{\underline{a}}\ \subset\ k[V_{r}(\mathbb{G})]$. Consider $k[\overline{Y_{\underline{a}}}]\ =\ k[V_{r}(\mathbb{G})]/I_{\underline{a}}$, equipped with the grading inherited from the grading of $k[V_{r}(\mathbb{G})]$ and set
$\Theta_{\mathbb{G},r;\overline{Y_{\underline{a}}}}\ \equiv\ k[\overline{Y_{\underline{a}}}]\otimes_{k[V_{r}(\mathbb{G})]}\Theta_{\mathbb{G},r}.$
So defined, $\Theta_{\mathbb{G},r;\overline{Y_{\underline{a}}}}$ determines a $p$-nilpotent $k[\overline{Y_{\underline{a}}}]$-linear operator $\Theta_{\mathbb{G},r,M:\overline{Y_{\underline{a}}}}$ on $k[\overline{Y_{\underline{a}}}]\otimes M$ as in Theorem 2.15. As in Theorem 2.18, $\Theta_{\mathbb{G},r;\overline{Y_{\underline{a}}}}$ determines
$\tilde{\Theta}_{\mathbb{G},r;\overline{Y_{\underline{a}}}}\ \in\ \Gamma(\overline{Y_{\underline{a}}},\mathcal{O}_{\overline{Y_{\underline{a}}}}(p^{r-1})\otimes k\mathbb{G}_{(r)})$
and the map of graded $k[\overline{Y_{\underline{a}}}]$-modules
$\tilde{\Theta}_{\mathbb{G},r,M;\overline{Y_{\underline{a}}}}:k[\overline{Y_{\underline{a}}}]\otimes M)^{\sim}\ \to\ (k[\overline{Y_{\underline{a}}}]\otimes M)^{\sim}(p^{r-1}).$
Then the restrictions of $\mathcal{K}\text{\it er}\{\tilde{\Theta}_{\mathbb{G},r,M}^{j}\},\ \mathcal{C}\text{\it oker}\{\tilde{\Theta}_{\mathbb{G},r,M}^{j}\},\ \mathcal{I}\text{\it m}\{\tilde{\Theta}_{\mathbb{G},r,M}^{j}\}$ along $Y_{\underline{a}}\subset\overline{Y_{\underline{a}}}\ \hookrightarrow\ \mathbb{P}V_{r}(\mathbb{G})$ are identified with the restrictions to the open subset $Y_{\underline{a}}\subset\overline{Y_{\underline{a}}}$ of $\mathcal{K}\text{\it er}\{\tilde{\Theta}_{\mathbb{G},r,M;\overline{Y_{\underline{a}}}}^{j}\},\ \mathcal{C}\text{\it oker}\{\tilde{\Theta}_{\mathbb{G},r,M;\overline{Y_{\underline{a}}}}^{j}\},\ \mathcal{I}\text{\it m}\{\tilde{\Theta}_{\mathbb{G},r,M;\overline{Y_{\underline{a}}}}^{j}\}$.
The proof that the restrictions to $Y_{\underline{a}}$ of $\mathcal{K}\text{\it er}\{\tilde{\Theta}_{\mathbb{G},r,M;\overline{Y_{\underline{a}}}}^{j}\},\ \mathcal{C}\text{\it oker}\{\tilde{\Theta}_{\mathbb{G},r,M;\overline{Y_{\underline{a}}}}^{j}\}$ and $\mathcal{I}\text{\it m}\{\tilde{\Theta}_{\mathbb{G},r,M;\overline{Y_{\underline{a}}}}^{j}\}$ are locally free coherent sheaves on $Y_{\underline{a}}$ is completed by repeating the proof of [16, Thm 4.13] with $U\subset\mathbb{P}(G),\ \tilde{\Theta}_{G}$ of that theorem replaced by $Y_{\underline{a}}\ \subset\ \overline{Y_{\underline{a}}},\ \tilde{\Theta}_{\mathbb{G},r;\overline{Y_{\underline{a}}}}$ as above.
If $(\mathbb{G},\mathcal{E}_{(r)})$ is an affine group scheme of exponential type of height $r$, then the pull-backs of these locally free sheaves along along the homeomoprhism $\Phi_{\mathfrak{g},r}:JT_{\mathfrak{g},r,M}^{-1}(\underline{a})\to JT_{\mathbb{G},r,M}^{-1}(\underline{a})$ are also locally free. Finally, only notational changes replacing $\Theta_{\mathfrak{g},r,M}$ by $\Theta_{\mathfrak{g},r,M}^{exp}$ are needed to obtains proofs of the assertions for $\mathcal{K}\text{\it er}\{(\tilde{\Theta}_{\mathfrak{g},r,M}^{exp})^{j}\},\ \mathcal{C}\text{\it oker}\{(\tilde{\Theta}_{\mathfrak{g},r,M}^{exp})^{j}\},\mathcal{I}\text{\it m}\{(\tilde{\Theta}_{\mathfrak{g},r,M}^{exp})^{j}\}$. ∎
Example 3.1 gives examples of $\mathbb{G}_{(r)}$ modules leading to a stratification of $\mathbb{P}V_{r}(\mathbb{G})$ with a single stratum and Example 3.2 gives examples of $\mathbb{G}_{(r)}$-modules with exactly two strata.
We give one “trivial” example of a $\mathbb{G}_{(r)}$ module with more than two strata, primarily to show that some examples can be easily computed.
Example 6.4.
Let $\mathbb{G}=\mathbb{G}_{a}^{\times p}$, so that $k\mathbb{G}_{(1)}=k[u_{0},\dots,u_{p-1}]$. Consider the $k\mathbb{G}_{(1)}$-module $M$ defined as the quotient of $k\mathbb{G}_{(1)}$ by the ideal generated by $\{u_{i}^{i}\}$. Consider points $\underline{b}=\langle b_{0},\ldots,b_{p-1}\rangle\in\mathbb{P}V_{1}(\mathbb{G})\simeq\mathbb{P}^{p-1}$. If $\underline{b}$ is given by $b_{i}=1,\ b_{j}=0,j\not=i$, then $JT_{\mathbb{G},1,M}(\underline{b})$ equals $\sum_{j\not=i}[j]+i[1]$.
We next investigate the relationship between the coherent sheaves $\mathcal{K}\text{\it er}\{\tilde{\Theta}_{\mathfrak{g},r,M}^{j}\},\\ \mathcal{C}\text{\it oker}\{\tilde{\Theta}_{\mathbb{G},r,M}^{j}\},\ \mathcal{I}\text{\it m}\{\tilde{\Theta}_{\mathbb{G},r,M}^{j}\}$ and the coherent sheaves $\mathcal{K}\text{\it er}\{(\tilde{\Theta}_{\mathfrak{g},r,M}^{exp})^{j}\},\\ \mathcal{C}\text{\it oker}\{(\tilde{\Theta}_{\mathbb{G},r,M}^{exp})^{j}\},\ \mathcal{I}\text{\it m}\{(\tilde{\Theta}_{\mathbb{G},r,M}^{exp})^{j}\}$.
Definition 6.5.
Let $(\mathbb{G},\mathcal{E}_{(r)})$ be an affine group scheme of exponential type of height $r$. Consider the bigraded algebra $k[s,t]\otimes k[\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))]$ where both $s\otimes 1$ and $t\otimes 1$ are given the bigrading $(1,0)$. We define $\Theta_{\mathfrak{g}/\mathbb{P}^{1},r}$ to be
(6.5.1) $\Theta_{\mathfrak{g}/\mathbb{P}^{1},r}\quad\equiv\quad(s\otimes 1)(1\otimes\Theta_{\mathfrak{g},r}^{\exp})\ +(t\otimes 1)(1\otimes\Phi_{\mathfrak{g},r}^{*}(\Theta_{\mathbb{G}_{(r)}}))\ \in\ k[s,t]\otimes k[\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))]\otimes k\mathbb{G}_{(r)}.$
So defined, $\Theta_{\mathbb{G}/\mathbb{P}^{1},r}$ is bihomogenous of bidegree $(1,p^{r-1})$, thereby determining a global section of $\mathcal{O}_{\mathbb{P}^{1}\times\mathbb{P}\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))}(1,p^{r-1})$.
Proposition 6.6.
Retain the notation and hypotheses of Definition 6.5. Let $M$ be a finite dimensional $\mathbb{G}_{(r)}$-module. For any scheme theoretic point $(\underline{B},\langle s,t\rangle)\in\mathcal{C}_{r}(\mathfrak{g})\times\mathbb{P}^{1}$,
$JT(\Theta^{exp}_{\mathfrak{g},r,\underline{B}},k(\underline{B})\otimes M)\ =\ \frac{m}{p}[p]\ \ \iff\\ JT(\Theta_{\mathfrak{g}/\mathbb{P}^{1},r,(\underline{B},\langle s,t\rangle)},k(\underline{B},\langle s,t\rangle)\otimes M)\ =\ \frac{m}{p}[p].$
Proof.
Observe that
$\Theta_{\mathfrak{g}/\mathbb{P}^{1},r,(\underline{B},\langle s,t\rangle)}\ =\ s\cdot\Theta_{\mathfrak{g},r,\underline{B}}^{\exp}\ +\ t\cdot\Theta_{\mathfrak{g},r,\underline{B}}\quad\in\quad k(\underline{B},\langle s,t\rangle)\otimes k\mathbb{G}_{(r)}.$
As recalled in the proof of Proposition 4.13, $\Theta_{\mathfrak{g},r,\underline{B}}-\Theta_{\mathfrak{g},r,\underline{B}}^{\exp}$ is a linear combination of terms each of which is a product of at least 2 $p$-nilpotent factors. Thus, for any $\langle s,t\rangle\in\mathbb{P}^{1}$, $s\cdot\Theta_{\mathfrak{g},r,\underline{B}}^{\exp}\ +\ t\cdot\Theta_{\mathfrak{g},r,\underline{B}}$ differs from $\Theta_{\mathfrak{g},r,\underline{B}}^{\exp}$ by a sum of terms each of which is a product of at least 2 $p$-nilpotent factors. Consequently, the proposition follows from [8, Prop 8]. ∎
If $(\mathbb{G},\mathcal{E}_{(r)})$ be an affine group scheme of exponential type of height $r$, if $j$ is a positive integer between 1 and r1r-1, and if $M$ is a finite dimensional $\mathbb{G}_{(r)}$-module, then $M$ is said to have constant $j$-rank if for every scheme-theoretic point $\underline{B}\in\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))$ the rank of $\Theta_{\mathfrak{g},r,\underline{B}}^{j}:k(\underline{B})\otimes M\to k(\underline{B})\otimes M$ is independent of $\underline{B}$. (See [16, Defn 3.12].) Thus, $M$ has constant Jordan type if and only if $M$ has constant $j$-type for each $j,\ 1\leq j<r$.
Proposition 6.7.
Let $(\mathbb{G},\mathcal{E}_{(r)})$ be an affine group scheme of exponential type of height $r$, and let $M$ be a $\mathbb{G}_{(r)}$-module. Then for any $j,0<j<p$, there are quasi-coherent sheaves ${\mathcal{K}\text{\it er}}\{(\tilde{\Theta}_{\mathfrak{g}/\mathbb{P}^{1},r,M})^{j}\},\ {\mathcal{C}\text{\it oker}}\{(\tilde{\Theta}_{\mathfrak{g}/\mathbb{P}^{1},r,M})^{j}\},\ {\mathcal{I}\text{\it m}}\{(\tilde{\Theta}_{\mathfrak{g}/\mathbb{P}^{1},r,M})^{j}\}$ on $\mathbb{P}^{1}\times\mathbb{P}\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))$ satisfying
  1. (1)
    The fibers over $\langle 0,t\rangle\in\mathbb{P}^{1}$ of these sheaves are naturally identified with the sheaves on $\mathbb{P}\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))$ given by the pull-backs along $\Phi_{\mathfrak{g},r}^{*}$ of the corresponding sheaves on $\mathbb{P}V_{r}(\mathbb{G})$ for $\tilde{\Theta}_{\mathbb{G},r,M}^{j}$.
  2. (2)
    The fibers over $\langle s,0\rangle\in\mathbb{P}^{1}$ of these sheaves are naturally identified with the corresponding sheaves on $\mathbb{P}\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))$ for $(\tilde{\Theta}_{\mathbb{G},r,M}^{exp})^{j}$.
  3. (3)
    If $M$ is a finite dimensional $\mathbb{G}_{(r)}$-module of constant $j$-rank, then these sheaves are locally free, coherent sheaves on $\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))\times\mathbb{P}^{1}$.
In other words, for any $\mathbb{G}_{(r)}$-module of constant $j$-type, taking kernels, cokernels, and images of $(\tilde{\Theta}_{\mathbb{G},r,M})^{j}$ and $(\tilde{\Theta}_{\mathfrak{g},r,M}^{exp})^{j}$ yield $\mathbb{A}^{1}$-homotopy equivalent vector bundles on $\mathbb{P}\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))$.
Proof.
By Proposition 4.5 and Theorem 4.8, $\Theta_{\mathfrak{g}/\mathbb{P}^{1},r}$ of Definition 6.5 is bihomogeneous of bidegree $(1,p^{r-1})$. Thus, the bigraded module
(6.7.1) $ker\{(\Theta_{\mathfrak{g}/\mathbb{P}^{1},r})^{j}:(k[s,t]\otimes k[\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))]\otimes M)\ \to\ (k[s,t]\otimes k[\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))]\otimes M)[1,p^{r-1}]$
determines the quasi-coherent sheaves ${\mathcal{K}\text{\it er}}\{(\tilde{\Theta}_{\mathfrak{g}/\mathbb{P}^{1},r,M})^{j}\},\ {\mathcal{C}\text{\it oker}}\{(\tilde{\Theta}_{\mathfrak{g}/\mathbb{P}^{1},r,M})^{j}\},\\ {\mathcal{I}\text{\it m}}\{(\tilde{\Theta}_{\mathfrak{g}/\mathbb{P}^{1},r,M})^{j}\}$ on $\mathbb{P}^{1}\times\mathbb{P}\mathcal{C}_{r}(\mathfrak{g})$. The first two assertions follow from the observation that the fiber above $\langle a,b\rangle\in\mathbb{P}^{1}$ of a sheaf on $\mathbb{P}^{1}\times\mathbb{P}\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))$ is the sheaf associated to the bigraded graded module (6.7.1) obtained by extending along $k[s,t]\to k(\langle a,b\rangle)\in\mathbb{P}^{1}$ which sends $s$ to $a$ and $t$ to $b$.
For $M$ a finite dimensional $\mathbb{G}_{(r)}$-module of constant $j$-type, the verification of [16, Thm 4.13] showing that $\mathcal{K}\text{\it er}\{\tilde{\Theta}_{\mathbb{G},r,M}^{j}\},\ \operatorname{Coker}\nolimits\{\tilde{\Theta}_{\mathbb{G},r,M}^{j}\},\ \mathcal{I}\text{\it m}\{\tilde{\Theta}_{\mathbb{G},r,M}^{j}\}$ are locally free sheaves of $\mathbb{P}V_{r}(\mathbb{G})$ applies equally to verify that the sheaves $\mathcal{K}\text{\it er}\{(\tilde{\Theta}_{\mathfrak{g},r,M}^{exp})^{j}\},\\ \mathcal{C}\text{\it oker}\{(\tilde{\Theta}_{\mathbb{G},r,M}^{exp})^{j}\},\ \mathcal{I}\text{\it m}\{(\tilde{\Theta}_{\mathbb{G},r,M}^{exp})^{j}\}$ are locally free sheaves on $\mathbb{P}\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))$, and applies with only notational changes to show that ${\mathcal{K}\text{\it er}}\{(\tilde{\Theta}_{\mathfrak{g}/\mathbb{P}^{1},r,M})^{j}\},\ {\mathcal{C}\text{\it oker}}\{(\tilde{\Theta}_{\mathfrak{g}/\mathbb{P}^{1},r,M})^{j}\},\\ {\mathcal{I}\text{\it m}}\{(\tilde{\Theta}_{\mathfrak{g}/\mathbb{P}^{1},r,M})^{j}\}$ are locally free sheaves on $\mathbb{P}^{1}\times\mathbb{P}\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))$
Remark 6.8.
As shown in [4], most vector bundles on $\mathbb{P}^{r-1}$ can be realized by kernel, cokernel, and image sheaves associated to modules of constant Jordan type for elementary abelian $p$-groups of rank $r$. For each such vector bundle on $\mathbb{P}^{r-1}$, there is a corresponding vector bundle on $w\mathbb{P}(1,p,\ldots,p^{r-1})$ associated to a module of constant Jordan type for $\mathbb{G}_{a(r)}$. It would be interesting to find examples in the context of Proposition 6.7(3) of locally free sheaves on $\mathbb{P}\mathcal{C}_{r}(\mathfrak{g})$ which are $\mathbb{A}^{1}$-equivalent but not isomorphic.
One possible source of such examples might come from $\mathbb{G}=\mathbb{G}_{a(r)}$, $I\subset k\mathbb{Z}/p^{\times r}$ the augmentation of $k\mathbb{Z}/p^{\times r}$, and $M$ the $\mathbb{G}$-module corresponding to the $k\mathbb{Z}/p^{\times r}$-module $I^{m}/I^{n}$ with $n>m$. In this case, we would be comparing $\mathbb{A}^{1}$-equivalent vector bundles on the singular variety $w\mathbb{P}(1,p,\ldots,p^{r-1})$. An approach might be to look at local invariants of these bundles at singular points of $w\mathbb{P}(1,p,\ldots,p^{r-1})$.
7. Jordan types for linear algebraic groups
In [12, Defn 4.4], we formulated the local Jordan type of a finite dimensional rational $\mathbb{G}$-module $M$ for $\mathbb{G}$ a linear algebraic group equipped with a structure of exponential type. There is some subtlety to this construction in order to “package” the Jordan types of $M_{|\mathbb{G}_{(r)}}$ of restrictions of $M$ to all Frobenius kernels of $\mathbb{G}$. This involves reordering $r$-tuples of $p$-nilpotent, pairwise commuting elements $(B_{0},\ldots,B_{r-1})$ of $\mathfrak{g}$ so that they appear as $(B_{r-1},\ldots,B_{0})$. The definition of the Jordan type of a finite dimensional rational $\mathbb{G}$-module $M$ is then formulated in terms of the action of a colimit of $p$-nilpotent operators, rather than as a colimit of Jordan types of restrictions $M_{|\mathbb{G}_{(r)}}$ of $M$ to the Frobenius kernels $\mathbb{G}_{(r)}$ of $\mathbb{G}$.
In this final section, we observe that this approach leads to Jordan type functions for finite dimensional rational $\mathbb{G}$-modules $M$ provided that $\mathbb{G}$ is of “exponential type of infinite height”. In Definition 7.3, we extend the construction of $JT^{exp}_{\mathfrak{g},r,M}(-):\mathcal{C}_{r}(N_{p}(\mathfrak{g}))\ \to\ \mathcal{Y}$ of Definition 4.16 for $\mathbb{G}_{(r)}$-modules to $JT_{\mathbb{G},\infty,M}^{exp}(-):\mathcal{C}_{\infty}(\mathfrak{g})\to\mathcal{Y}$ for finite dimensional rational $\mathbb{G}$-modules. As shown in Proposition 7.4 (correcting an error of [12, Prop 4.8]), $JT_{\mathbb{G},\infty,M}^{exp}(-)$ is a twisted stabilization of $JT_{\mathbb{G},r,M_{|\mathbb{G}_{(r)}}}^{exp}(-)$.
[…]
where $(B_{0},\ldots,B_{r-1})\mapsto(B_{0},\ldots,B_{r-1},0).$ We define the “coordinate algebra”
[…]
We define a scheme-theoretic (respectively, geometric) point of $\mathcal{C}_{\infty}(\mathcal{N}_{p}(\mathfrak{g}))$ to be an element in the direct limit of scheme-theoretic (respectively, geometric) points of $\{\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))\}$. We can view such a point $\iota_{x}:\operatorname{Spec}\nolimits K\to\mathcal{C}_{\infty}(\mathcal{N}_{p}(\mathfrak{g}))$ as given by a “continuous” map of algebras $\iota_{x}^{*}:k[\mathcal{C}_{\infty}(\mathcal{N}_{p}(\mathfrak{g}))]\to K$.
We provide analogues of the constructions given in Construction 4.3.
Definition 7.1.
Let $\mathbb{G}$ be a linear algebraic group with Lie algebra $\mathfrak{g}$, equipped with a morphism
$\mathcal{E}:\mathbb{G}_{a}\times\mathcal{N}_{p}(\mathfrak{g})\quad\to\quad\mathbb{G}$
such that $(\mathbb{G}_{(r)},\ \mathcal{E}_{|r})$ constitutes an affine group scheme of exponential type of height $r$ for each $r>0$. Here, $\mathcal{E}_{|r}:\mathbb{G}_{a(r)}\times\mathcal{N}_{p}(\mathfrak{g})\to\mathbb{G}_{(r)}$ is the restriction of $\mathcal{E}$. We say that $\mathbb{G}$ is of exponential type of infinite height.
For any (geometric) point $B$ of $\mathcal{N}_{p}(\mathfrak{g})$, we denote by $\mathcal{E}_{|r,B}:k(B)\times\mathbb{G}_{a(r)}\ \to\ k(B)\otimes\mathbb{G}$ the restriction of $\mathcal{E}_{|r}$ along $\operatorname{Spec}\nolimits k(B)\to\mathcal{N}_{p}(\mathfrak{g})$.
We define
$\mathcal{E}_{\infty|r}\ =\ \mathcal{E}_{|r}\circ(1\times pr_{s}):\mathbb{G}_{a(r)}\times\mathcal{C}_{\infty}(\mathcal{N}_{p}(\mathfrak{g}))\ \to\ \mathbb{G}_{a(r)}\times\mathcal{N}_{p}(\mathfrak{g})\ \to\ \mathbb{G}_{(r)}$
and set
$\mathcal{U}_{\mathfrak{g},\infty|r}\ \equiv\ \mathcal{E}_{\infty|r}\times pr_{\mathcal{C}_{\infty}(\mathcal{N}_{p}(\mathfrak{g}))}:\mathbb{G}_{a(r)}\times\mathcal{C}_{\infty}(\mathcal{N}_{p}(\mathfrak{g}))\ \to\ \mathbb{G}_{(r)}\times\mathcal{C}_{\infty}(\mathcal{N}_{p}(\mathfrak{g})).$
We define the operator (a formal limit of well defined partial sums)
(7.1.1) $\Theta_{\mathfrak{g},\infty}^{exp}\ \equiv\ \sum_{s\geq 0}(\mathcal{U}_{\mathfrak{g},\infty|s})_{*}(1\otimes u_{s})\ =\sum_{s\geq 0}\in\ k[\mathcal{C}_{\infty}(\mathcal{N}_{p}(\mathfrak{g}))]\otimes k\mathbb{G}.$
As verified in [12, Prop 2.6], for any linear algebraic group $(\mathbb{G},\mathcal{E})$ of exponential type of infinite height, any finite dimensional $\mathbb{G}$-module $M$, any element $m\in M$, and any $B\in\mathcal{N}_{p}(\mathfrak{g})$, the element $(\mathcal{E}_{|r,B_{s}})_{*}(1\otimes u_{s})$ in $k[\mathcal{C}_{\infty}(\mathcal{N}_{p}(\mathfrak{g}))]\otimes k\mathbb{G}$ acts trivially on $m$ for all sufficiently large $s$. This justifies the following proposition which should be compared with Definition 4.16.
Proposition 7.2.
For any finite dimensional $\mathbb{G}$-module $M$, the formal sum $\Theta_{\mathfrak{g},\infty}^{exp}$ of (7.1.1) determines a well defined $p$-nilpotent $k[\mathcal{C}_{\infty}(\mathfrak{g})]$-linear operator
$\Theta_{\mathfrak{g},\infty}^{exp}:k[\mathcal{C}_{\infty}(\mathcal{N}_{p}(\mathfrak{g}))]\otimes M\ \to\ k[\mathcal{C}_{\infty}(\mathcal{N}_{p}(\mathfrak{g}))]\otimes M.$
The specialization at a scheme-theoretic point $\underline{B}\in\mathcal{C}_{\infty}(\mathcal{N}_{p}(\mathfrak{g}))$ of $\Theta_{\mathfrak{g},\infty}^{exp}$ determines the operator
$\Theta_{\mathfrak{g},\infty,\underline{B}}^{exp}\ =\ \sum_{s\geq 0}(\mathcal{E}_{|s,B_{s}})_{*}(1\otimes u_{s-1}):k[\underline{B}]\otimes M\ \to\ k[\underline{B}]\otimes M,$
(denoted $\sum_{s\geq 0}(\mathcal{E}_{B_{s}})_{*}(u_{s})$ in [12]).
The following definition of $JT_{\mathbb{G},\infty,M}(\underline{B})$ is essentially that of [12, Defn 4.4].
Definition 7.3.
For any finite dimensional $\mathbb{G}$-module $M$, we define the Jordan type of $M$ at $\underline{B}\in\mathcal{C}_{\infty}(\mathfrak{g})$ to be
(7.3.1) $JT_{\mathfrak{g},\infty,M}^{exp}(\underline{B})\quad\equiv\quad JT(\sum_{s\geq 0}(\mathcal{E}_{|s,B_{s}})_{*}(1\otimes u_{s}),k(\underline{B})\otimes M),$
the Jordan type of the $p$-nilpotent action of $k(\underline{B})\otimes_{k[\mathcal{C}_{\infty}(\mathcal{N}_{p}(\mathfrak{g}))]}\Theta_{\mathfrak{g},\infty}^{exp}\in k(\underline{B})\otimes k\mathbb{G}$ acting on $k(\underline{B})\otimes M$.
Thus, (7.3.1) determines the Jordan type function
$JT_{\mathfrak{g},\infty,M}^{exp}(-):\ \mathcal{C}_{\infty}(\mathcal{N}_{p}(\mathfrak{g}))\quad\to\quad\mathcal{Y}.$
We point out that [12, Prop 4.8] is wrong as stated. The following proposition provides a correction. The proof of Proposition 7.4 is an immediate consequence of the observation above that $(\mathcal{E}_{|r,B_{s}})_{*}(1\otimes u_{s})$ in $k[\mathcal{C}_{\infty}(\mathfrak{g})]\otimes k\mathbb{G}$ acts trivially on $m\in M$ for all sufficiently large $s$.
Proposition 7.4.
Adopt the notation and hypotheses of Definition 7.3. If $M$ is finite dimensional, then
$JT_{\mathbb{G},\infty,M}^{exp}(-)\ =\ JT_{\mathbb{G},r,M_{|\mathbb{G}_{(r)}}}^{exp}(-)\circ\Lambda^{r}:\ \mathcal{C}_{\infty}(\mathcal{N}_{p}(\mathfrak{g}))\ \to\ \mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))\ \to\ \mathcal{Y},\quad r\gg 0,$
where $\Lambda^{r}:\mathcal{C}_{\infty}(\mathcal{N}_{p}(\mathfrak{g}))\to\mathcal{C}_{r}(\mathcal{N}_{p}(\mathfrak{g}))$ is given by sending $(B_{0},\ldots,B_{n},\ldots)$ to $(B_{r-1},\ldots,B_{0})$.
In particular, $JT_{\mathbb{G},\infty,M}(-)$ is continuous.
One can interpret the Proposition 7.4 as saying for $M$ a finite dimensional $\mathbb{G}$-module that the Jordan type function $JT_{\mathbb{G},\infty,M}^{exp}(-)$ is a twisted stabilization (with respect to increasing $r$) of the exponential Jordan type functions $JT_{\mathbb{G},r,M_{|\mathbb{G}_{(r)}}}^{exp}(-)$ of Definition 4.16.